### v2

# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models
Savi Virolainen
University of Helsinki
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss labelling of the shocks, maximum likelihood estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty.
Keywords: smooth transition vector autoregression, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P. O. Box 17, FI–00014 University of Helsinki, Finland; e-mail: savi.virolainen@helsinki.fi. ORCiD ID: $0$-0002-5075-6821.
1 Introduction
Smooth transition vector autoregressive (STVAR) models (Anderson and Vahid, 1998) have been widely adopted in empirical macroeconomics due to their ability to flexibly capture nonlinear data generating dynamics and gradual shifts in parameter values. Such variation may arise due to wars, business cycle fluctuations, or policy shifts, for example. Structural STVAR models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock. In order to estimate their effects, the economic shocks need to be identified. Conventional identification methods typically rely on restrictive assumptions such as zero contemporaneous interactions among some of the variables or long-term neutrality of the shocks. Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification. To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
There are two main branches in the statistical identification literature: identification by heteroskedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, forthcoming, and others) and identification by non-Gaussianity (Lanne et al., 2017, Maxand, 2020, Lanne and Luoto, 2021, Lanne et al., 2023, Anttonen et al., forthcoming, and others). Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroskedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see the discussion in Virolainen, forthcoming, Sections 3.1 and 4). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses. Bacchiocchi and Fanelli (2015) do propose an alternative framework for identification by heteroskedasticity that allows for variation in the impact effects of the shocks across the volatility regimes, but their method requires imposing a number of untestable zero impact effect restrictions that can be difficult to justify economically. To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to STVAR models. We also consider threshold vector autoregressive models of Tsay (1998) as special cases of STVAR models. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Similarly to Lanne et al. (2017), however, it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations.
In line with statistical identification literature, labelling the identified structural shocks as economic shocks requires external information. This requires labelling the same shock to the same column of the impact matrix across the regimes. In our experience, however, the identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes. This can be problematic because if the shocks are ordered differently in some regimes, the (generalized) impulse response functions do not estimate the effects of the intended shock but of a linear combination of structural shocks. As a solution, we propose a procedure that utilizes the almost equally well fitting local solutions with different orderings and signs of the columns of the impact matrices obtainable by our estimation method. If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). Morioka et al. (2021) propose using IIA for directly estimating the (conditionally) contemporaneously independent innovations of a general nonlinear vector autoregressive (VAR) model without assuming a specific functional form for the model. They show that under a certain set of (fairly general) assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. In applied macroeconometrics, the main interest is, however, typically in computing the (generalized) impulse response functions or performing other analyses that require obtaining an estimate for the parameters of the model. Estimating the parameters of a STVAR model using the framework of Morioka et al. (2021) requires first estimating the independent innovations using a specific machine learning algorithm (see Morioka et al., 2021), and then using the estimates of the innovations to obtain estimates of the parameters. However, it is not obvious how to estimate the parameters of the model from the innovations obtained by IIA, as they are identified only up to permutation and component-wise invertible transformations, and different parameter estimates would be obtained with different transformations. Moreover, Morioka et al. (2021) assume that the innovations are from a distribution in an exponential family, thereby excluding various distributions such as the popular Student’s $t$ distribution. Our approach, in turn, facilitates estimating the parameters directly by the method of maximum likelihood, and it does not exclude distributions of the independent shocks other than Gaussian distribution and those with infinite variance.
Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S. data from 1987:4 to 2024:2. Following Gavriilidis et al. (2023), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to climate policy uncertainty. We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty. Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable, and allowing for smooth transitions in the intercept parameters as well as in the impact matrix. We find that a positive CPU shock decreases production in times of low economic policy uncertainty and slightly increases it in times of high economic policy uncertainty. Consumer prices, in turn, increase in response to a positive CPU shock more significantly when the level of economic policy uncertainty is high than when it is low. Our results are, hence, complementary to Gavriilidis et al. (2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and describes a procedure for labelling the shocks in the presence of weak identification. Section 4 discusses estimation of the parameters by the method of maximum likelihood and stationarity of the model. Section 5 presents the empirical application and Section 6 concludes. Appendices provide further discussion on stationarity of the model; proofs for the stated lemmas, propositions, and theorem; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
2 Reduced form STVAR models
Let $y_{t}$, $t=1,2,...$, be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$. We consider STVAR models with $M$ regimes and autoregressive order $p$ assumed to satisfy
$y_{t}$ $=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu),$ (2.1)
$\mu_{m,t}$ $=\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either nonrandom and exogenous or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and they satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
It is easy to see that, conditional on $\mathcal{F}_{t-1}$, the conditional mean of the above-described process is $\mu_{y,t}\equiv E[y_{t}|\mathcal{F}_{t-1}]=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$, a weighted sum the regime-specific means $\mu_{m,t}$ with the weights given by the transition weights $\alpha_{m,t}$. The conditional covariance matrix is $\Omega_{y,t}$, but discussion on its specification is postponed to Section 3. See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
It is often assumed that the coefficient matrices $A_{m,i}$, $i=1,...,p$, satisfy the usual stability condition $\det\left(I_{d}-\sum_{i=1}^{p}A_{m,i}z^{i}\right)\neq 0\ \ \text{for}\ \ |z|\leq 1,$ for each of the regimes $m=1,...,M$. However, this condition does not necessarily ensure that the model is stationary (see Kheifets and Saikkonen, 2020, and the references therein). If the transition weights are exogenous, it might be particularly challenging to establish the stationarity of the model. Assuming endogenous transition weights, we present sufficient conditions for establishing ergodicity, stationary, and mixing properties of our structural STVAR model in Appendix A (see also Section 4). Nevertheless, our main identification results discussed in Section 3 do not require stationarity of the model.
The above-described framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right.$ (2.3)
where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable. At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be either exogenous or a lagged endogenous variable.
Another popular specification is to assume two regimes ($M=2$) and logistic transition weights. In the logistic STVAR model (Anderson and Vahid, 1998), the transition weights vary according to a logistic function as
$\alpha_{1,t}=1-\alpha_{2,t}\ \ \text{and}\ \ \alpha_{2,t}=[1+\exp\{-\gamma(z_{t}-c)\}]^{-1},$ (2.4)
where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the mid point of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$. The switching variable $z_{t}$ can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables.
In addition specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
$y_{t}=\phi_{0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), so it is more feasible in large scale VAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
Imposing constant AR matrices across the regimes but allowing the intercepts to vary in time is a particularly useful special case of (2.1)-(2.2), as constant AR matrices substantially decrease the number of parameters compared to the fully nonlinear specification. However, the model still allows for a time-varying conditional mean via time-varying intercepts as well as shifts in the volatility regime through the error process $u_{t}$. This specification is obtained from (2.1)-(2.2) by assuming $A_{m,i}=A_{i}$ for all $i=1,...,p$ and $M=1,...,M$, simplifying the model to:
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.6)
Finally, note that while we focus on TVAR and STVAR models here for concreteness and because these models are implemented to the accompanying R package sstvars (Virolainen, 2024), the identification results of this paper directly extend to many other nonlinear SVAR models as well. This includes models incorporating discrete regime switches such as Markov-Switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007), and more generally models whose impact matrix is defined as a weighted sum of the impact matrices of the regimes and the time $t$ weights are known at the time $t$.
3 Structural STVAR models identified by non-Gaussianity
3.1 Structural STVAR models
A structural STVAR model is obtained from the reduced form model defined in Section 2 by identifying the serially and mutually uncorrelated structural shocks $e_{t}=(e_{1t},...,e_{dt})$ $(d\times 1)$ from the reduced form innovations $u_{t}$. Specifically, the structural shocks are recovered from the reduced form shocks with the transformation
$e_{t}=B_{y,t}^{-1}u_{t},$ (3.1)
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$. Various types of restrictions that recover the structural shocks have been proposed in the literature. Many of them are discussed in Ramey (2016) and Kilian and Lütkepohl (2017), for example.
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian. With Gaussian shocks, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ lead to observationally equivalent models. However, when the shocks are mutually independent and at most of them is Gaussian, different decompositions generally lead to models that are not observationally equivalent, as shown in Lanne et al. (2017) for a static impact matrix. We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with the impact matrix $B_{y,t}$.
In order to parametrize the model directly with the impact matrix, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
where $e_{t}$ are independent and identically distributed structural errors with identity covariance matrix and a distribution that may depend on the parameter $\nu$.
Since we incorporate time-variation in the impact matrix, its functional form needs to be specified. A natural specification for a STVAR model is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
$B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m},$ (3.3)
where $B_{1},...,B_{M}$ are invertible $(d\times d)$ impact matrices of the regimes. Also the impact matrix $B_{y,t}$ needs to be invertible to ensure positive definiteness of the conditional covariance matrix, which does not automatically follow from the invertibility of $B_{1},...,B_{M}$. Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix B.1).
Lemma 1.
Suppose $B_{1},...,B_{M}$ are invertible $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and $m=1,...,M$. Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ that are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, observe that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to an impact matrix $B_{t}$ that is singular for some $t=1,...,T$ are excluded.
Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
$\Omega_{y,t}=\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}^{2}\Omega_{m}+\sum_{m=1}^{M}\sum_{n=1,n\neq m}^{M}\alpha_{m,t}\alpha_{n,t}\Omega_{m,n},$ (3.4)
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thereby described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. To facilitate the interpretation of the conditional covariance matrix (3.4), suppose the process is completely in one of the regimes at some $t$, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$ and $\alpha_{m,t}=0$ for $m\neq\tilde{m}$. Then, $\Omega_{y,t}$ reduces to $B_{\tilde{m}}B_{\tilde{m}}^{\prime}=\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of the regime $\tilde{m}$. When the process is not completely in any of the regimes, $\alpha_{\tilde{m},t}\in(0,1)$ for some regime $\tilde{m}$$\Omega_{y,t}$ depends both on the conditional covariance matrices of the regimes $\Omega_{1},...,\Omega_{M}$ and on the cross terms $\Omega_{m,n}$. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ has the form $B_{y,t}B_{y,t}^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for the covariance matrices $\Omega_{1},...,\Omega_{M}$ of the regimes.
Specifying the conditional covariance matrix as $\Omega_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ would either require parametrizing the model directly with $\Omega_{y,t}$ or imposing some specific structure on $B_{y,t}$ that would ensure that this identity holds at all $t$. In the former case, identification of the shocks requires further restrictions to be imposed on the model (even when the shocks are non-Gaussian) as discussed at the beginning of this section. In the latter case, the restrictions are implicitly imposed by the specific form of the impact matrix ensuring that $B_{y,t}B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for all $t$. One of such parametrizations of $B_{y,t}$ aligns with the identification by heteroskedasticity (Lütkepohl and Netšunajev, 2017), which restricts the relative impact responses of the variables to each shock time-invariant (see also the related discussion in Virolainen, forthcoming). In either case, the identifying restrictions may be challenging to justify economically. Our specification of $\Omega_{y,t}$ in (3.3) and (3.4), in turn, facilitates exploiting non-Gaussianity of the shocks in their identification, as we show the next section.
3.2 Identification of the shocks by non-Gaussianity
The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are a mutually independent and at most one of them is Gaussian. We also assume that $e_{t}$ is an IID sequence with zero mean and identity covariance matrix as is formally stated in the following assumption.
Assumption 1.
  1. (i)
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, has zero mean and unit variance.
  2. (ii)
    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one them has a Gaussian marginal distribution.
Note that the assumption of a unit variance of $e_{it}$, $i=1,...,d$, is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
The next lemma then shows that, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns at each $t$. This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent. This lemma, which is mostly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the parameters in the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
Lemma 2.
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Lemma 2 is proven in Appendix B.2.
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering and signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to a model that is not observationally equivalent. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, and it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. In order to fix the ordering and signs of the columns of $B_{y,t}$ it then suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$. This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending of the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others. Also, constraints that are identifying in some specifications may not be sufficient for identification in others. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
3.3 Identification of the shocks in structural TVAR models
We consider structural threshold VAR (TVAR) models here, as they are special cases of STVAR models, but the introduced identification results also directly extend to other models incorporating discrete regime switches, such as Markov-switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007). Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$. Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$. That is, we have structural TVAR model. By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained by fixing the ordering and signs of the columns of $B_{1},...,B_{M}$. However, whether this is overidentifying depends on the distributions of the shocks $e_{t}=(e_{1t},...,e_{dt})$.
For instance, suppose the shocks follow independent Student’s $t$ distributions. Then, the signs of the columns of $B_{m}$ can be fixed in each regime separately without loss of generality due to the symmetry of the Student’s $t$ distribution. But fixing the ordering of the columns in each regime separately would generally be overidentifying because the $i$th column of $B_{m}$ is related to the $i$th shock, and each shock has its own degrees-of-freedom parameter that affects the density of the distribution (and is in our specification shared across the regimes). Fixing the ordering of the columns of $B_{1}$ would thereby effectively fix the ordering of the columns of $B_{2},...,B_{M}$ without restricting the model when the values of the degrees-of-freedom parameter are not equal for any of the shocks. On the other hand, if the degrees-of-freedom parameter values of some of the shocks are equal, the ordering of the corresponding columns can be changed in each regime separately without changing the underlying model. If the degrees of freedom parameter values are close to each other, reordering the columns of $B_{m}$ in one of the regimes without reordering them in the other regimes accordingly would have only a small effect of the density of the model (or to the fit of an estimated model), thereby resulting in weak identification with respect to the ordering of the columns of $B_{2},...,B_{M}$.
As the required identification conditions depend on the distributions of the shocks, it useful to make some assumptions on them. Namely, we assume that the distributions of the shocks are continuous and none of the them are identical to each other in the sense that the values of their density functions are unequal almost everywhere in $\mathbb{R}$. This assumption allows to effectively fix the ordering of the columns of $B_{2},...,B_{M}$ by fixing the ordering of the columns of $B_{1}$, since changing the ordering would change the distribution of the corresponding shocks in some of the regimes relative to Regime 1. If this assumption is close to break (e.g., when the degrees-of-freedom parameter values of independent Student’s $t$ distributions are close to each other or very large), identification is weak with respect to the ordering of the columns. Labelling the shocks based on the estimates of the columns of each $B_{m}$ can then be unreliable, as small changes in the data could result in changes in the ordering of columns in some of the regimes. Therefore, we propose a procedure to facilitate reliable labelling of the shocks in the presence of such weak identification in Section 3.5.
The following proposition, which is proved in Appendix B.3, formally states the identification results for TVAR models.
Proposition 1.
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied and $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$. Denote the density function of the distribution of the shock $e_{it}$ as $d_{e_{i}}(\cdot;0,1,\nu_{i})$, where $\nu_{i}$ denotes the parameters of the distribution other than mean and variance. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and assume the following conditions hold:
  1. (a)
    the signs of the columns of $B_{m}$ are fixed for each $m=1,...,M$,
  2. (b)
    the ordering of the columns of $B_{1}$ is fixed, and
  3. (c)
    $e_{it}$ follow continuous distributions with $d_{e_{i}}(x;0,1,\nu_{i})\neq d_{e_{j}}(x;0,1,\nu_{j})$ almost everywhere in $x\in\mathbb{R}$ for $i\neq j=1,....,d$.
Then, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
Condition (a) of Proposition 1 grants arbitrary identification with respect to the signs of the shocks. If the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, any fixed signs can be assumed without loss of generality. However, if the distributions of the shocks are not symmetric, the signs of the columns of the impact matrices may affect the density of the model, implying that fixing the signs is an overidentifying restriction. Condition (b) fixes the ordering of the columns of $B_{1}$. Any fixed ordering of the columns can be assumed without loss generality. For instance, if the first nonzero entry in each column of $B_{1}$ is assumed positive, Condition (a) is satisfied by the positivity assumption, and Condition (b) is satisfied by assuming that these entries are in a decreasing order, given that none of the entries are exactly equal to each other.
Condition (c) is concerned with the distributions of the shocks, and it is satisfied by many of the popular distributions such as Student’s $t$ distributions with different values of the degrees-of-freedom parameters. It ensures that changing the ordering of the columns of $B_{2},...,B_{M}$ while keeping the ordering of the columns of $B_{1}$ fixed would change the joint density of the model. Thus, combined with Condition (b), it leads to identification with respect to ordering of the columns of $B_{1},...,B_{M}$. Finally, the result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ would not change the joint density of the model. Since the probability of obtaining such special cases is zero, this does not have any practical implications.
3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, or more generally to models incorporating discrete regime switches, we now assume that the transition weights exhibit more variation. Because the variation in the transition weights then translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns (by Lemma 2), the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1. Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Proposition 2.
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Suppose the following conditions hold:
  1. (A)
    the ordering and signs of the columns of $B_{1}$ are fixed,
  2. (B)
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero, and
  3. (C)
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values.
Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that there can exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. Switching the ordering of the columns of $B_{y,t}$ merely equals switching the ordering of the shocks, so any fixed ordering of the columns of $B_{1}$ can be assumed without loss of generality. Furthermore, if the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, also the signs of the columns of $B_{1}$ can be fixed without loss of generality, as switching the sign of a column of $B_{y,t}$ equals switching the sign of the corresponding shock. The ordering and signs of the columns of $B_{1}$ can be fixed, for example, by assuming that the first nonzero element in each column of $B_{1}$ is positive, these elements are not exactly equal to each other, and they are in a decreasing ordering. The assumption of unequal first nonzero elemenents is not restrictive in practice, as the probability of obtaining an estimate such that some of the elements are exactly equal to each other within the numerical accuracy of modern statistical software is extremely small. See Lanne et al. (2017) for a more complicated strategy for fixing the ordering and signs (or more generally scalings) of the columns of a static impact matrix that is valid without such an additional assumption.
Condition (B) of Proposition 2 states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete (in which case Proposition 1 would be more relevant). Condition (C), in turn, states that the transition weights should exhibit at least a certain small degree of variation. For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$. In practice, Condition (C) should be satisfied for the observed time period $t=1,...,T$. This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete).
The concerns of weak identification with respect to the ordering (and signs) of the columns of $B_{2},...,B_{M}$ with binary transition weights, discussed in Section 3.3, also translate to smooth transition models when the estimates of the transition weights are such that the weights are mostly close to zero or one. Surprisingly, our experience shows that even when there is a high degree of variation in the transition weights, the identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ often (if not virtually always) appears to be very weak in practice, at least if the parameters are estimated by the method of maximum likelihood. Specifically, there are frequently a number of local maxima of the log-likelihood function with log-likelihoods very close to each other, corresponding to mostly very similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
Weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ can be a problem in practice for labelling the shocks, as a given column in $B_{y,t}$ need not correspond to the same shock for all $t=1,...,T$. In this case, the (generalized) impulse response functions would not estimate of the effects of the intended shock but of a linear combination of the structural shocks. As a solution, in the next section, we propose a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method (discussed in Section 4). If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
3.5 Labelling the shocks
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroskedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017). Additionally, in line with the statistical identification literature, any economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, because the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, due to the problem of weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, discussed in Sections 3.3 and 3.4, the ordering and signs of the columns of $B_{2},...,B_{M}$ is often practically arbitrary even when the ordering and signs are, in theory, uniquely identified by Proposition 1 or 2.
Assuming that the parameters are estimated by the method of maximum likelihood (ML), weak identification with respect to the ordering and signs of the columns often results in multiple local maxima of the log-likelihood function with log-likelihoods very close to each other. Our estimation method, discussed in Section 4 and implemented to the R package sstvars (Virolainen, 2024), can produce a set of local solutions to the estimation problem. This allows the researcher to check whether there are multiple local solutions with log-likelihoods very close to each other. In practice, one can then simply choose from the almost equally well fitting local solutions the one that, based on economic reasoning, facilitates labelling the shock(s) of interest to the same column of the impact matrix in all regimes.
To exemplify, consider the simplistic setup of two variables, an output variable and a price variable, two shocks, the aggregate supply shock and the aggregate demand shock, and two regimes ($M=2$). Suppose the estimates of the impact matrices of the two regimes with the highest log-likelihood are the following:
$\hat{B}_{1}=\begin{bmatrix}0.67&\phantom{-}0.21\\ 0.06&-0.21\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}-0.55&1.14\\ \phantom{-}0.49&0.09\\ \end{bmatrix}.$ (3.5)
Then, suppose there exists another local solution solution with the second highest likelihood that is almost equal to the highest log-likelihood, but the estimates of the impact matrices are the following:
$\hat{B}_{1}=\begin{bmatrix}0.59&\phantom{-}0.16\\ 0.05&-0.17\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}1.05&\phantom{-}0.58\\ 0.18&-0.41\\ \end{bmatrix}.$ (3.6)
In the estimates (3.5), the first shock moves output and prices in the same direction at impact in Regime 1 but in opposite directions in Regime 2. The second shock in, turn, moves output and prices in opposite directions in Regime 1 but in the same direction in Regime 2. Thus, neither of the shocks can be labelled as the aggregate supply shock nor as the aggregate demand shock. On the other hand, in the estimates (3.6), the first shock moves output and prices in the same direction satisfies the same signs of the impact effects in both regimes, so it can be labelled as the aggregate supply shock. The second shock moves output and prices in opposite directions and satisfies the same signs of the impact effects in both regimes, so it can be labelled as the supply shock.
Since the estimates in (3.5) do not align with what is expected from the aggregate supply and demand shocks but the estimates in (3.6) do, the latter solution can be chosen based on economic reasoning despite having marginally smaller log-likelihood. [^1] It is also possible to obtain the estimates in (3.6) more formally by using restricted ML estimation, imposing the restriction that the first shock moves both variables in the same direction and the second shock in opposite directions on impact. Since the solution (3.6) yields the second highest log-likelihood among all local solutions, it is obtained from the restricted ML estimation. Note that also other types of information than just signs of the impact responses of the variable can be utilized to label the shocks. For instance, if the shock of interest should or should not significantly move certain variables on impact, the corresponding estimates of $B_{1},...,B_{M}$ should satisfy this property. Lanne et al. (2017) go further and exploit more information by labelling the shocks based on the estimated impulse response functions.
If none of the alternative local solutions with almost equal log-likelihoods allows the shock of interest to be labelled to same column across $B_{1},...,B_{M}$, the identification can be strengthened by imposing economic overidentifying restrictions on the model. For example, if economic reasoning suggests that certain variables should respond to the shock of interest with a lag, the appropriate zero restrictions can be imposed on the corresponding entries of $B_{y,t}$. Due to the statistical identification, these zero restrictions are overidentifying and thereby testable, which a major advantage compared to using conventional short-run restrictions alone.
4 Estimation
The parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of maximum likelihood. Obtaining the log-likelihood function requires specifying a distribution for the structural errors $e_{t}=(e_{1t},...,e_{dt})$ satisfying Assumption 1. In our empirical application presented in Section 5, we assume that each shock $e_{it}$, $i=1,...,d$, follows a Student’s $t$ distribution with mean zero, variance one, and degrees of freedom $\nu_{i}\in(2,\infty)$. The assumption $\nu_{i}>2$ is made to ensure the existence of the second moments, and the assumption $\nu_{i}<\infty$ is made to distinguish the distribution from the Gaussian distribution.
In order to normalize the variances of the structural shocks to unity, we parametrize the Student’s $t$ distribution using its variance as a parameter together with the mean and degrees-of-freedom. Similar parametrization has been utilized previously in Virolainen (2022) and Meitz et al. (2023), for example. The density function of such a univariate $t$-distribution with mean $\mu_{i}$, variance $\sigma_{i}^{2}>0$, and $\nu_{i}>2$ degrees-of-freedom is
$t_{1}(y_{it};\mu_{i},\sigma_{i}^{2},\nu_{i})=\frac{\Gamma\left(\frac{1+\nu_{i}}{2}\right)}{\sqrt{\pi(\nu_{i}-2)}\Gamma\left(\frac{\nu_{i}}{2}\right)}\sigma_{i}^{-1}\left(1+\frac{(y_{it}-\mu_{i})^{2}}{(\nu_{i}-2)\sigma_{i}^{2}}\right)^{-(1+\nu_{i})/2},$ (4.1)
where $\Gamma\left(\cdot\right)$ is the gamma function.
We collect the parameters to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),....,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(vec(B_{1}),...,vec(B_{M}))$, $\alpha$ contains the transition weight parameters (and is omitted if exogenous weights are used), and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters. Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (4.2)
where the impact matrix $B_{y,t}$ is defined in Equation (3.3), $I_{d,i}$ is the $i$th column of the $(d\times d)$ identity matrix $I_{d}$, and the Student’s $t$ density $t_{1}(\cdot;0,1,\nu_{i})$ is defined in (4.1). The ML estimator maximizes the log-likelihood function (4.2) with respect to the parameter $\boldsymbol{\theta}$.
To facilitate obtaining desirable asymptotic properties for the ML estimator, such as consistency and a conventional limiting distribution, it is useful to establish the ergodicity and stationarity of the structural STVAR model. Based on the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one. However, since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in a numerical ML estimation. Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition. After the estimation, the sufficient condition can be checked for the solutions of interest.
In practice, maximizing the log-likelihood function typically requires numerical optimization techniques. Particularly if the transition weights are endogenous, there are often a very large number of modes in the surface of the log-likelihood function, and large areas in the parameter space where it is flat in multiple directions. To address these complexities, we follow Meitz et al. (2023) and Virolainen (2022, forthcoming) and adopt a two-phase estimation procedure, which is ran a large number of times. In the first phase, a genetic algorithm is employed to find starting values for a gradient based variable metric algorithm. In the second phase, the variable metric algorithm is then initialized from each starting value found in the first phase, thereby often converging to a nearby local maximum. Some of the estimation rounds may end up in saddle points or points near the boundary of the parameters space (e.g., in points where $B_{y,t}$ is nearly singular for some $t$, or Condition A.1 in Appendix A, if imposed, is close to breaking), and some of the local solutions might be inappropriate for statistical inference (e.g., there might be only few observations from some of the regimes).
As the log-likelihood function is highly multimodal and the estimation algorithm is ran a large number of times, it often results in a variety of local solutions. After the estimation rounds have been ran, the solution that maximizes the log-likelihood function among the appropriate local solutions can be selected. Then, if the transition weights are endogenous and a stationary solution is preferred, the sufficient condition for ergodic stationarity of the model (Assumption A.1(I) in Appendix A) can be checked for the local solutions of interest to select the best solution whose stationarity can be verified. If there are multiple such solutions with almost equal log-likelihoods, the best local solution that facilitates labelling the shocks of interest can selected based on the procedure discussed in Section 3.5. The accompanying CRAN distributed R package sstvars (Virolainen, 2024) implements the two-phase estimation procedure utilizing a modified genetic algorithm.
5 Empirical application
Our empirical application studies the macroeconomic effects of the climate policy uncertainty shock. Currently, there exist only a few studies on the subject. Fried et al. (2021) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy risk decreases total capital stock, causing output and carbon emissions to decrease. Gavriilidis et al. (2023), in turn, propose a news based index on measuring climate policy uncertainty (CPU), and the study the effects of the CPU shock using a linear recursively identified SVAR model and local projections. They find that a positive shock to their CPU index (CPUI) decreases real activity and increases prices, thus, transmitting to the U.S. economy as a supply shock.
We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty. Such variation can be accommodated by our structural STVAR model by specifying a logistic transition weight function with the lagged level of economic policy uncertainty as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more nuanced results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
5.1 Data
We consider a monthly U.S. dataset consisting of six variables and covering the time period from 1987:4 to 2024:2. For measuring climate policy uncertainty, we make use of the CPUI data (Gavriilidis, 2021), whose availability determines the beginning of our sample period. The CPUI is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty.
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking the first differences. For measuring the price level, we use the consumer price index (CPI) and the producer price index (PPI), which are detrended by taking the first differences. As a forward looking variable, the producer price index should also help to distinguish the effects of climate policy uncertainty shocks from climate policy actions (see Gavriilidis et al., 2023). Finally, as a measure of the monetary policy stance, we include an interest rate variable (RATE), which is the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods. The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, and the Federal funds rate are retrieved from the Federal reserve bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:2. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$), economic policy uncertainty index (divided by $10$), the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$), and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the estimated transition weights of the first regime of the fitted two-regime fifth-order LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
The series of the included variables are presented in Figure 1. Both CPUI and EPUI show a generally increased level of uncertainty towards the end of the sample period, suggesting that a model accommodating shifts in the mean could be appropriate. Moreover, there is clearly variation in the volatility of the variables, so shifts in the volatility regime should also be incorporated into the model.
5.2 Logistic STVAR model identified by non-Gaussianity
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks need to be specified. As transition weights, we employ the popular logistic transition weights (Anderson and Vahid, 1998) given in Equation (2.4) for a two-regime model. They facilitate smooth transitions between the regimes depending the level of the switching variable, and they also nest the TVAR model (Tsay, 1998) as a special case when the scale parameter $\gamma$ tends to infinity. To study how the effects of the CPU shock vary depending on the level of economic policy uncertainty, we specify the first lag of EPUI as the switching variable. This approach accommodates shifts in the mean and volatility depending on the level of EPUI, and hence allows the effects of the CPU shock to vary with the level of economic policy uncertainty.
We assume that the structural shocks follow independent Student’s $t$ distributions, which is a fairly standard choice for modelling fat tailed data. Moreover, as the Gaussian distribution is obtained as a special case of the $t$-distribution, with the value of the degrees-of-freedom parameter tending to infinity, the validity of the identifying Assumption 1 can be assessed based on the estimates of the degrees-of-freedom parameters. In order to maintain a reasonable number of parameters in the model and to avoid the problems of potentially imprecise estimates and overfitting, we assume that the autoregression matrices are constant across the regimes, i.e., employ the specification in Equation (2.6). Due to the time-varying intercepts and impact matrices, the model is still fairly general, allowing for shifts in the mean, volatility, and impact responses of the variables to the shocks.
Our structural two-regime logistic STVAR (LSTVAR) model can be written as
[…]
where $c$ is the location parameter, $\gamma>0$ is the scale parameter, $y_{2t-1}$ is the lagged EPUI, the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ follow independent $t$-distributions with zero mean and unit variance, and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters of the structural shocks. Finally, we select the autoregressive order $p=5$ for our model based on AIC.
The estimates of the transition weights of Regime 1 are presented in the bottom panel of Figure 1. They show that Regime 1 dominates in the periods of low economic policy uncertainty, whereas Regime 2 conversely dominates in the periods of high economic policy uncertainty. The location parameter estimate is $\hat{c}=9.80$ and the scale parameter estimate is $\hat{\gamma}=1.00$. Since the sample mean of EPUI over our sample period is $11.48$, Regime 2 starts dominating roughly when EPUI surpasses its mean, and based on the small estimate of the scale parameter, the transitions between the regimes are smooth. Therefore, we label Regime 1 as the low economic policy uncertainty (EPU) regime and Regime 2 as the high EPU regime.
The estimates of the degrees of freedom parameters are small, $\hat{\nu}=(3.86,5.54,2.94,2.61,2.02,2.23)$, and therefore none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the matrices $\boldsymbol{A}_{m}$ (A.2), $m=1,2$, using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Hence, the ML estimator is expected to be well-behaved.
To assess the adequacy of our model, we utilize residual based model diagnostics to examine how well our model captures the autocorrelation structure, conditional heteroskedasticity, and distribution of the data. To that end, we plot several diagnostic figures, which are presented and discussed in more detail in Appendix B.6. They show that our model captures the autocorrelation structure of the data reasonably well, but there is a large correlation coefficient in the crosscorrelation functions between two of the residual series. Also conditional heteroskedasticity is somewhat reasonably well captured, although there are a few relatively large correlation coefficients in the auto- and crosscorrelation functions of squared standardized residuals. The marginal distributions of the series are also reasonably well captured, but the distributions of two of the residual series are somewhat asymmetric. Nonetheless, in our view, the overall adequacy of the model is reasonable enough for impulse responses analysis.
5.3 Labelling the CPU shock
The shocks of our structural LSTVAR model (5.1) are readily identified by Proposition 2. However, labelling the shocks requires external information. Moreover, due to the problem of weak identification discussed in Section 3.5, we should choose among the practically equally well fitting local solutions the one that facilitates labelling the CPU shock to the same column of the impact matrix in all of the regimes.
The estimates of $B_{1}$ and $B_{2}$ for the local solution (discussed in Section 5.2) with the highest obtained log-likelihood are the following:
[…]
where the ordering of the variables is $y_{t}=(\text{CPUI}_{t},\text{EPUI}_{t},\text{IPI}_{t},\text{CPI}_{t},\text{PPI}_{t},\text{RATE}_{t})$, and the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are fixed by normalizing the entries in the first row of $\hat{B_{1}}$ to be positive and in a decreasing order.
In terms of labelling the CPU shock based on the impact responses of the variables in the estimates of $B_{1}$ and $B_{2}$, there are several properties that are reasonable to expect from it. First, since the CPU shock is the shock to CPUI, the impact response of CPUI should be substantial, and it should have the same sign in $B_{1}$ and $B_{2}$. Second, as CPU contributes to the overall economic policy uncertainty, EPUI should increase or at least not significantly decrease in response to a positive CPU shock. The effects of the CPU shock on IPI are more ambiguous. Increased uncertainty may postpone investments, decreasing production, while some firms may accelerate production to avoid future costs or invest in future technologies to gain a competitive edge. The CPU shock’s impact on the price level, in turn, may be indirect, translated through the effects on the aggregate demand, production costs, or production levels. Similarly, the Fed may not directly respond to increased uncertainty but rather to its economic impacts.
Based on the estimates of $B_{1}$ in Equation (5.2), the only variable that moves substantially at impact in response to the first shock is CPUI, and the first shock has also clearly the highest impact effect on CPUI. The first shock also moves EPUI in the same direction with CPUI, in line with increased CPU contributing to the level of general economic policy uncertainty. Also the second shock moves CPUI considerably on impact, but it moves EPUI much more, so since CPUI and EPUI vary in the same scale, it does not appear to be a plausible CPU shock candidate. Therefore, the first shock seems to be the most reasonable candidate as the CPU shock. The estimates of $B_{2}$ in Equation (5.2) show that the first shock substantially moves CPUI at impact, while the response of CPUI to the other shocks is clearly smaller. Importantly, the sign of the response of CPUI to the first shock is the same as in $\hat{B}_{1}$. Moreover, the first shock moves EPUI in the same direction with CPUI, thus, in line with what can be expected from a CPU shock. Overall, the first shock seems to be the most plausible CPU shock candidate, and hence, we deem it as the CPU shock. Since we labelled the CPUI shock in the local solution that obtained the highest likelihood, there is no need to consider other local solutions incorporating practically equal fit and mostly very similar estimates but different ordering and signs of the columns of $B_{2}$ (see Section 3.5).
5.4 Generalized impulse response functions
The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, which makes the conventional way of calculating impulse responses unsuitable. Therefore, we consider the general impulse response function (GIRF) (Koop et al., 1996) that accommodates such features and is defined as
$\text{GIRF}(h,\delta_{i},\mathcal{F}_{t-1})=\text{E}[y_{t+h}|\delta_{i},\mathcal{F}_{t-1}]-\text{E}[y_{t+h}|\mathcal{F}_{t-1}],$ (5.3)
where $h$ is the horizon. The first term on the right side of (5.3) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the structural shock of sign and size $\delta_{i}$ in the $i$th element of $e_{t}$ hits the system at time $t$ as opposed to all shocks being random. The generalized impulse response functions also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.3).
As our structural LSTVAR model has a $p$-step Markov property, conditioning on (the $\sigma$-algebra generated by) the $p$ previous observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is effectively the same as conditioning on $\mathcal{F}_{t-1}$ at time $t$ and later. In order to study state dependence of the effects of the CPU shock, we obtain the histories $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ from low EPU and high EPU regimes separately. Specifically, for each regime, we take all of the length $p$ histories in the data that imply transition weights larger than $0.75$ for the given regime, so that the given regime is clearly dominant, and the number of histories from both of the regimes is reasonably large. For low EPU regime, there are $112$ such histories and for high EPU regime there are $202$. Then, for each regime separately, we calculate the GIRFs for these length $p$ histories, and take the sample mean to obtain a point estimate for the GIRF conditional on dominance of the given regime, whereas taking sample quantiles gives confidence intervals that reflect uncertainty about the initial state of the economy within the dominating regime. This way the GIRFs estimate the effects of the CPU shock in times of low EPU and high EPU in a way that reflects the observed data. Our Monte Carlo algorithm for computing the GIRF is similar to the one described in Lanne and Virolainen (2024), Appendix B.
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of Regime 1 are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to a positive one-standard-error CPU shock with initial values from Regime 1, whereas the right column shows the responses with initial values from Regime 2. All GIRFs have been scaled so that the instantaneous increase in CPUI is $5$. The shaded areas depict the $95\%$ confidence intervals that reflect uncertainty about the initial state of the economy within the regime.
The GIRFs to a one-standard-error positive CPU shock are presented in Figure 2. We also computed the GIRFs for two-standard-error and negative CPU shocks, but the responses are similar to the positive one-standard-error shock (after adjusting for the different sign or size), so they are not presented for brevity. The responses of CPUI and EPUI follow similar patterns during times of low and high EPU, with CPUI responding substantially and EPU more weakly. The responses of the macroeconomic variables, however, differ across the regimes.
When the level of EPU is low, production decreases following a positive CPU shock, possibly because of firms postponing investments (see Fried et al., 2021). On the other hand, when EPU is high, production slightly increases. A possible explanation is that some firms, already operating in a more unpredictable environment, continue with planned investments, or hedge against multiple types of uncertainties by ramping up production in sectors less likely to be immediately affected or by investing in a broader range of technologies.
Consumer prices initially decrease and then slightly increase when EPU is low, while they increase more significantly when EPU is high. The initial decrease in prices under low EPU could be due to reduced consumer demand as people postpone consumption in anticipation of new policies. Under high EPU, prices rise more substantially, possibly because of firms adapting their operations to meet potential new regulations, increasing production costs and passing these costs onto consumers.
Producer prices rise permanently in both regimes, possibly due to anticipatory adjustments to future costs. The effects of the CPU shock on the interest rate are minor but consistent with the responses of the other variables, with interest rates increasing as prices rise, more so during high EPU. Finally, the responses of the transition weights indicate a shift towards the high EPU regime, which is expected, as a positive CPU shock increases EPUI.
Overall, our results suggest that a positive CPU shock decreases production in times of low EPU but slightly increases it in times of high EPU. Prices and the interest rate increase in both regimes, but the increase in consumer prices (and interest rate) is more significant when level of EPU is high. Therefore, our results are complementary to Gavriilidis et al. (2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.
6 Conclusions
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations. Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
The identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes, and an incorrect ordering or signs of the columns results in impulse response functions that do not estimate the effects of the intended shock but of a linear combination of the structural shocks. In other words, there is often multiple local maxima of the log-likelihood function with almost equal log-likelihood that are very similar to each other but incorporate different ordering and signs of the columns of the impact matrices in some regimes. As a solution, we proposed a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method. If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
In addition to discussing the identification of the shocks, we discussed the estimation of the model parameters by the method maximum likelihood in case of independent Student’s $t$ shocks, and proposed a two-phase estimation procedure that can produce a set of alternative local solutions. By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model. While useful for obtaining a well-behaving estimator, stationarity is not, however, required for applicability of our identification results. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S. data from 1987:4 to 2024:2. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable and allowing for smooth transitions in the intercept parameters as well as in the impact matrix. We found that a positive CPU shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty. Consumer prices, in turn, increase more significantly when the level of economic policy uncertainty is high than when it is low. Our results are, therefore, complementary to Gavriilidis et al. (2023), who found a positive CPU decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.
Appendix A Stationarity and ergodicity
Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties for the estimator, such as consistency and a conventional limiting distribution. Saikkonen (2008) derives a sufficient condition for ergodicity and stationarity of a smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model cannot be directly obtained from the model of Saikkonen (2008). The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (although the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights). Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
The sufficient stationarity condition is expressed in terms of the joint spectral radius (JSR) of certain matrices. The JSR of a finite set of square matrices $\mathcal{A}$ is defined by
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (A.1)
where $\mathcal{A}^{j}=\{A_{1}A_{2}...A_{j}:A_{i}\in\mathcal{A}\}$ and $\rho(A)$ is the spectral radius of the square matrix $A$.
Consider the companion form AR matrices of the regimes defined as
$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (A.2)
The following assumption collects sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
Assumption A.1.
Suppose the following conditions hold:
  1. (I)
    $\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})<1$,
  2. (II)
    The distribution of the $IID(0,I_{d},\nu)$ random vector $e_{t}$ has a (Lebesgue) density that is bounded away from zero on compact subsets of $\mathbb{R}^{d}$.
  3. (III)
    The transition weight $\alpha_{m,t}$ is an $\mathcal{F}_{t-1}$-measurable function of $\{y_{t-j},j=1,...,p\}$ for all $t$ and $m=1,...,M$.
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008). Note that this condition is sufficient but not (necessarily) necessary for ergodic stationarity of the process. Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied. This condition rules out nonrandom exogenous weights, which are perfectly fine in terms of our main identification results discussed in Section 3.
The following theorem, which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
Theorem A.1.
Consider the STVAR process $y_{t}$ defined in Equations (3.2) and (3.3), and suppose Assumptions 1 and A.1 are satisfied. Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the process $\boldsymbol{y}_{t}=(y_{t},...,y_{t-p+1})$ is a $(1+||x||^{2})$-geometrically ergodic Markov chain. Thus, there exists a choice of initial values $y_{-p+1},...,y_{0}$ such that the process $y_{t}$ is strictly stationary, second-order stationary, and $\beta$-mixing with geometrically decaying mixing numbers.
Theorem A.1 is proven in Appendix A.1. A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere. To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a zero Lebesgue measure.
Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical maximum likelihood estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large of times, thus, substantially increasing the computational cost of the estimation, making it extremely tedious. Therefore, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
Condition A.1.
$\max\{\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M})\}<1$,
where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, $m=1,...,M$.
Condition A.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). Since our estimation method discussed in Section 4 can produce a number of local solutions, the researcher can choose the best appropriate local solution whose stationarity can be verified by Assumption A.1(I).
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2024) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
Appendix B Proofs
B.1 Proof of Lemma 1
We start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M})\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero.
It then follows from Implicit function theorem that in the points $B^{0}=(vec(B_{1,0}),...,vec(B_{M,0}))\in\mathbb{R}^{Md^{2}}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold locally in a neighbourhood of $B^{0}$. Because this argument applies at all points $B^{0}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, it follows that the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold in regions where this condition holds. Therefore, the Lebesque measure of the set $S_{t}$ is zero in $\mathbb{R}^{Md^{2}}$ in the region where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P_{t}(B_{1},...,B_{M})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$ for each $t$. Hence, the union $\cup_{t=1}^{\infty}S_{t}$ is a union of a countable number of measure zero sets and thus itself has a Lebesgue measure zero. It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
B.2 Proof of Lemma 2
For most parts, our proof is similar to the proof of Proposition 1 in Lanne et al. (2017) making use of results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
$y_{t}-\phi_{t}-\boldsymbol{A}_{t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.1)
where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.1) holds also for some other coefficients, impact matrix, and shocks $\tilde{\phi}_{t},\tilde{\boldsymbol{A}}_{t}$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$, i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.1), we obtain
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}$ $=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}$ (B.2)
Taking conditional expectation conditional on $\mathcal{F}_{t-1}$ gives
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}$ $=0,$ (B.3)
as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (B.2), we get
$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.4)
Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most on column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$) so that $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
Now, from the invertibility of $C$ it follows that each column of $C$ as exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element. Therefore, there exists a permutation matrix $P$ and a diagonal matrix $D=\text{diag}(d_{1},...,d_{d})$ with nonzero diagonal elements such that $C=DP$. Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.
To complete proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.∎
B.3 Proof of Proposition 1
We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ would lead to model that is not observationally equivalent with the original one. Specifically, we say that two models are observational equivalent if and only if their joint densities are equal. After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
Condition (a) readily fixes the signs of the columns of $B_{1},...,B_{M}$, so it suffices the check that under Conditions (c) and (b), changing the ordering of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{t})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (B.5)
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $d_{e_{i}}(\cdot;0,1,\nu_{i})$ is the density function the $i$th shock. At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$, implying the that the density (B.5) can be expressed as
$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}),\ \ \text{and}$ (B.6)
$f_{y}(y_{t}-\mu_{t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i}),$ (B.7)
where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices that satisfy $\alpha_{m,t}=1$ for the regime $m$, and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but its dependence on the other parameters is omitted).
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $B_{m}(\boldsymbol{\tau}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$. By the invertibility of $B_{m}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{t})$ are equal to each other almost everywhere in $Y_{T}\in\mathbb{R}^{T}$. Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it then follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m})(y_{t}-\mu_{t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m})(y_{t}-\mu_{t})$ are in a different order. Moreover, as $|\det(B_{m})|$ is invariant to the ordering of the columns of $B_{m}$, it follows from this and Condition (c) that, almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, $f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m}))$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$.
Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.8)
and thus also
$f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\boldsymbol{\tau}_{2}),...,B_{M}(\boldsymbol{\tau}_{M}))\neq f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\tilde{\boldsymbol{\tau}}_{2})...,B_{M}(\tilde{\boldsymbol{\tau}}_{M}))$ (B.9)
almost everywhere in $Y_{T}\in\mathbb{R}^{T}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ for any $m=2,...,M$. Since the this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.
Finally, by Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the signs of the columns of $B_{1},...,B_{M}$ and the ordering of the columns $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.∎
B.4 Proof of Proposition 2
To proof consists of two parts. First, we show that almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, if the ordering and signs of the columns of $B_{1}$ is fixed, changing ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the invertible impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$. We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.10)
for some $(d\times d)$ diagonal matrix $D_{t}$ with ±$\pm 1$ diagonal elements and some $(d\times d)$ permutation matrix $P_{t}$ (that may vary in time). Rearranging the above equation gives
$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})D_{t}P_{t})=-\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t}).$ (B.11)
Suppose first $D_{t}P_{t}=I_{d}$. Equation (B.12) then simplifies to
$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))=0.$ (B.12)
By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, Equation (B.12) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other. Hence, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.12) forms a subspace in $\mathbb{R}^{d\times d(M-1)}$ with the dimension of this subspace being less than $(M-1)d^{2}$, as it must satisfy a nontrivial linear equation. Consequently, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.12) has a Lebesgue measure zero in $[B_{2}:\cdots:B_{M}]\in\mathbb{R}^{d\times d(M-1)}$, implying that the set of matrices $B_{1},...,B_{M}$ satisfying this equation have a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.11) as $C_{t}\in\mathbb{R}^{d\times d}$. By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, the right side of Equation (B.11) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros. Equation (B.11) is thereby nontrivial and can be expressed as
$B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t})=\alpha_{1,t}^{-1}C_{t}$ (B.13)
The set of matrices $B_{1}$ that satisfy Equation (B.13) forms a subspace in $\mathbb{R}^{d\times d}$ with the dimension of this subspace being less than $d^{2}$, as it must satisfy a nontrivial linear equation. It follows that the set of matrices $B_{1}$ that satisfy Equation (B.13) has Lebesgue measure zero on $\mathbb{R}^{d\times d}$, and therefore, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.13) has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$.
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.11) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.12) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$. By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$). Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A)), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$. Hence, reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ would lead to a model that is not observationally equivalent with the original one. The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A), implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$). Thus, $B_{y,t}$ is uniquely identified at all $t$.
Since $B_{y,t}$ is uniquely identified and two models are observationally equivalent if and only if they are observationally equivalent at all $t$, a unique identification of the matrices $B_{1},...,B_{M}$ is achieved if an impact matrix that is identical to $B_{y,t}$ for all $t$ cannot be obtained with any other invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$. Suppose an impact matrix identical to $B_{y,t}$ for all $t$ is obtained with some invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$. Then,
$\sum_{m=1}^{M}\alpha_{m,t}(B_{m}-\tilde{B}_{m})=0,$ (B.14)
for all $t$.
Denote the $ij$th element of $\tilde{B}_{m}$ as $\tilde{b}_{ij}^{(m)}$. At each $t$, Equation (B.14) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$. The weights $\alpha_{m,t}$ are treated as constant multipliers here, as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here. If $\alpha_{m,t}=0$ for some regimes $m$, the corresponding variables in $\tilde{B}_{m}$ drop out of the system of equations, however.
Under Condition (C), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values (plus one accounting for transition weights that equal to zero). Therefore, Equation (B.14) defined together for all $t$ defines a system of linear equations involving $Md^{2}$ variables and at least $Md^{2}$ linearly independent equations. Moreover, as $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values, each variable $\tilde{b}_{ij}^{(m)}$ appears in at least $M$ linearly independent the equations. Thus, there exists at most one solution to this system of linear equations. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.∎
B.5 Proof of Theorem A.1
Saikkonen (2008) establishes stationarity, ergodicity, and mixing properties of a STVEC model by making use of the theory of Markov chains. Kheifets and Saikkonen (2020), in turn, show that the results of Saikkonen (2008) are applicable to their STVAR model, as it can be obtained as a special case of the STVEC model of Saikkonen (2008). However, the results of Saikkonen (2008) and Kheifets and Saikkonen (2020) are not directly applicable to our structural STVAR model, because our model defines the conditional covariance matrix differently to the STVEC model of Saikkonen (2008) and STVAR model of Kheifets and Saikkonen (2020). Hence, certain steps need to be taken show the applicability of the results of Saikkonen (2008) to our structural STVAR model.
Note that by Assumption A.1(II), Assumption 1 in Saikkonen (2008) is satisfied. Moreover, by the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$, Assumption 2(a) in Saikkonen (2008) satisfied, and as we assume that the set $\text{l}_{2}$ in Saikkonen (2008) is an empty set, also Assumption 2(b) is satisfied. Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption A.1(I) also Condition (19). The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows. First, we assume that the number of unit roots in the model is zero, so $n=r$ in Saikkonen (2008). Consequently, $z_{t}=y_{t}$ and, in the definition of the matrix $J$ defined above Equation (17) of Saikkonen (2008), $\beta=c=I_{d}$ (see Saikkonen, 2008, Note 2). The matrix $J$ is thereby nonsingular, and following Kheifets and Saikkonen (2020), we can choose the matrix $S$ in Equation (17) to be the inverse of $J^{\prime}$. Second, we assume the transition functions $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})$ of Saikkonen (2008) are independent of the random variable $\eta_{t}$. Hence, the transition functions are functions of $Z_{t-1}=(y_{t-1},...,y_{t-p})$, and we can write $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})=\alpha_{m,t}$. Using the notation of this paper, the first term on the right side of Equation (17) in Saikkonen (2008) then reduces to $\sum_{m=1}^{M}\alpha_{m,t}\sum_{i=1}^{p}A_{m,i}y_{t-i}$. Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ (with the notation of this paper).
That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), reduce to the first term on the right hand side of Equation (3.2) defining our structural STVAR model. The third term on the right side of Equation (17) in Saikkonen (2008), however, reduces to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes. This is different to the second term of Equation (3.2), which takes the form $\sum_{m=1}^{M}\alpha_{m,t}B_{m}e_{t}$ when the impact matrix (3.3) is substituted to it. Nonetheless, it turns out that with only slight modifications, the proof of Theorem 1 in Saikkonen (2008) applies to our model as well.
The conditional covariance matrix is relevant in only a few places in the proof of Theorem 1 in Saikkonen (2008). Firstly, on page 312, Saikkonen (2008) makes use of the positive definiteness of the conditional covariance matrix $H(S^{\prime}JZ_{t-1})$ (or $\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ in our notation) to conclude that the smallest eigenvalue of $H(S^{\prime}Jx)$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. Saikkonen (2008) argues that thereby by Theorem 2.2(ii) of Cline and Phu (1998), the Markov chain $Z_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, which is then used to conclude further results. Concerning our model, note that by Lemma 1, $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, implying that the conditional covariance matrix $B_{y,t}(\boldsymbol{y}_{t-1})B_{y,t}(\boldsymbol{y}_{t-1})^{\prime}$ is positive definite, where the dependence of the conditional covariance matrix on the $(dp\times 1)$ random vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is made explicit in the notation. Hence, the smallest eigenvalue of $B_{y,t}(x)B_{y,t}(x)^{\prime}$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. It follows from this, Equation (3.2), and the invertibility of $B_{y,t}$ that we can apply Theorem 2.2(ii) of Cline and Phu (1998) to conclude that the Markov chain $\boldsymbol{y}_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, similarly to the process $Z_{t-1}$ in Saikkonen (2008).
Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p. 358) is satisfied. To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$, where $j$ is a positive integer, is bounded from above by a finite real number.
Similarly to Saikkonen (2008, p. 313), clearly,
$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (B.15)
where |||||$\cdot$||| | ⋅ | | is Eucledian norm, $N$ is a positive integer, and $c$ a finite real number. Hence, because $B_{y,t}(x)$ is a constant here, by Jensen’s inequality and submultiplicativity of norms, we have
$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E[||e_{t}||]^{2},$ (B.16)
where the right side is bounded from above by some finite constant (by (B.15) and the finite second moment of $e_{t}$ given by Assumption 1). Thus, an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model. Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
B.6 Details on the empirical application
The ML estimation of the models, residual diagnostics, and estimation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2024) in which the introduced methods have been implemented. The R package sstvars also contains the dataset studied in the empirical application to facilitate reproduction of our results. In the ML estimation, we consider only solutions that can be deemed local maximums of the log-likelihood function.
In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because identification with respect to the ordering of the columns of $B_{2}$ is weak (see Section 3.5). One the other hand, changing the ordering of the columns of $B_{y,t}$ changes the ordering of the variables in the density function $t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})$, thus, changing the ordering of the residuals in the given regime. The residual time series can, therefore, be labelled only for the variables whose structural shocks have been labelled. Since we labelled the CPU shock as the first shock in Section 5.3, the first residual series constitutes the residuals of CPUI. Since the other structural shocks were not labelled, the other residual series cannot be labelled for any of the variables.
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the residuals for the first $24$ lags in Figure B.1. As the figure shows, there is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the first lag of the CCFs between the fourth and fifth residual series. There are also some moderate sized CCFs in the ACFs of the third residual series, but these are not particularly large. Since rest of the CCs are reasonable, our LSTVAR model seems to capture the autocorrelation structure of the data reasonably well.
In order to study the model’s adequacy to capture the conditional heteroskedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure B.2. The figure shows that there are several moderate and relative large CCs in the ACFs and CCFs of some of the residual series. Thus, there is some but not particularly severe inadequacies in terms of capturing the conditional heteroskedasticity.
The series of standardized residuals, presented in the top panels of Figure B.3, also show some remaining heteroskedasticity and several outliers, many of which are accommodated by the heavy tailed $t$-distributions. Finally, the quantile-quantile-plots based on the independent $t$-distributions with mean zero, variance one, and degrees-of-freedom given by the degrees-of-freedom parameter estimate are presented in the bottom panel of Figure B.3. They show that the marginal distributions of the series are relatively well captured, although the distributions of two of the residual series are slightly asymmetric. Nonetheless, in our view, the overall adequacy of the model seems reasonable enough.
Figure B.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
Figure B.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
Figure B.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime fifth-order LSTVAR model. The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees of freedom given by the ML estimate of the degrees of freedom parameter.

### v3

@@
 We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes.
-We also discuss labelling of the shocks, maximum likelihood estimation of the parameters, and stationarity the model.
+We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model.
 The introduced methods are implemented to the accompanying R package sstvars.
@@
 macroeconomy.
-In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty.
-Keywords: smooth transition vector autoregression, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
+In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
+Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
 The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P.
@@
 To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
-There are two main branches in the statistical identification literature: identification by heteroskedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, forthcoming, and others) and identification by non-Gaussianity (Lanne et al., 2017, Maxand, 2020, Lanne and Luoto, 2021, Lanne et al., 2023, Anttonen et al., forthcoming, and others).
+There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others).
 Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions.
-However, identification by heteroskedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see the discussion in Virolainen, forthcoming, Sections 3.1 and 4).
+However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017).
 This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
-Bacchiocchi and Fanelli (2015) do propose an alternative framework for identification by heteroskedasticity that allows for variation in the impact effects of the shocks across the volatility regimes, but their method requires imposing a number of untestable zero impact effect restrictions that can be difficult to justify economically.
 To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
@@
 In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification.
-Similarly to Lanne et al.
-(2017), however, it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian.
+Nonetheless, similarly to Lanne et al.
+(2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian.
 We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs.
-The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations.
-In line with statistical identification literature, labelling the identified structural shocks as economic shocks requires external information.
-This requires labelling the same shock to the same column of the impact matrix across the regimes.
-In our experience, however, the identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes.
-This can be problematic because if the shocks are ordered differently in some regimes, the (generalized) impulse response functions do not estimate the effects of the intended shock but of a linear combination of structural shocks.
-As a solution, we propose a procedure that utilizes the almost equally well fitting local solutions with different orderings and signs of the columns of the impact matrices obtainable by our estimation method.
-If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
+The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well.
+In line with statistical identification literature, however, external information is required to labelling the identified structural shocks as economic shocks.
+Since the identification can, in our experience, be also weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional information (cf.
+Carriero et al., 2024).
 Developments related to ours include the new framework of Morioka et al.
 (2021) called independent innovation analysis (IIA).
-Morioka et al.
-(2021) propose using IIA for directly estimating the (conditionally) contemporaneously independent innovations of a general nonlinear vector autoregressive (VAR) model without assuming a specific functional form for the model.
-They show that under a certain set of (fairly general) assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations.
-In applied macroeconometrics, the main interest is, however, typically in computing the (generalized) impulse response functions or performing other analyses that require obtaining an estimate for the parameters of the model.
-Estimating the parameters of a STVAR model using the framework of Morioka et al.
-(2021) requires first estimating the independent innovations using a specific machine learning algorithm (see Morioka et al., 2021), and then using the estimates of the innovations to obtain estimates of the parameters.
-However, it is not obvious how to estimate the parameters of the model from the innovations obtained by IIA, as they are identified only up to permutation and component-wise invertible transformations, and different parameter estimates would be obtained with different transformations.
+The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form.
+Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations.
+While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis.
 Moreover, Morioka et al.
-(2021) assume that the innovations are from a distribution in an exponential family, thereby excluding various distributions such as the popular Student’s $t$ distribution.
-Our approach, in turn, facilitates estimating the parameters directly by the method of maximum likelihood, and it does not exclude distributions of the independent shocks other than Gaussian distribution and those with infinite variance.
+(2021) assume innovations follow a distribution in the exponential family, thereby excluding various alternatives.
+In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
 Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S.
-data from 1987:4 to 2024:2.
-Following Gavriilidis et al.
-(2023), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to climate policy uncertainty.
-We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty.
-Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable, and allowing for smooth transitions in the intercept parameters as well as in the impact matrix.
-We find that a positive CPU shock decreases production in times of low economic policy uncertainty and slightly increases it in times of high economic policy uncertainty.
-Consumer prices, in turn, increase in response to a positive CPU shock more significantly when the level of economic policy uncertainty is high than when it is low.
-Our results are, hence, complementary to Gavriilidis et al.
-(2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S.
-economy as a supply shock.
+data from 1987:4 to 2024:12.
+Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU.
+We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU).
+Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable.
+We find that a positive CPU shock decreases production and increases inflation in times both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
+Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
 The rest of this paper is organized as follows.
 Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models.
-Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and describes a procedure for labelling the shocks in the presence of weak identification.
-Section 4 discusses estimation of the parameters by the method of maximum likelihood and stationarity of the model.
+Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks.
+Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-phase procedure.
 Section 5 presents the empirical application and Section 6 concludes.
-Appendices provide further discussion on stationarity of the model; proofs for the stated lemmas, propositions, and theorem; and details related to the empirical application.
-Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
+Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application.
+Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
 2 Reduced form STVAR models
@@
 where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
-The transition weights $\alpha_{m,t}$ are assumed to be either nonrandom and exogenous or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and they satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$.
+The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$.
 Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
@@
 See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
-It is often assumed that the coefficient matrices $A_{m,i}$, $i=1,...,p$, satisfy the usual stability condition $\det\left(I_{d}-\sum_{i=1}^{p}A_{m,i}z^{i}\right)\neq 0\ \ \text{for}\ \ |z|\leq 1,$ for each of the regimes $m=1,...,M$.
-However, this condition does not necessarily ensure that the model is stationary (see Kheifets and Saikkonen, 2020, and the references therein).
-If the transition weights are exogenous, it might be particularly challenging to establish the stationarity of the model.
-Assuming endogenous transition weights, we present sufficient conditions for establishing ergodicity, stationary, and mixing properties of our structural STVAR model in Appendix A (see also Section 4).
-Nevertheless, our main identification results discussed in Section 3 do not require stationarity of the model.
-The above-described framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics.
+Our framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics.
 To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
 where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable.
-At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be either exogenous or a lagged endogenous variable.
+At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables, for example.
 Another popular specification is to assume two regimes ($M=2$) and logistic transition weights.
@@
 The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
-The switching variable $z_{t}$ can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables.
-In addition specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2).
+In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2).
 For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$.
@@
 The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), so it is more feasible in large scale VAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
-Imposing constant AR matrices across the regimes but allowing the intercepts to vary in time is a particularly useful special case of (2.1)-(2.2), as constant AR matrices substantially decrease the number of parameters compared to the fully nonlinear specification.
-However, the model still allows for a time-varying conditional mean via time-varying intercepts as well as shifts in the volatility regime through the error process $u_{t}$.
-This specification is obtained from (2.1)-(2.2) by assuming $A_{m,i}=A_{i}$ for all $i=1,...,p$ and $M=1,...,M$, simplifying the model to:
-$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.6)
-Finally, note that while we focus on TVAR and STVAR models here for concreteness and because these models are implemented to the accompanying R package sstvars (Virolainen, 2024), the identification results of this paper directly extend to many other nonlinear SVAR models as well.
-This includes models incorporating discrete regime switches such as Markov-Switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007), and more generally models whose impact matrix is defined as a weighted sum of the impact matrices of the regimes and the time $t$ weights are known at the time $t$.
 3 Structural STVAR models identified by non-Gaussianity
@@
 The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$.
-Various types of restrictions that recover the structural shocks have been proposed in the literature.
-Many of them are discussed in Ramey (2016) and Kilian and Lütkepohl (2017), for example.
+Various types of restrictions that recover the structural shocks have been proposed in the literature, many of which are discussed in Kilian and Lütkepohl (2017), for example.
 Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al.
 (2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian.
-With Gaussian shocks, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ lead to observationally equivalent models.
-However, when the shocks are mutually independent and at most of them is Gaussian, different decompositions generally lead to models that are not observationally equivalent, as shown in Lanne et al.
-(2017) for a static impact matrix.
-We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with the impact matrix $B_{y,t}$.
+Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017).
+We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$.
 In order to parametrize the model directly with the impact matrix, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
@@
 where $e_{t}$ are independent and identically distributed structural errors with identity covariance matrix and a distribution that may depend on the parameter $\nu$.
-Since we incorporate time-variation in the impact matrix, its functional form needs to be specified.
-A natural specification for a STVAR model is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
+To incorporate time-variation in the impact matrix, we specify its functional form.
+A natural specification for STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
 $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m},$ (3.3)
@@
 Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ that are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero.
+The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero.
 This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$).
-To further improve intuition, observe that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
-In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to an impact matrix $B_{t}$ that is singular for some $t=1,...,T$ are excluded.
+To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
+In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
 Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
@@
 where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$.
-The conditional covariance matrix of $y_{t}$ can be thereby described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-To facilitate the interpretation of the conditional covariance matrix (3.4), suppose the process is completely in one of the regimes at some $t$, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$ and $\alpha_{m,t}=0$ for $m\neq\tilde{m}$.
-Then, $\Omega_{y,t}$ reduces to $B_{\tilde{m}}B_{\tilde{m}}^{\prime}=\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of the regime $\tilde{m}$.
-When the process is not completely in any of the regimes, $\alpha_{\tilde{m},t}\in(0,1)$ for some regime $\tilde{m}$$\Omega_{y,t}$ depends both on the conditional covariance matrices of the regimes $\Omega_{1},...,\Omega_{M}$ and on the cross terms $\Omega_{m,n}$.
-Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ has the form $B_{y,t}B_{y,t}^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for the covariance matrices $\Omega_{1},...,\Omega_{M}$ of the regimes.
-Specifying the conditional covariance matrix as $\Omega_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ would either require parametrizing the model directly with $\Omega_{y,t}$ or imposing some specific structure on $B_{y,t}$ that would ensure that this identity holds at all $t$.
-In the former case, identification of the shocks requires further restrictions to be imposed on the model (even when the shocks are non-Gaussian) as discussed at the beginning of this section.
-In the latter case, the restrictions are implicitly imposed by the specific form of the impact matrix ensuring that $B_{y,t}B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for all $t$.
-One of such parametrizations of $B_{y,t}$ aligns with the identification by heteroskedasticity (Lütkepohl and Netšunajev, 2017), which restricts the relative impact responses of the variables to each shock time-invariant (see also the related discussion in Virolainen, forthcoming).
-In either case, the identifying restrictions may be challenging to justify economically.
-Our specification of $\Omega_{y,t}$ in (3.3) and (3.4), in turn, facilitates exploiting non-Gaussianity of the shocks in their identification, as we show the next section.
+The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$.
+If the process is completely in one of the regimes, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of Regime $\tilde{m}$.
+When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends both on the regime-specific covariance matrices and the cross terms.
+Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes.
+Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.
 3.2 Identification of the shocks by non-Gaussianity
@@
   1. (i)
-    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, has zero mean and unit variance.
+    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, follows a continuous distribution with zero mean and unit variance.
   2. (ii)
     The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one them has a Gaussian marginal distribution.
-Note that the assumption of a unit variance of $e_{it}$, $i=1,...,d$, is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$.
+Note that the assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$.
 Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
-The next lemma then shows that, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns at each $t$.
+The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns.
 This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent.
-This lemma, which is mostly similar to Proposition 1 in Lanne et al.
-(2017), is used in the following sections to show that the parameters in the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
+This lemma, which is partly similar to Proposition 1 in Lanne et al.
+(2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
+Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) as well as the weight function parameters are identified.
+Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
 Lemma 2.
@@
 Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$. If the following conditions hold:
+  1. (1)
+    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter), and
+  2. (2)
+    $(\phi_{m,0},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n,0},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
+the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+For threshold models (2.3), also the parameters $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+For logistic models (2.4), the parameters $c,\gamma$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
 Lemma 2 is proven in Appendix B.2.
 We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1.
-In other words, at each $t$‚ changing the ordering and signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to a model that is not observationally equivalent.
+In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model.
 The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1.
+Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, because in some special cases of degenerate data (e.g., if the regime does not switch), there is insufficient variation to distinguish the parameters.
 If the impact matrix is time-invariant as in Lanne et al.
-(2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows that the structural shocks are identified up to ordering and signs.
-Such statistically identified shocks can then be labelled as economic shocks based on external information (see Lanne et al., 2017).
+(2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs.
+Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017).
 However, when the impact matrix varies in time, two complications arise.
@@
 To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-In order to fix the ordering and signs of the columns of $B_{y,t}$ it then suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$.
+To fix the ordering and signs of the columns of $B_{y,t}$, it suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$.
 This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
-It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending of the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others.
+It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending on the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others.
 Also, constraints that are identifying in some specifications may not be sufficient for identification in others.
@@
 3.3 Identification of the shocks in structural TVAR models
-We consider structural threshold VAR (TVAR) models here, as they are special cases of STVAR models, but the introduced identification results also directly extend to other models incorporating discrete regime switches, such as Markov-switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007).
 Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$.
 Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$.
-That is, we have structural TVAR model.
-By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained by fixing the ordering and signs of the columns of $B_{1},...,B_{M}$.
-However, whether this is overidentifying depends on the distributions of the shocks $e_{t}=(e_{1t},...,e_{dt})$.
-For instance, suppose the shocks follow independent Student’s $t$ distributions.
-Then, the signs of the columns of $B_{m}$ can be fixed in each regime separately without loss of generality due to the symmetry of the Student’s $t$ distribution.
-But fixing the ordering of the columns in each regime separately would generally be overidentifying because the $i$th column of $B_{m}$ is related to the $i$th shock, and each shock has its own degrees-of-freedom parameter that affects the density of the distribution (and is in our specification shared across the regimes).
-Fixing the ordering of the columns of $B_{1}$ would thereby effectively fix the ordering of the columns of $B_{2},...,B_{M}$ without restricting the model when the values of the degrees-of-freedom parameter are not equal for any of the shocks.
-On the other hand, if the degrees-of-freedom parameter values of some of the shocks are equal, the ordering of the corresponding columns can be changed in each regime separately without changing the underlying model.
-If the degrees of freedom parameter values are close to each other, reordering the columns of $B_{m}$ in one of the regimes without reordering them in the other regimes accordingly would have only a small effect of the density of the model (or to the fit of an estimated model), thereby resulting in weak identification with respect to the ordering of the columns of $B_{2},...,B_{M}$.
-As the required identification conditions depend on the distributions of the shocks, it useful to make some assumptions on them.
-Namely, we assume that the distributions of the shocks are continuous and none of the them are identical to each other in the sense that the values of their density functions are unequal almost everywhere in $\mathbb{R}$.
-This assumption allows to effectively fix the ordering of the columns of $B_{2},...,B_{M}$ by fixing the ordering of the columns of $B_{1}$, since changing the ordering would change the distribution of the corresponding shocks in some of the regimes relative to Regime 1.
-If this assumption is close to break (e.g., when the degrees-of-freedom parameter values of independent Student’s $t$ distributions are close to each other or very large), identification is weak with respect to the ordering of the columns.
-Labelling the shocks based on the estimates of the columns of each $B_{m}$ can then be unreliable, as small changes in the data could result in changes in the ordering of columns in some of the regimes.
-Therefore, we propose a procedure to facilitate reliable labelling of the shocks in the presence of such weak identification in Section 3.5.
-The following proposition, which is proved in Appendix B.3, formally states the identification results for TVAR models.
+That is, we have a structural TVAR model.
+By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained with constraints that fix the ordering and signs of the columns the impact matrix in each regime.
+To state the identification result, a specific shock distribution is assumed here for simplicity.
+Namely, we assume that each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness in the data, nests the Gaussian distribution as a special case, and is discussed in detail in Section 4.111The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
+The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR models.
 Proposition 1.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied and $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$.
-Denote the density function of the distribution of the shock $e_{it}$ as $d_{e_{i}}(\cdot;0,1,\nu_{i})$, where $\nu_{i}$ denotes the parameters of the distribution other than mean and variance.
-Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and assume the following conditions hold:
+Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied, $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$, and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively.
+Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
   1. (a)
-    the signs of the columns of $B_{m}$ are fixed for each $m=1,...,M$,
+    the ordering and signs of the columns of $B_{1}$ are fixed,
   2. (b)
-    the ordering of the columns of $B_{1}$ is fixed, and
+    $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j\in\{1,...,d\}$,
   3. (c)
-    $e_{it}$ follow continuous distributions with $d_{e_{i}}(x;0,1,\nu_{i})\neq d_{e_{j}}(x;0,1,\nu_{j})$ almost everywhere in $x\in\mathbb{R}$ for $i\neq j=1,....,d$.
-Then, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
-Condition (a) of Proposition 1 grants arbitrary identification with respect to the signs of the shocks.
-If the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, any fixed signs can be assumed without loss of generality.
-However, if the distributions of the shocks are not symmetric, the signs of the columns of the impact matrices may affect the density of the model, implying that fixing the signs is an overidentifying restriction.
-Condition (b) fixes the ordering of the columns of $B_{1}$.
-Any fixed ordering of the columns can be assumed without loss generality.
-For instance, if the first nonzero entry in each column of $B_{1}$ is assumed positive, Condition (a) is satisfied by the positivity assumption, and Condition (b) is satisfied by assuming that these entries are in a decreasing order, given that none of the entries are exactly equal to each other.
-Condition (c) is concerned with the distributions of the shocks, and it is satisfied by many of the popular distributions such as Student’s $t$ distributions with different values of the degrees-of-freedom parameters.
-It ensures that changing the ordering of the columns of $B_{2},...,B_{M}$ while keeping the ordering of the columns of $B_{1}$ fixed would change the joint density of the model.
-Thus, combined with Condition (b), it leads to identification with respect to ordering of the columns of $B_{1},...,B_{M}$.
-Finally, the result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{T}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ would not change the joint density of the model.
-Since the probability of obtaining such special cases is zero, this does not have any practical implications.
+    $\lambda_{i}\neq 0$ for all $i=1,...,d$.
+Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively.
+Any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal).
+If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering corresponding of the columns of $B_{2},...,B_{M}$.
+In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification.
+If Condition (c), in turn, fails for some $i\in\{1,...,d\}$, the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification.
+The result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
+If Condition (b) or (c) is close to break, the identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively.
+Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) are not close to break, the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4).
+Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
+As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
 3.4 Identification of the shocks in structural STVAR models
-Having established the identification result for structural TVAR models, or more generally to models incorporating discrete regime switches, we now assume that the transition weights exhibit more variation.
-Because the variation in the transition weights then translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns (by Lemma 2), the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1.
-Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Having established the identification result for structural TVAR models, we now consider the transition weights that exhibit more variation.
+The transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited.
+Because the variation in the transition weights translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns by Lemma 2, the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1.
+Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
 Proposition 2.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Suppose the following conditions hold:
+Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed).
+Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
   1. (A)
@@
   3. (C)
-    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values.
-Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the matrices $B_{1},...,B_{M}$ are uniquely identified.
+    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values.
+Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
+For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
 Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified.
-The result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that there can exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
+The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as there may exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
+For logistic models, the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, as special cases of degenerate data may lack sufficient variation to uniquely identify the transition weights (see Lemma 2).
 Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$.
-Switching the ordering of the columns of $B_{y,t}$ merely equals switching the ordering of the shocks, so any fixed ordering of the columns of $B_{1}$ can be assumed without loss of generality.
-Furthermore, if the distributions of the shocks are symmetric, which is typically assumed in nonlinear SVAR models, also the signs of the columns of $B_{1}$ can be fixed without loss of generality, as switching the sign of a column of $B_{y,t}$ equals switching the sign of the corresponding shock.
-The ordering and signs of the columns of $B_{1}$ can be fixed, for example, by assuming that the first nonzero element in each column of $B_{1}$ is positive, these elements are not exactly equal to each other, and they are in a decreasing ordering.
-The assumption of unequal first nonzero elemenents is not restrictive in practice, as the probability of obtaining an estimate such that some of the elements are exactly equal to each other within the numerical accuracy of modern statistical software is extremely small.
-See Lanne et al.
-(2017) for a more complicated strategy for fixing the ordering and signs (or more generally scalings) of the columns of a static impact matrix that is valid without such an additional assumption.
-Condition (B) of Proposition 2 states that the transition weights should, for some $t$, be strictly positive for all of the regimes.
-If there are only two regimes, the requirement is merely that the regime-switches are not always discrete (in which case Proposition 1 would be more relevant).
-Condition (C), in turn, states that the transition weights should exhibit at least a certain small degree of variation.
+If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related skewness parameter.
+Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes.
+If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be more relevant.
+Condition (C) of Proposition 2 states that the transition weights should exhibit at least a certain small degree of variation.
 For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$.
-In practice, Condition (C) should be satisfied for the observed time period $t=1,...,T$.
 This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete).
-The concerns of weak identification with respect to the ordering (and signs) of the columns of $B_{2},...,B_{M}$ with binary transition weights, discussed in Section 3.3, also translate to smooth transition models when the estimates of the transition weights are such that the weights are mostly close to zero or one.
-Surprisingly, our experience shows that even when there is a high degree of variation in the transition weights, the identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ often (if not virtually always) appears to be very weak in practice, at least if the parameters are estimated by the method of maximum likelihood.
-Specifically, there are frequently a number of local maxima of the log-likelihood function with log-likelihoods very close to each other, corresponding to mostly very similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
-Weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$ can be a problem in practice for labelling the shocks, as a given column in $B_{y,t}$ need not correspond to the same shock for all $t=1,...,T$.
-In this case, the (generalized) impulse response functions would not estimate of the effects of the intended shock but of a linear combination of the structural shocks.
-As a solution, in the next section, we propose a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method (discussed in Section 4).
-If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
-3.5 Labelling the shocks
+Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
+3.5 Labelling the shocks and blended identification
 As in the linear SVAR model of Lanne et al.
@@
 Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables.
-This is a major advantage compared to statistical identification by heteroskedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
-Additionally, in line with the statistical identification literature, any economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
-Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, because the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes.
-For example, if a positive supply shock increases output and decreases prices on impact, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs.
-However, due to the problem of weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, discussed in Sections 3.3 and 3.4, the ordering and signs of the columns of $B_{2},...,B_{M}$ is often practically arbitrary even when the ordering and signs are, in theory, uniquely identified by Proposition 1 or 2.
-Assuming that the parameters are estimated by the method of maximum likelihood (ML), weak identification with respect to the ordering and signs of the columns often results in multiple local maxima of the log-likelihood function with log-likelihoods very close to each other.
-Our estimation method, discussed in Section 4 and implemented to the R package sstvars (Virolainen, 2024), can produce a set of local solutions to the estimation problem.
-This allows the researcher to check whether there are multiple local solutions with log-likelihoods very close to each other.
-In practice, one can then simply choose from the almost equally well fitting local solutions the one that, based on economic reasoning, facilitates labelling the shock(s) of interest to the same column of the impact matrix in all regimes.
-To exemplify, consider the simplistic setup of two variables, an output variable and a price variable, two shocks, the aggregate supply shock and the aggregate demand shock, and two regimes ($M=2$).
-Suppose the estimates of the impact matrices of the two regimes with the highest log-likelihood are the following:
-$\hat{B}_{1}=\begin{bmatrix}0.67&\phantom{-}0.21\\ 0.06&-0.21\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}-0.55&1.14\\ \phantom{-}0.49&0.09\\ \end{bmatrix}.$ (3.5)
-Then, suppose there exists another local solution solution with the second highest likelihood that is almost equal to the highest log-likelihood, but the estimates of the impact matrices are the following:
-$\hat{B}_{1}=\begin{bmatrix}0.59&\phantom{-}0.16\\ 0.05&-0.17\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}1.05&\phantom{-}0.58\\ 0.18&-0.41\\ \end{bmatrix}.$ (3.6)
-In the estimates (3.5), the first shock moves output and prices in the same direction at impact in Regime 1 but in opposite directions in Regime 2.
-The second shock in, turn, moves output and prices in opposite directions in Regime 1 but in the same direction in Regime 2.
-Thus, neither of the shocks can be labelled as the aggregate supply shock nor as the aggregate demand shock.
-On the other hand, in the estimates (3.6), the first shock moves output and prices in the same direction satisfies the same signs of the impact effects in both regimes, so it can be labelled as the aggregate supply shock.
-The second shock moves output and prices in opposite directions and satisfies the same signs of the impact effects in both regimes, so it can be labelled as the supply shock.
-Since the estimates in (3.5) do not align with what is expected from the aggregate supply and demand shocks but the estimates in (3.6) do, the latter solution can be chosen based on economic reasoning despite having marginally smaller log-likelihood.
-[^1] It is also possible to obtain the estimates in (3.6) more formally by using restricted ML estimation, imposing the restriction that the first shock moves both variables in the same direction and the second shock in opposite directions on impact.
-Since the solution (3.6) yields the second highest log-likelihood among all local solutions, it is obtained from the restricted ML estimation.
-Note that also other types of information than just signs of the impact responses of the variable can be utilized to label the shocks.
-For instance, if the shock of interest should or should not significantly move certain variables on impact, the corresponding estimates of $B_{1},...,B_{M}$ should satisfy this property.
+This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
+Additionally, in line with the statistical identification literature, many economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
+Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes.
+For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs.
+However, as discussed in Sections 3.3 and 3.4, the identification is often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
+Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other.
+If the identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
+As our estimation method, discussed in Section 4, can produce a set of local solutions, the possible presence of such local solutions can be easily checked.
+To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf.
+Carriero et al., 2024).
+As a concrete example, if a supply shock should increase output and decrease prices at impact in all regimes, selecting the best local solution where the same column in all $B_{1},...,B_{M}$ satisfies such signs amounts to combining identification by non-Gaussianity with the appropriate short-run sign restrictions.
+In addition to short-run sign restrictions, also other types of information can be exploited in blended identification, including narrative restrictions and testable zero impact effect restrictions, for instance.
+4 Estimation
+4.1 Maximum likelihood estimation and ergodic stationarity
+The parameters of the structural STVAR models discussed in Section 3 can be estimated by the method of maximum likelihood (ML), which amounts to maximizing the log-likelihood function over the parameter space.
+To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary.
+Based on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one.
+But since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in estimation.
+Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition.
+After estimation, the sufficient condition can be checked for the solutions of interest.
+Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics.
+Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum.
+To overcome this issue, we propose allowing for unstable estimates, but adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close the boundary of the stability region, and then maximizing the penalized log-likelihood function.
+In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance of the optimization algorithms, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
+Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024).
+They impose nonnegativity constraints in the autoregressive conditional heteroscedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space.
+Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
+4.2 The penalized log-likelihood function
+Before introducing the penalized log-likelihood function, the log-likelihood function is presented.
+We assume that each shock $e_{it}$, $i=1,...,d$, follow the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$.
+The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
+$st(e_{it};\nu_{i},\lambda_{i})=b_{i}c_{i}\left(1+\frac{1}{\nu_{i}-2}\left(\frac{b_{i}e_{it}+a_{i}}{1-\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}\right)^{-(\nu_{i}+1)/2}$ (4.1)
+where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise.
+The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is Gamma function.
+Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
+$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (4.2)
+where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3).
+The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{M,0},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters.
+[^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights $\alpha$ is omitted from the parameter vector.
+The penalized likelihood function is then defined as
+$PL(\boldsymbol{\theta})=L(\boldsymbol{\theta})-P(\boldsymbol{\theta}),$ (4.3)
+where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space.
+Focusing on avoiding unstable estimates, we define the penalization term as
+$P(\boldsymbol{\theta})=\kappa Td\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (4.4)
+where $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization.
+Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues.
+Our penalization term incorporates the multiplicative coefficient $Td$ in attempt to standardize the strength of the penalization with respect to the number of observations $T$ and variables $d$, as the value of the log-likelihood function (4.2) tends to scale with these quantities.
+In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
+4.3 Three-phase estimation procedure
 Lanne et al.
-(2017) go further and exploit more information by labelling the shocks based on the estimated impulse response functions.
-If none of the alternative local solutions with almost equal log-likelihoods allows the shock of interest to be labelled to same column across $B_{1},...,B_{M}$, the identification can be strengthened by imposing economic overidentifying restrictions on the model.
-For example, if economic reasoning suggests that certain variables should respond to the shock of interest with a lag, the appropriate zero restrictions can be imposed on the corresponding entries of $B_{y,t}$.
-Due to the statistical identification, these zero restrictions are overidentifying and thereby testable, which a major advantage compared to using conventional short-run restrictions alone.
-4 Estimation
-The parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of maximum likelihood.
-Obtaining the log-likelihood function requires specifying a distribution for the structural errors $e_{t}=(e_{1t},...,e_{dt})$ satisfying Assumption 1.
-In our empirical application presented in Section 5, we assume that each shock $e_{it}$, $i=1,...,d$, follows a Student’s $t$ distribution with mean zero, variance one, and degrees of freedom $\nu_{i}\in(2,\infty)$.
-The assumption $\nu_{i}>2$ is made to ensure the existence of the second moments, and the assumption $\nu_{i}<\infty$ is made to distinguish the distribution from the Gaussian distribution.
-In order to normalize the variances of the structural shocks to unity, we parametrize the Student’s $t$ distribution using its variance as a parameter together with the mean and degrees-of-freedom.
-Similar parametrization has been utilized previously in Virolainen (2022) and Meitz et al.
-(2023), for example.
-The density function of such a univariate $t$-distribution with mean $\mu_{i}$, variance $\sigma_{i}^{2}>0$, and $\nu_{i}>2$ degrees-of-freedom is
-$t_{1}(y_{it};\mu_{i},\sigma_{i}^{2},\nu_{i})=\frac{\Gamma\left(\frac{1+\nu_{i}}{2}\right)}{\sqrt{\pi(\nu_{i}-2)}\Gamma\left(\frac{\nu_{i}}{2}\right)}\sigma_{i}^{-1}\left(1+\frac{(y_{it}-\mu_{i})^{2}}{(\nu_{i}-2)\sigma_{i}^{2}}\right)^{-(1+\nu_{i})/2},$ (4.1)
-where $\Gamma\left(\cdot\right)$ is the gamma function.
-We collect the parameters to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),....,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(vec(B_{1}),...,vec(B_{M}))$, $\alpha$ contains the transition weight parameters (and is omitted if exogenous weights are used), and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters.
-Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
-$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (4.2)
-where the impact matrix $B_{y,t}$ is defined in Equation (3.3), $I_{d,i}$ is the $i$th column of the $(d\times d)$ identity matrix $I_{d}$, and the Student’s $t$ density $t_{1}(\cdot;0,1,\nu_{i})$ is defined in (4.1).
-The ML estimator maximizes the log-likelihood function (4.2) with respect to the parameter $\boldsymbol{\theta}$.
-To facilitate obtaining desirable asymptotic properties for the ML estimator, such as consistency and a conventional limiting distribution, it is useful to establish the ergodicity and stationarity of the structural STVAR model.
-Based on the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one.
-However, since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in a numerical ML estimation.
-Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition.
-After the estimation, the sufficient condition can be checked for the solutions of interest.
-In practice, maximizing the log-likelihood function typically requires numerical optimization techniques.
-Particularly if the transition weights are endogenous, there are often a very large number of modes in the surface of the log-likelihood function, and large areas in the parameter space where it is flat in multiple directions.
-To address these complexities, we follow Meitz et al.
-(2023) and Virolainen (2022, forthcoming) and adopt a two-phase estimation procedure, which is ran a large number of times.
-In the first phase, a genetic algorithm is employed to find starting values for a gradient based variable metric algorithm.
-In the second phase, the variable metric algorithm is then initialized from each starting value found in the first phase, thereby often converging to a nearby local maximum.
-Some of the estimation rounds may end up in saddle points or points near the boundary of the parameters space (e.g., in points where $B_{y,t}$ is nearly singular for some $t$, or Condition A.1 in Appendix A, if imposed, is close to breaking), and some of the local solutions might be inappropriate for statistical inference (e.g., there might be only few observations from some of the regimes).
-As the log-likelihood function is highly multimodal and the estimation algorithm is ran a large number of times, it often results in a variety of local solutions.
-After the estimation rounds have been ran, the solution that maximizes the log-likelihood function among the appropriate local solutions can be selected.
-Then, if the transition weights are endogenous and a stationary solution is preferred, the sufficient condition for ergodic stationarity of the model (Assumption A.1(I) in Appendix A) can be checked for the local solutions of interest to select the best solution whose stationarity can be verified.
-If there are multiple such solutions with almost equal log-likelihoods, the best local solution that facilitates labelling the shocks of interest can selected based on the procedure discussed in Section 3.5.
-The accompanying CRAN distributed R package sstvars (Virolainen, 2024) implements the two-phase estimation procedure utilizing a modified genetic algorithm.
+(2017) propose a three-phase estimation procedure for their linear SVAR model identified by non-Gaussianity.
+In the first phase, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent phases to maximize the log-likelihood function.
+The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function.
+To address these challenges, we propose a modified version of the three-phase estimation procedure in which the multimodality is explicitly taken into account.
+Our three-phase estimation procedure operates with the following steps:
+  1. 1.
+    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1,0},...,\phi_{M,0},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) as described in Appendix C.
+Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
+  2. 2.
+    Estimate the error distribution parameters $\boldsymbol{\theta}_{2}\equiv(\sigma,\nu)$ by (penalized) ML with a genetic algorithm conditionally on the (penalized) NLS estimate $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
+Denote the obtained estimate of $\boldsymbol{\theta}_{2}$ as $\hat{\boldsymbol{\theta}}_{2}^{GA}$.
+  3. 3.
+    Estimate the full parameter vector $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ by (penalized) ML by initializing a gradient based optimization algorithm from $(\hat{\boldsymbol{\theta}}_{1}^{NLS},\hat{\boldsymbol{\theta}}_{2}^{GA})$.
+Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al.
+(2017) but with the nonlinearity of the model explicitly accounted for.
+Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-phase procedure.
+Step 2 is comparable to Step 2 of Lanne et al.
+(2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able escape from local maxima.
+Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
+Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results.
+Since multiple estimation rounds are ran, the estimation procedure produces a set of estimates, possibly corresponding to different local maximums of the (penalized) log-likelihood function.
+If there are multiple solutions with (penalized) log-likelihoods almost equal to the highest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest.
+The proposed three-phase estimation procedure is implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
+To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-phase estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D.
+Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$.
+To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
+According to the results (Tables D.2 and D.3 in Appendix D), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias.
+Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models.
+While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for a degrees-of-freedom parameter, and for a certain entry in each impact matrix, explainable by the statistical properties of the true data generating process.
+However, the bias diminishes for all of the parameters as the sample size increases.
+Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator.
+Overall, the performance of the proposed PML estimator and the three-phase estimation procedure seems reasonable.
 5 Empirical application
-Our empirical application studies the macroeconomic effects of the climate policy uncertainty shock.
-Currently, there exist only a few studies on the subject.
-Fried et al.
-(2021) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy risk decreases total capital stock, causing output and carbon emissions to decrease.
-Gavriilidis et al.
-(2023), in turn, propose a news based index on measuring climate policy uncertainty (CPU), and the study the effects of the CPU shock using a linear recursively identified SVAR model and local projections.
-They find that a positive shock to their CPU index (CPUI) decreases real activity and increases prices, thus, transmitting to the U.S.
-economy as a supply shock.
-We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty.
-Such variation can be accommodated by our structural STVAR model by specifying a logistic transition weight function with the lagged level of economic policy uncertainty as the switching variable.
-This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more nuanced results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
+Our empirical application studies the macroeconomic effects of the climate policy uncertainty (CPU) shock.
+In the related literature, Fried et al.
+(2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output.
+Khalil and Strobel (2023) reach similar conclusions in their DSGE model, attributing the effects to financial institutions’ aversion to uncertainty.
+Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
+We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU).
+Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable.
+This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
 5.1 Data
 We consider a monthly U.S.
-dataset consisting of six variables and covering the time period from 1987:4 to 2024:2.
-For measuring climate policy uncertainty, we make use of the CPUI data (Gavriilidis, 2021), whose availability determines the beginning of our sample period.
-The CPUI is constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
-In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty.
-As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking the first differences.
-For measuring the price level, we use the consumer price index (CPI) and the producer price index (PPI), which are detrended by taking the first differences.
-As a forward looking variable, the producer price index should also help to distinguish the effects of climate policy uncertainty shocks from climate policy actions (see Gavriilidis et al., 2023).
-Finally, as a measure of the monetary policy stance, we include an interest rate variable (RATE), which is the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods.
-The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, and the Federal funds rate are retrieved from the Federal reserve bank of St.
-Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
+dataset consisting of six variables and covering the time period from 1987:4 to 2024:12.
+Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
+The availability of the CPUI data determines the beginning of our sample period.
+In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
+As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences.
+For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences.
+Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE).
+The series of the included variables are presented in Figure 1.
 Figure 1: Monthly U.S.
-time series covering the period from 1987:4 to 2024:2.
-From top to bottom, the variables are the climate policy uncertainty index (divided by $10$), economic policy uncertainty index (divided by $10$), the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$), and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods.
-The bottom panel shows the estimated transition weights of the first regime of the fitted two-regime fifth-order LSTVAR model.
+time series covering the period from 1987:4 to 2024:12.
+From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods.
+The bottom panel shows the transition weights of High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model.
 The shaded areas indicate the U.S.
 recessions defined by the NBER.
-The series of the included variables are presented in Figure 1.
-Both CPUI and EPUI show a generally increased level of uncertainty towards the end of the sample period, suggesting that a model accommodating shifts in the mean could be appropriate.
-Moreover, there is clearly variation in the volatility of the variables, so shifts in the volatility regime should also be incorporated into the model.
-5.2 Logistic STVAR model identified by non-Gaussianity
-In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks need to be specified.
-As transition weights, we employ the popular logistic transition weights (Anderson and Vahid, 1998) given in Equation (2.4) for a two-regime model.
-They facilitate smooth transitions between the regimes depending the level of the switching variable, and they also nest the TVAR model (Tsay, 1998) as a special case when the scale parameter $\gamma$ tends to infinity.
-To study how the effects of the CPU shock vary depending on the level of economic policy uncertainty, we specify the first lag of EPUI as the switching variable.
-This approach accommodates shifts in the mean and volatility depending on the level of EPUI, and hence allows the effects of the CPU shock to vary with the level of economic policy uncertainty.
-We assume that the structural shocks follow independent Student’s $t$ distributions, which is a fairly standard choice for modelling fat tailed data.
-Moreover, as the Gaussian distribution is obtained as a special case of the $t$-distribution, with the value of the degrees-of-freedom parameter tending to infinity, the validity of the identifying Assumption 1 can be assessed based on the estimates of the degrees-of-freedom parameters.
-In order to maintain a reasonable number of parameters in the model and to avoid the problems of potentially imprecise estimates and overfitting, we assume that the autoregression matrices are constant across the regimes, i.e., employ the specification in Equation (2.6).
-Due to the time-varying intercepts and impact matrices, the model is still fairly general, allowing for shifts in the mean, volatility, and impact responses of the variables to the shocks.
-Our structural two-regime logistic STVAR (LSTVAR) model can be written as
-[…]
-where $c$ is the location parameter, $\gamma>0$ is the scale parameter, $y_{2t-1}$ is the lagged EPUI, the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ follow independent $t$-distributions with zero mean and unit variance, and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters of the structural shocks.
-Finally, we select the autoregressive order $p=5$ for our model based on AIC.
-The estimates of the transition weights of Regime 1 are presented in the bottom panel of Figure 1.
-They show that Regime 1 dominates in the periods of low economic policy uncertainty, whereas Regime 2 conversely dominates in the periods of high economic policy uncertainty.
-The location parameter estimate is $\hat{c}=9.80$ and the scale parameter estimate is $\hat{\gamma}=1.00$.
-Since the sample mean of EPUI over our sample period is $11.48$, Regime 2 starts dominating roughly when EPUI surpasses its mean, and based on the small estimate of the scale parameter, the transitions between the regimes are smooth.
-Therefore, we label Regime 1 as the low economic policy uncertainty (EPU) regime and Regime 2 as the high EPU regime.
-The estimates of the degrees of freedom parameters are small, $\hat{\nu}=(3.86,5.54,2.94,2.61,2.02,2.23)$, and therefore none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied.
-Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the matrices $\boldsymbol{A}_{m}$ (A.2), $m=1,2$, using the JSR toolbox (Jungers, 2023) in MATLAB.
+5.2 The structural LSTVAR model
+In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified.
+We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4).
+They facilitate smooth transitions between the regimes depending the level of the switching variable, enabling an analysis of how the effects of the CPU shock vary depending on the level of EPU by specifying its first lag as the switching variable.
+The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating selection between these two weight functions within estimation.
+We estimate the structural LSTVAR model by PML with the three-phase procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions.
+The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case of the skewed $t$-distribution, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters.
+To avoid overfitting, particularly in the regime prevailing when the level of EPU is high, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
+To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that avoids making restrictive assumptions about the effects of the CPU shock.
+We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes.
+Then, to support our identification, we check that the recovered CPU shocks align with several major historical events affecting CPU.
+In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively).
+The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition.
+Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
+The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU.
+Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime.
+The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1 where the fitted transition weights of High EPU Regime are presented.
+The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively.
+Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied.
+Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes using the JSR toolbox (Jungers, 2023) in MATLAB.
 The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1.
-Hence, the ML estimator is expected to be well-behaved.
-To assess the adequacy of our model, we utilize residual based model diagnostics to examine how well our model captures the autocorrelation structure, conditional heteroskedasticity, and distribution of the data.
-To that end, we plot several diagnostic figures, which are presented and discussed in more detail in Appendix B.6.
-They show that our model captures the autocorrelation structure of the data reasonably well, but there is a large correlation coefficient in the crosscorrelation functions between two of the residual series.
-Also conditional heteroskedasticity is somewhat reasonably well captured, although there are a few relatively large correlation coefficients in the auto- and crosscorrelation functions of squared standardized residuals.
-The marginal distributions of the series are also reasonably well captured, but the distributions of two of the residual series are somewhat asymmetric.
-Nonetheless, in our view, the overall adequacy of the model is reasonable enough for impulse responses analysis.
-5.3 Labelling the CPU shock
-The shocks of our structural LSTVAR model (5.1) are readily identified by Proposition 2.
-However, labelling the shocks requires external information.
-Moreover, due to the problem of weak identification discussed in Section 3.5, we should choose among the practically equally well fitting local solutions the one that facilitates labelling the CPU shock to the same column of the impact matrix in all of the regimes.
-The estimates of $B_{1}$ and $B_{2}$ for the local solution (discussed in Section 5.2) with the highest obtained log-likelihood are the following:
-[…]
-where the ordering of the variables is $y_{t}=(\text{CPUI}_{t},\text{EPUI}_{t},\text{IPI}_{t},\text{CPI}_{t},\text{PPI}_{t},\text{RATE}_{t})$, and the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are fixed by normalizing the entries in the first row of $\hat{B_{1}}$ to be positive and in a decreasing order.
-In terms of labelling the CPU shock based on the impact responses of the variables in the estimates of $B_{1}$ and $B_{2}$, there are several properties that are reasonable to expect from it.
-First, since the CPU shock is the shock to CPUI, the impact response of CPUI should be substantial, and it should have the same sign in $B_{1}$ and $B_{2}$.
-Second, as CPU contributes to the overall economic policy uncertainty, EPUI should increase or at least not significantly decrease in response to a positive CPU shock.
-The effects of the CPU shock on IPI are more ambiguous.
-Increased uncertainty may postpone investments, decreasing production, while some firms may accelerate production to avoid future costs or invest in future technologies to gain a competitive edge.
-The CPU shock’s impact on the price level, in turn, may be indirect, translated through the effects on the aggregate demand, production costs, or production levels.
-Similarly, the Fed may not directly respond to increased uncertainty but rather to its economic impacts.
-Based on the estimates of $B_{1}$ in Equation (5.2), the only variable that moves substantially at impact in response to the first shock is CPUI, and the first shock has also clearly the highest impact effect on CPUI.
-The first shock also moves EPUI in the same direction with CPUI, in line with increased CPU contributing to the level of general economic policy uncertainty.
-Also the second shock moves CPUI considerably on impact, but it moves EPUI much more, so since CPUI and EPUI vary in the same scale, it does not appear to be a plausible CPU shock candidate.
-Therefore, the first shock seems to be the most reasonable candidate as the CPU shock.
-The estimates of $B_{2}$ in Equation (5.2) show that the first shock substantially moves CPUI at impact, while the response of CPUI to the other shocks is clearly smaller.
-Importantly, the sign of the response of CPUI to the first shock is the same as in $\hat{B}_{1}$.
-Moreover, the first shock moves EPUI in the same direction with CPUI, thus, in line with what can be expected from a CPU shock.
-Overall, the first shock seems to be the most plausible CPU shock candidate, and hence, we deem it as the CPU shock.
-Since we labelled the CPUI shock in the local solution that obtained the highest likelihood, there is no need to consider other local solutions incorporating practically equal fit and mostly very similar estimates but different ordering and signs of the columns of $B_{2}$ (see Section 3.5).
+Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
 5.3 Generalized impulse response functions
@@
 The latter term on the right side is the expected realization of the process conditionally on the previous observations only.
-The GIRF thus expresses the expected difference in the future outcomes when the structural shock of sign and size $\delta_{i}$ in the $i$th element of $e_{t}$ hits the system at time $t$ as opposed to all shocks being random.
-The generalized impulse response functions also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.3).
-As our structural LSTVAR model has a $p$-step Markov property, conditioning on (the $\sigma$-algebra generated by) the $p$ previous observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is effectively the same as conditioning on $\mathcal{F}_{t-1}$ at time $t$ and later.
-In order to study state dependence of the effects of the CPU shock, we obtain the histories $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ from low EPU and high EPU regimes separately.
-Specifically, for each regime, we take all of the length $p$ histories in the data that imply transition weights larger than $0.75$ for the given regime, so that the given regime is clearly dominant, and the number of histories from both of the regimes is reasonably large.
-For low EPU regime, there are $112$ such histories and for high EPU regime there are $202$.
-Then, for each regime separately, we calculate the GIRFs for these length $p$ histories, and take the sample mean to obtain a point estimate for the GIRF conditional on dominance of the given regime, whereas taking sample quantiles gives confidence intervals that reflect uncertainty about the initial state of the economy within the dominating regime.
-This way the GIRFs estimate the effects of the CPU shock in times of low EPU and high EPU in a way that reflects the observed data.
-Our Monte Carlo algorithm for computing the GIRF is similar to the one described in Lanne and Virolainen (2024), Appendix B.
+The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random.
+The GIRFs also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
+Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$.
+To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives.
+Specifically, for each Regime $\tilde{m}\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{\tilde{m},t}$ is greater than $0.75$, indicating that Regime $\tilde{m}$ is clearly dominant (there are $366$ such histories for Low EPU Regime and $80$ for High EPU Regime).
+For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B.
+Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
 Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead.
-From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of Regime 1 are depicted in each row.
+From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of High EPU Regime are depicted in each row.
 The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels.
-The left column shows the responses to a positive one-standard-error CPU shock with initial values from Regime 1, whereas the right column shows the responses with initial values from Regime 2.
-All GIRFs have been scaled so that the instantaneous increase in CPUI is $5$.
-The shaded areas depict the $95\%$ confidence intervals that reflect uncertainty about the initial state of the economy within the regime.
-The GIRFs to a one-standard-error positive CPU shock are presented in Figure 2.
-We also computed the GIRFs for two-standard-error and negative CPU shocks, but the responses are similar to the positive one-standard-error shock (after adjusting for the different sign or size), so they are not presented for brevity.
-The responses of CPUI and EPUI follow similar patterns during times of low and high EPU, with CPUI responding substantially and EPU more weakly.
-The responses of the macroeconomic variables, however, differ across the regimes.
-When the level of EPU is low, production decreases following a positive CPU shock, possibly because of firms postponing investments (see Fried et al., 2021).
-On the other hand, when EPU is high, production slightly increases.
-A possible explanation is that some firms, already operating in a more unpredictable environment, continue with planned investments, or hedge against multiple types of uncertainties by ramping up production in sectors less likely to be immediately affected or by investing in a broader range of technologies.
-Consumer prices initially decrease and then slightly increase when EPU is low, while they increase more significantly when EPU is high.
-The initial decrease in prices under low EPU could be due to reduced consumer demand as people postpone consumption in anticipation of new policies.
-Under high EPU, prices rise more substantially, possibly because of firms adapting their operations to meet potential new regulations, increasing production costs and passing these costs onto consumers.
-Producer prices rise permanently in both regimes, possibly due to anticipatory adjustments to future costs.
-The effects of the CPU shock on the interest rate are minor but consistent with the responses of the other variables, with interest rates increasing as prices rise, more so during high EPU.
-Finally, the responses of the transition weights indicate a shift towards the high EPU regime, which is expected, as a positive CPU shock increases EPUI.
-Overall, our results suggest that a positive CPU shock decreases production in times of low EPU but slightly increases it in times of high EPU.
-Prices and the interest rate increase in both regimes, but the increase in consumer prices (and interest rate) is more significant when level of EPU is high.
-Therefore, our results are complementary to Gavriilidis et al.
-(2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S.
-economy as a supply shock.
-6 Conclusions
+The left column shows the responses to conditionally on High EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on High EPU Regime dominating.
+All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points.
+Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
+The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRFs using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf.
+Lanne and Virolainen, 2025).
+The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in Low EPU Regime and decreases in High EPU Regime.
+A possible explanation for the decrease in the EPUI could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
+Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes.
+Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al.
+(2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital.
+Although many of the GIRFs exhibit a short-term increase of IPI in High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
+In both regimes, a positive CPU shock increases consumer prices, aligning with Huang and Punzi (2024) and Khalil and Strobel (2023).
+Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a positive supply shock, whereas according Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup.
+However, we find that the inflationary effects are stronger in High EPU Regime than in Low EPU Regime.
+One possible explanation is that the firms are more risk averse in the periods of high EPU, thereby reacting more sensitively and passing the increased costs or markups onto consumers.
+Producer prices, in turn, mostly decrease in Low EPU Regime but increase in High EPU Regime.
+The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases both regimes, but clearly more so in High EPU Regime.
+Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU.
+Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
+6 Conclusion
 Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017).
 We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes.
-The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations.
+The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well.
 Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
-The identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes, and an incorrect ordering or signs of the columns results in impulse response functions that do not estimate the effects of the intended shock but of a linear combination of the structural shocks.
-In other words, there is often multiple local maxima of the log-likelihood function with almost equal log-likelihood that are very similar to each other but incorporate different ordering and signs of the columns of the impact matrices in some regimes.
-As a solution, we proposed a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method.
-If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
-In addition to discussing the identification of the shocks, we discussed the estimation of the model parameters by the method maximum likelihood in case of independent Student’s $t$ shocks, and proposed a two-phase estimation procedure that can produce a set of alternative local solutions.
-By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model.
-While useful for obtaining a well-behaving estimator, stationarity is not, however, required for applicability of our identification results.
-The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
+In addition to discussing the identification of the shocks, we discussed the problem of labelling them, and how to utilize blended identification (cf.
+Carriero et al., 2024) to solve the problem of weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices.
+We discussed estimation of the model parameters by the method of penalized maximum likelihood, and proposed a three-phase estimation procedure by adapting the procedure of Lanne et al.
+(2017) to nonlinear SVAR models.
+By making use of the results of Saikkonen (2008), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model.
+While useful for obtaining a well-behaving estimator, stationarity is not, however, required for the applicability of our identification results.
+The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
 Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S.
-data from 1987:4 to 2024:2.
+data from 1987:4 to 2024:12.
 As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
-We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable and allowing for smooth transitions in the intercept parameters as well as in the impact matrix.
-We found that a positive CPU shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty.
-Consumer prices, in turn, increase more significantly when the level of economic policy uncertainty is high than when it is low.
-Our results are, therefore, complementary to Gavriilidis et al.
-(2023), who found a positive CPU decreasing real activity and increasing prices, thus, transmitting to the U.S.
-economy as a supply shock.
+We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable.
+We found that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
+Our results are, therefore, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
 Appendix A Stationarity and ergodicity
@@
 Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013).
-The accompanying R package sstvars (Virolainen, 2024) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required.
+The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required.
 The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@
 B.1 Proof of Lemma 1
-We start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Before proving Lemma 1, we establish the following lemma that made use of in the proof:
+Lemma B.1.
+Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variable $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
+$P(x_{1},...,x_{k})=0$ (B.1)
+as $\mathcal{X}\equiv\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$.
+Then, the set $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$, unless Equation (B.1) is identically true.
+Proof.
+If $k=1$, the proof is trivial.
+Suppose $k>1$.
+It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},...,x_{k,0})\in\mathbb{R}^{k}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, the set $\mathcal{X}$ is an $k-1$-dimensional manifold locally in a neighbourhood of $x^{0}$.
+Because this argument applies at all points $x^{0}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, it follows that the set $\mathcal{X}$ is an $k-1$-dimensional manifold in regions where this condition holds.
+Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{k}$ in the region where the gradient of $P(x_{1},...,x_{k})$ is nonzero.
+On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P(x_{1},...,x_{k})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{k}$.
+Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$.
+∎
+To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
 First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$.
 Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible.
-The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M})\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
+The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
 Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$.
@@
 Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero.
-It then follows from Implicit function theorem that in the points $B^{0}=(vec(B_{1,0}),...,vec(B_{M,0}))\in\mathbb{R}^{Md^{2}}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold locally in a neighbourhood of $B^{0}$.
-Because this argument applies at all points $B^{0}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, it follows that the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold in regions where this condition holds.
-Therefore, the Lebesque measure of the set $S_{t}$ is zero in $\mathbb{R}^{Md^{2}}$ in the region where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero.
-On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P_{t}(B_{1},...,B_{M})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
-Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
+It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
 Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$ for each $t$.
@@
 B.2 Proof of Lemma 2
-For most parts, our proof is similar to the proof of Proposition 1 in Lanne et al.
+The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al.
 (2017) making use of results from independent component analysis (Comon, 1994).
@@
 where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
-Then, suppose Equation (B.1) holds also for some other coefficients, impact matrix, and shocks $\tilde{\phi}_{t},\tilde{\boldsymbol{A}}_{t}$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$, i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
+Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{t}$ and $\tilde{\boldsymbol{A}}_{t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
 By subtracting this identify from (B.2), we obtain
-$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}$ $=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}$ (B.2)
+$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
 Taking conditional expectation conditional on $\mathcal{F}_{t-1}$ gives
@@
 Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.
-To complete proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$.
+To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$.
 By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$.
@@
 That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$.
-Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.∎
+Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Threshold weights.
+To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3).
+Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$.
+That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
+At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{0,m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$.
+Consequently, Equation (B.4) can be written as
+$\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
+where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$.
+Now, Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$.
+Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true.
+Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t},0}=\phi_{m_{t},0}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$.
+Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so that the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$.
+For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches.
+Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$.
+Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+This concludes that the parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+In other words, two distinct parameter sets give at most a measure-zero set of solutions for $(y_{1},...,y_{T})$.
+Logistic weights.
+Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$.
+Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$.
+The proof proceed in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then we use these results to show that (3) the intercepts and autoregression matrices are uniquely identified.
+By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{t}=\phi_{t}$ and $\tilde{\boldsymbol{A}}_{t}=\boldsymbol{A}_{t}$ (irrespective of the weight function dynamics).
+By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
+$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\phi_{2,0})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1,0}-\tilde{\phi}_{2,0})=0$ (B.7)
+and
+$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-A_{2,i})-\tilde{\alpha}_{2,t}(\tilde{A}_{1,i}-\tilde{A}_{2,i})=0,\ \ i=1,...,p,$ (B.8)
+respectively.
+Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in the (two) variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$.
+Therefore, for given coefficients, Equations (B.7) and (B.8) have at most one solution unless at least $d(1+pd)-1$ scalar equations are identically true.
+All the $d(1+pd)$ scalar equations are identically true only if $\phi_{1,0}=\phi_{2,0}$, $\tilde{\phi}_{1,0}=\tilde{\phi}_{2,0}$, $A_{1,i}=A_{2,i}$, and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$, which violates Condition (2) of Lemma 2.
+If exactly $d(1+pd)-1$ scalar equations are identically true, the coefficients for all but one of the scalar equations are zero, and $\alpha_{2,t}$ can be expressed in the form $\alpha_{2}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some constants $a_{0},a_{1}\in\mathbb{R}$.
+Since any single solution of $\alpha_{2,t},\tilde{\alpha}_{2,t}$ can be expressed in the form $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some $a_{0},a_{1}\in\mathbb{R}$, it follows that the solution to Equations (B.7) and (B.8) must satisfy $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ (for all $t$).
+Making the dependence of the transition weights of $z_{t}$ explicit in the notation, observe that in the identity $\alpha_{2,t}(z_{t})=a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$, the logistic functions $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are both continuous and strictly increasing maps of $z_{t}$ into $(0,1)$ (as $\gamma,\tilde{\gamma}>0$).
+Specifically, $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ increase smoothly from 00 to $1$ in an ”S”-shaped curve as $z_{t}$ goes from $\infty$- ∞ to $\infty$.
+Denote $h(z_{t})\equiv\alpha_{2,t}(z_{t})-a_{0}-a_{1}\tilde{\alpha}_{2,t}(z_{t})$, which is continuous in $z_{t}$ as $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are continuous.
+Suppose $h(z_{t})=0$ is not identically true, and suppose $a_{1}\neq 1$.
+Then, $a_{1}\tilde{\alpha}_{2,t}(z_{t})$ takes values from 00 to $a_{1}$.
+If $a_{1}=0$, then $h(z_{t})=0$ at most in a single point by strict monotonicity of $\alpha_{2,t}(z_{t})$.
+If $a_{1}<0$, then $h(z_{t})$ is strictly increasing, so $h(z_{t})=0$ at most in a single point.
+If $a_{1}\in(0,1)$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})<\tilde{\alpha}_{2,t}(z_{t})$, whereas if $a_{1}>1$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})>\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}\in\mathbb{R}$.
+In either scenario, $h(z_{t})$ can cross zero only a finite number of times, as otherwise repeated crossing would force repeated sign changes, clearly contradicting the smooth logistic shapes that are “similar” but scaled.
+On the other hand, $h(z_{t})=0$ cannot hold on any set of positive Lebesgue measure because $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are strictly increasing logistic functions, and two such functions that coincide on an interval must coincide on $\mathbb{R}$, thereby forcing $a_{0}=0$ and $a_{1}=1$.
+On the other hand, if $a_{1}=1$ and $a_{0}\neq 0$, then $\alpha_{2,t}(z_{t})$ and $a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$ differ by constant shifting, and they can intersect in at most finitely many points.
+Consequently, $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ only in sets of Lebesgue measure zero in $z_{t}\in\mathbb{R}$, implying that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ (and thereby also $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$) almost everywhere in $z_{t}\in\mathbb{R}$.
+Thus, also $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ and $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$ for all $t$ almost everywhere in $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+Next, we make use of the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ to show that $\tilde{c}=c$ and $\tilde{\gamma}=\gamma$.
+By substituting the logistic weight function and simplifying, the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ can be expressed in the form
+$(\tilde{\gamma}-\gamma)z_{t}=\tilde{\gamma}\tilde{c}-\gamma c.$ (B.9)
+If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$.
+If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$.
+Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $z_{t}\in\mathbb{R}$.
+We have established that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$, $\tilde{\gamma}=\gamma$, and $\tilde{c}=c$, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+Consequently, Equations (B.7) and (B.8) simplify to
+$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\tilde{\phi}_{1,0})+\alpha_{2,t}(\tilde{\phi}_{2,0}-\phi_{2,0})=0$ (B.10)
+and
+$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-\tilde{A}_{1,i})+\alpha_{2,t}(\tilde{A}_{2,i}-A_{2,i})=0,\ \ i=1,...,p,$ (B.11)
+respectively.
+Treating and $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$ as fixed constants, Equations (B.10) and (B.11) define a system of $d(1+pd)$ linear equations and $2d(1+pd)$ variables ($\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$).
+Hence, there exists at most one solution to this system linear equations.
+Since one solution is $\tilde{\phi}_{m,0}=\phi_{m,0},\tilde{A}_{m,1}=A_{m,1},...,\tilde{A}_{m,p}=A_{m,p}$, $m=1,2$, this is the unique solution.
+This concludes the proof that the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$ as well as the parameter $r_{1},...,r_{m-1}$ models with threshold and the parameters $c,\gamma$ for models with logistics transition weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+Exogenous nonrandom weights.
+Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom).
+Then, arguments similar to those used for logistic weights, but applied for any finite number of regimes, show that the parameters $\phi_{m,0},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.∎
 B.3 Proof of Proposition 1
@@
 After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
-Condition (a) readily fixes the signs of the columns of $B_{1},...,B_{M}$, so it suffices the check that under Conditions (c) and (b), changing the ordering of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model.
+First, we establish the following lemma and its corollary, the latter of which is used in the proof:
+Lemma B.2.
+Suppose $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ (Condition (b) of Proposition 1).
+Then, $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$, where the density function $st(\cdot;\nu_{i},\lambda_{i})$ is given in Equation (4.1).
+Moreover, if $\lambda_{i}\neq 0$, then the equation $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$, and thus, $st(x;\nu_{i},\lambda_{i})\neq st(-x;\nu_{i},\lambda_{i})$ almost everywhere in $x\in\mathbb{R}$.
+Proof.
+Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$.
+It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable.
+It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is an one-dimensional manifold locally in a neighbourhood of $x^{0}$.
+Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds.
+Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero.
+On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$.
+Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$.
+That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
+To prove the moreover part, suppose $\lambda_{i}\neq 0$ and $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$, which simplifies to
+$\left(\frac{b_{i}x+a_{i}}{1-\mathbbm{1}\{x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}=\left(\frac{-b_{i}x+a_{i}}{1-\mathbbm{1}\{-x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{-x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2},$ (B.12)
+where $a_{i}\neq 0$ and $b_{i}\neq 1$.
+If $x<-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$, a direct calculation shows that the only solution to Equation (B.12) is $x=0$.
+If $x<-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$, Equation (B.12) is a second degree polynomial equation in $x$ and has thereby at most two real solutions.
+Combining all the four possible cases, in can be concluded that $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$.
+∎
+Corollary B.1.
+Suppose $D$ is a diagonal matrix with ±$\pm 1$ diagonal entries and $P$ is a permutation matrix, and denote by $I_{d,i}$ the $i$th column of the $d$-dimensional identity matrix.
+Moreover, suppose the entries $x$ are distinct as is and after changing any of their signs.
+If $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j=1,...,d$ and $\lambda_{i}\neq 0$ for all $i=1,...,d$, and in addition, $D\neq I_{d}$ or $P\neq I_{d}$, then
+$\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})\neq\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$ (B.13)
+almost everywhere in $x\in\mathbb{R}^{d}$.
+Proof.
+Suppose $\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})=\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$. Rearranging and taking logarithm yields
+$\sum_{i=1}^{d}(\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i}))=0.$ (B.14)
+Since $DPx$ reorders or swaps the signs of some of the elements of $x$ (but does not leave all the elements unchanged), by Lemma B.2 (and strict monotonicity of logarithm),
+$f_{i}(x)\equiv\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})\neq 0$ (B.15)
+for at least one $i\in\{1,...,d\}$ over sets of positive Lebesgue measure. Therefore, Equation (B.14) can hold almost everywhere only if the nonzero terms in the sum cancel each other out.
+Denote $\mathcal{I}(x)\equiv\{i\in\{1,...,d\}|f_{i}(x)\neq 0\}$ and $g(x^{*})\equiv\sum_{i\in\mathcal{I}}f_{i}(x)$, where $x^{*}$ is the $(d^{*}\times 1)$-dimensional ($d^{*}\leq d$) subvector of $x$ containing the components of $x$ for which $I_{d,i}^{\prime}x\neq\pm I_{d,i}^{\prime}DPx$, $i=1,...,d$ (i.e., $x^{*}$ captures all variables contributing to any nonzero $f_{i}(x)$).
+It is straightforward to show (by induction) that the function $\log st(x_{i};\nu_{i},\lambda_{i})$, $x_{i}\in\mathbb{R}$, is infinitely continuously differentiable.
+It follows that also $g(x^{*})$ is infinitely continuously differentiable, and thus $d^{*}$ times continuously differentiable.
+Then, arguments based on Implicit function theorem and Sard’s theorem (Sard, 1994) can be used similarly to the proof of Lemma B.2 to conclude that the set of solutions to $g(x^{*})=0$ has a Lebesgue measure zero in $x^{*}\in\mathbb{R}^{d^{*}}$.
+Hence, as sets of zero Lebesgue measure in $\mathbb{R}^{d^{*}}\subset\mathbb{R}^{d}$ have zero Lebesgue measure in $\mathbb{R}^{d}$, Equation (B.14) can hold only in sets of zero Lebesgue measure in $x\in\mathbb{R}^{d}$, concluding the proof.
+∎
+Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model.
 Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$.
 Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
-$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{t})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (B.5)
-where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $d_{e_{i}}(\cdot;0,1,\nu_{i})$ is the density function the $i$th shock.
-At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$, implying the that the density (B.5) can be expressed as
-$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}),\ \ \text{and}$ (B.6)
-$f_{y}(y_{t}-\mu_{t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i}),$ (B.7)
-where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices that satisfy $\alpha_{m,t}=1$ for the regime $m$, and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but its dependence on the other parameters is omitted).
-Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $B_{m}(\boldsymbol{\tau}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ of its columns.
-Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$.
-By the invertibility of $B_{m}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{t})$ are equal to each other almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.
-Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it then follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m})(y_{t}-\mu_{t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m})(y_{t}-\mu_{t})$ are in a different order.
-Moreover, as $|\det(B_{m})|$ is invariant to the ordering of the columns of $B_{m}$, it follows from this and Condition (c) that, almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, $f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m}))$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$.
+$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (B.16)
+where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $st(\cdot;\nu_{i},\lambda_{i})$ is the density function the $i$th shock given in Equation (4.1).
+At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $r_{m-1}<z_{t}\leq r_{m}$, where the threshold parameters, as well as $B_{y,t}$ up to ordering and signs of its columns, are identified by Lemma 2 regardless of whether $B_{1},...,B_{M}$ are identified.
+Consequently the density (B.16) can be expressed as
+$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}),\ \ \text{and}$ (B.17)
+$f_{y}(y_{t}-\mu_{y,t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i}),$ (B.18)
+where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices (identified by Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
+Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$.
+Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns.
+Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well as the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
+By the invertibility of $B_{m}$, almost everywhere in $y_{t}\in\mathbb{R}^{d}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{y,t})$ are equal to each other as is nor after changing the signs of some of the entries.
+Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ are in a different order.
+Likewise, changing the signs of the columns of $B_{m}$ changes the signs of the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ have different signs (in addition to possibly being in a different order).
+By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1.
+Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$.
 Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
@@
 $f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\boldsymbol{\tau}_{2}),...,B_{M}(\boldsymbol{\tau}_{M}))\neq f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\tilde{\boldsymbol{\tau}}_{2})...,B_{M}(\tilde{\boldsymbol{\tau}}_{M}))$ (B.20)
-almost everywhere in $Y_{T}\in\mathbb{R}^{T}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ for any $m=2,...,M$.
-Since the this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.
-Finally, by Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$.
+almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$.
+Since this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.
+By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$.
 Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns.
-Under the conditions of Proposition 1, the signs of the columns of $B_{1},...,B_{M}$ and the ordering of the columns $B_{1}$ are directly fixed.
-Moreover, by the above discussion, different permutations of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, i.e., to models that are not observationally equivalent.
-Thus, only models corresponding to the same permutations of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.∎
+Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed.
+Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$, i.e., to models that are not observationally equivalent.
+Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.∎
 B.4 Proof of Proposition 2
 To proof consists of two parts.
-First, we show that almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, if the ordering and signs of the columns of $B_{1}$ is fixed, changing ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix.
+First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix.
 In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$.
-Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified.
+For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$ (where $z_{t}$ is the switching variable), whereas for models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given.
+In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
 Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively.
@@
 Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
-We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
+We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
 At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
@@
 It follows that the set of matrices $B_{1}$ that satisfy Equation (B.24) has Lebesgue measure zero on $\mathbb{R}^{d\times d}$, and therefore, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.24) has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$.
-By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.11) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$.
+By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.22) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
 Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.23) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
@@
 Thus, there exists at most one solution to this system of linear equations.
-Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.∎
+Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified.
+As already stated, the transition weights weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$ for logistic models and almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ for models with nonrandom exogenous weight, concluding the proof.∎
 B.5 Proof of Theorem A.1
@@
 Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
-B.6 Details on the empirical application
-The ML estimation of the models, residual diagnostics, and estimation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2024) in which the introduced methods have been implemented.
-The R package sstvars also contains the dataset studied in the empirical application to facilitate reproduction of our results.
-In the ML estimation, we consider only solutions that can be deemed local maximums of the log-likelihood function.
-In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because identification with respect to the ordering of the columns of $B_{2}$ is weak (see Section 3.5).
-One the other hand, changing the ordering of the columns of $B_{y,t}$ changes the ordering of the variables in the density function $t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})$, thus, changing the ordering of the residuals in the given regime.
-The residual time series can, therefore, be labelled only for the variables whose structural shocks have been labelled.
-Since we labelled the CPU shock as the first shock in Section 5.3, the first residual series constitutes the residuals of CPUI.
-Since the other structural shocks were not labelled, the other residual series cannot be labelled for any of the variables.
-To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the residuals for the first $24$ lags in Figure B.1.
-As the figure shows, there is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the first lag of the CCFs between the fourth and fifth residual series.
-There are also some moderate sized CCFs in the ACFs of the third residual series, but these are not particularly large.
-Since rest of the CCs are reasonable, our LSTVAR model seems to capture the autocorrelation structure of the data reasonably well.
-In order to study the model’s adequacy to capture the conditional heteroskedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure B.2.
-The figure shows that there are several moderate and relative large CCs in the ACFs and CCFs of some of the residual series.
-Thus, there is some but not particularly severe inadequacies in terms of capturing the conditional heteroskedasticity.
-The series of standardized residuals, presented in the top panels of Figure B.3, also show some remaining heteroskedasticity and several outliers, many of which are accommodated by the heavy tailed $t$-distributions.
-Finally, the quantile-quantile-plots based on the independent $t$-distributions with mean zero, variance one, and degrees-of-freedom given by the degrees-of-freedom parameter estimate are presented in the bottom panel of Figure B.3.
-They show that the marginal distributions of the series are relatively well captured, although the distributions of two of the residual series are slightly asymmetric.
-Nonetheless, in our view, the overall adequacy of the model seems reasonable enough.
-Figure B.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$.
+Appendix C Penalized nonlinear least squares estimation of STVAR models
+This appendix describes how the parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of nonlinear least squares (NLS) as explained in Hubrich and Teräsvirta (2013, Section 4.2) for logistic STVAR models.
+After describing the NLS estimation, penalized nonlinear least squares (PNLS) estimation is considered by introducing a penalization term to the sum of squares of residuals function.
+C.1 Nonlinear least squares estimation
+Consider the following sum of squares of residuals function
+$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (C.1)
+where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters.
+[^3] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$.
+If the transition weights are exogenous, we simply treat them as known constants.
+The NLS estimates of the parameters $\phi_{m,0},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
+If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ they depend on needs to be estimated.
+Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only.
+However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that not able escape from local maxima are not sufficient.
+Since we only need approximate estimates to serve as starting values for Phase 2 and 3 estimation in the three-phase procedure described in Appendix 4.3, the following procedure is adopted.
+First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model.
+As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model and its parameters are not identified.
+Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
+One way to asses the contribution of each regime is by summing over the transition weights for each regime, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$.
+In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime.
+Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$.
+Denote the large enough sum as $\tilde{T}$.
+We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough.
+[^4] Our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
+Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range.
+For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches.
+With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$.
+Hence, it is advisable to select a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time.
+Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$.
+Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$.
+Denote this subset as $\mathcal{W}$.
+Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1,0},....,\phi_{M,0}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$.
+Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
+C.2 Penalized nonlinear least squares estimation
+Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (C.1), the penalized sum of squares function is minimized.
+We define the penalized sum of squares function as
+$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (C.2)
+where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space.
+Here we are interested only in avoiding estimates that do not satisfy the usual stability condition for all the regimes.
+Therefore, we define the penalty term as
+$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (C.3)
+where
+$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (C.4)
+is the minimized residual sum of squares, $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization.
+Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues.
+The penalty term (C.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
+The penalized NLS estimates of $\phi_{m,0},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2).
+In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
+Appendix D A Monte Carlo study
+This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-phase estimation procedure discussed in Section 4.
+We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable.
+To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$.
+For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
+For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1.
+In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1).
+In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region.
+To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus).
+The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specification (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
+Regime 1 Regime 2
+$\phi_{1,0}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2,0}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
+[…]
+[…]
+[…]
+[…]
+Table D.1: The parameters values of $\phi_{1,0},\phi_{2,0},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study.
+Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m,0}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
+For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$.
+Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-phase estimation procedure based on the penalized log-likelihood function, proposed in Section 4.
+After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values.
+To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
+$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
+TVAR 1 $\phi_{1,0}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$0.00$ $(0.07)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$\phi_{2,0}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
+$0.01$ $(0.28)$ $0.00$ $(0.19)$ $0.00$ $(0.13)$ $0.00$ $(0.09)$ $0.00$ $(0.04)$
+$A_{1,1}$ $-0.01$ $(0.08)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$A_{2,1}$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$
+$0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$B_{1}$ $0.20$ $(0.52)$ $0.09$ $(0.33)$ $0.05$ $(0.21)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
+$-0.08$ $(0.34)$ $-0.04$ $(0.19)$ $-0.02$ $(0.11)$ $-0.01$ $(0.06)$ $0.00$ $(0.02)$
+$-0.02$ $(0.09)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.03$ $(0.10)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$B_{2}$ $0.30$ $(0.61)$ $0.14$ $(0.39)$ $0.07$ $(0.24)$ $0.03$ $(0.13)$ $0.00$ $(0.04)$
+$0.04$ $(0.15)$ $0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+$-0.03$ $(0.18)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$r_{1}$ $0.00$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$ $0.00$ $(0.00)$
+$\nu_{1}$ $0.05$ $(0.53)$ $0.02$ $(0.34)$ $0.01$ $(0.23)$ $0.00$ $(0.16)$ $0.00$ $(0.07)$
+$\nu_{2}$ $165.03$ $(700.13)$ $105.25$ $(757.09)$ $13.98$ $(174.19)$ $1.48$ $(6.32)$ $0.16$ $(1.40)$
+$\lambda_{1}$ $-0.03$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$\lambda_{2}$ $0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+TVAR 2 $\phi_{1,0}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
+$-0.05$ $(0.14)$ $-0.04$ $(0.11)$ $-0.02$ $(0.07)$ $-0.02$ $(0.04)$ $-0.01$ $(0.02)$
+$\phi_{2,0}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
+$-0.01$ $(0.30)$ $0.00$ $(0.18)$ $0.01$ $(0.12)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$
+$A_{1,1}$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.00)$
+$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.03$ $(0.06)$ $-0.02$ $(0.04)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
+$A_{2,1}$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$0.00$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.02$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$B_{1}$ $0.38$ $(0.64)$ $0.16$ $(0.40)$ $0.08$ $(0.24)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
+$-0.15$ $(0.43)$ $-0.07$ $(0.25)$ $-0.04$ $(0.15)$ $-0.02$ $(0.08)$ $0.00$ $(0.03)$
+$-0.04$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.05$ $(0.18)$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
+$B_{2}$ $0.55$ $(0.73)$ $0.24$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
+$0.06$ $(0.17)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.13)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+$-0.05$ $(0.25)$ $-0.01$ $(0.08)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$r_{1}$ $0.08$ $(0.45)$ $0.02$ $(0.24)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+$\nu_{1}$ $0.05$ $(2.77)$ $-0.07$ $(0.29)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
+$\nu_{2}$ $107.17$ $(542.02)$ $93.32$ $(1273.27)$ $10.09$ $(81.02)$ $1.29$ $(5.10)$ $0.09$ $(1.47)$
+$\lambda_{1}$ $-0.04$ $(0.11)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
+$\lambda_{2}$ $-0.01$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4.
+The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
+The results are based on $500$ Monte Carlo repetitions.
+$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
+LSTVAR 1 $\phi_{1,0}$ $0.01$ $(0.17)$ $0.01$ $(0.08)$ $0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
+$0.00$ $(0.25)$ $0.00$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
+$\phi_{2,0}$ $-0.01$ $(0.17)$ $-0.01$ $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
+$-0.02$ $(0.39)$ $-0.02$ $(0.24)$ $-0.01$ $(0.15)$ $0.00$ $(0.11)$ $0.00$ $(0.05)$
+$A_{1,1}$ $0.00$ $(0.15)$ $0.01$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.02$ $(0.26)$ $-0.01$ $(0.13)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
+$0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.01$ $(0.10)$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$A_{2,1}$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$0.00$ $(0.14)$ $0.00$ $(0.09)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
+$-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$B_{1}$ $0.24$ $(0.60)$ $0.10$ $(0.37)$ $0.04$ $(0.20)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
+$-0.09$ $(0.46)$ $-0.05$ $(0.24)$ $-0.02$ $(0.13)$ $-0.01$ $(0.08)$ $0.00$ $(0.03)$
+$-0.03$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.05$ $(0.14)$ $-0.03$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$B_{2}$ $0.44$ $(0.72)$ $0.19$ $(0.42)$ $0.08$ $(0.22)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
+$0.05$ $(0.17)$ $0.02$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+$-0.03$ $(0.19)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$c$ $0.03$ $(0.19)$ $0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$\gamma$ $28.86$ $(121.66)$ $2.90$ $(25.52)$ $0.09$ $(0.75)$ $0.05$ $(0.47)$ $0.01$ $(0.19)$
+$\nu_{1}$ $-0.02$ $(0.49)$ $-0.02$ $(0.32)$ $-0.02$ $(0.21)$ $-0.01$ $(0.15)$ $0.00$ $(0.07)$
+$\nu_{2}$ $152.80$ $(597.75)$ $43.27$ $(266.76)$ $8.00$ $(52.02)$ $1.46$ $(5.55)$ $0.11$ $(1.37)$
+$\lambda_{1}$ $-0.04$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$\lambda_{2}$ $0.02$ $(0.10)$ $0.01$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+LSTVAR 2 $\phi_{1,0}$ $-0.01$ $(0.18)$ $-0.01$ $(0.11)$ $-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
+$-0.10$ $(0.19)$ $-0.06$ $(0.13)$ $-0.03$ $(0.08)$ $-0.02$ $(0.06)$ $-0.02$ $(0.02)$
+$\phi_{2,0}$ $0.07$ $(0.67)$ $0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
+$0.12$ $(1.18)$ $0.02$ $(0.19)$ $0.01$ $(0.13)$ $0.01$ $(0.08)$ $0.01$ $(0.04)$
+$A_{1,1}$ $-0.05$ $(0.13)$ $-0.02$ $(0.07)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $0.00$ $(0.01)$
+$-0.03$ $(0.11)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$0.03$ $(0.13)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$-0.03$ $(0.08)$ $-0.03$ $(0.06)$ $-0.02$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
+$A_{2,1}$ $-0.02$ $(0.11)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.02$ $(0.21)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.06)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+$-0.03$ $(0.13)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$B_{1}$ $0.44$ $(0.74)$ $0.16$ $(0.41)$ $0.08$ $(0.24)$ $0.03$ $(0.12)$ $0.01$ $(0.05)$
+$-0.14$ $(0.55)$ $-0.06$ $(0.27)$ $-0.04$ $(0.18)$ $-0.02$ $(0.09)$ $0.00$ $(0.03)$
+$-0.04$ $(0.13)$ $-0.02$ $(0.08)$ $-0.01$ $(0.06)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
+$-0.03$ $(0.19)$ $-0.02$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$
+$B_{2}$ $0.64$ $(0.84)$ $0.25$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
+$0.08$ $(0.19)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
+$-0.01$ $(0.09)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+$-0.04$ $(0.23)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$c$ $0.23$ $(0.80)$ $0.05$ $(0.31)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
+$\gamma$ $79.82$ $(349.54)$ $22.77$ $(90.88)$ $5.71$ $(44.05)$ $0.39$ $(4.28)$ $0.10$ $(0.30)$
+$\nu_{1}$ $-0.11$ $(0.57)$ $-0.07$ $(0.30)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
+$\nu_{2}$ $192.21$ $(889.55)$ $38.30$ $(233.50)$ $9.58$ $(79.39)$ $1.37$ $(4.71)$ $0.15$ $(1.42)$
+$\lambda_{1}$ $-0.05$ $(0.09)$ $-0.02$ $(0.06)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
+$\lambda_{2}$ $0.00$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4.
+The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
+The results are based on $500$ Monte Carlo repetitions.
+The results of the Monte Carlo study are presented in Table D.2 for the TVAR models and in Table D.3 for the LSTVAR models.
+The results show that there is some but mostly very small bias, which diminishes when the sample size increases.
+Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with consistency of the PML estimator.
+Nonetheless, for certain parameters, there appears to be a notable bias in small samples.
+First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large.
+This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution.
+Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values, and hence, possibly very large degrees-of-freedom parameter values drawn by random in Phase 2 estimation may remain large through Phase 3 estimation.
+Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the upward bias should not be a concern (for example, if only the Monte Carlo repetitions in which the estimate of $\nu_{2}$ is less than $100$ are considered, then for the specification TVAR 1, the estimated bias reduces to $5.37$ and the standard deviation to $17.63$ when $T=250$).
+Second, in the LSTVAR models, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small.
+A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are smooth in the true data generating process ($\gamma=5$).
+The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
+Third, somewhat surprisingly, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small.
+A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters.
+In particular, the estimates may be sensitive to extreme values drawn for the skewed $t$ shocks, potentially amplifying small-sample bias.
+In general, the small-sample bias of the PML estimator seems to be slightly larger the specifications where the true AR matrices are in the penalization region (TVAR 2 and LSTVAR 2) than the specifications where the true AR matrices lie well inside the stability region (TVAR 1 and LSTVAR 1).
+The differences are, however, only small, and even in small samples, there does not appear to be any particular issues with the estimation accuracy of the AR matrices lying inside the penalization region.
+Also, the TVAR models seem to be slightly more accurately estimated than the LSTVAR models, possibly because, unlike in TVAR models, in LSTVAR models also the smoothness of regime-switches needs to be estimated.
+Appendix E Details on the empirical application
+The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a) in which the introduced methods have been implemented.
+The R package sstvars also contains the dataset studied in the empirical application.
+In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because labelling them effectively requires labelling the corresponding structural shocks, and only the CPU shock has been labelled.
+To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure E.1.
+There is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the lag zero CCFs between the fourth and fifth residual series.
+In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small.
+Hence, with the exception of the single large CC in the lag zero CCFs of the fourth and fifth residual series, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
+To study our model’s adequacy to capture the conditional heteroscedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.2.
+There is a very large CC in the lag zero CCFs between the second and third residual series, which, however, disappears if the COVID-19 residual from March 2020 is excluded (not shown).
+In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series.
+Thus, there are some inadequacies in capturing the conditional heteroscedasticity, but with the exception of the large CC in CCFs between the second and third residual series (due to the extreme COVID-19 volatility), the inadequacies do not seem particularly severe.
+Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates are presented in Figure E.3 for each residual series.
+They show that the marginal distributions of the series are reasonably well captured, although there is a single large outlier in the third residual series (which disappears if the March 2020 residual is removed; not shown).
+Therefore, the overall adequacy of the model seems, in our view, reasonable.
+Figure E.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$.
 The lag zero autocorrelation coefficients are omitted, as they are one by convention.
 The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
-Figure B.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$.
+Figure E.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$.
 The lag zero autocorrelation coefficients are omitted, as they are one by convention.
 The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
-Figure B.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime fifth-order LSTVAR model.
-The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees of freedom given by the ML estimate of the degrees of freedom parameter.
+Figure E.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime sixth-order LSTVAR model.
+The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates.

### v4

@@
 macroeconomy.
-In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
+In a structural logistic smooth transition vector autoregressive model with two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
 Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
@@
 1 Introduction
-Smooth transition vector autoregressive (STVAR) models (Anderson and Vahid, 1998) have been widely adopted in empirical macroeconomics due to their ability to flexibly capture nonlinear data generating dynamics and gradual shifts in parameter values.
-Such variation may arise due to wars, business cycle fluctuations, or policy shifts, for example.
-Structural STVAR models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock.
-In order to estimate their effects, the economic shocks need to be identified.
-Conventional identification methods typically rely on restrictive assumptions such as zero contemporaneous interactions among some of the variables or long-term neutrality of the shocks.
+Linear vector autoregressive (VAR) models are a standard tool in empirical macroeconomics due to their ability to effectively capture constant dynamic relationships of the included variables.
+Macroeconomic systems may, however, exhibit variation in their dynamics, induced by wars, crises, policy shifts, or business cycle fluctuations, for example.
+Such structural changes cannot be accommodated by linear VAR models, and therefore, nonlinear VAR models that are able capture such dynamics are often employed.
+Nonlinear structural VAR (SVAR) models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock (Kilian and Lütkepohl, 2017, Chapter 18).
+To estimate these effects, the economic shocks need to be identified.
+Conventional identification methods typically employed in nonlinear SVAR models rely on restrictive assumptions, such as zero contemporaneous interactions among some of the variables (see Kilian and Lütkepohl, 2017, Section 18.2.2).
 Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification.
@@
 There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others).
+To the best our knowledge, this paper is the first to study identification by non-Gaussianity in nonlinear SVAR models.
 Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions.
@@
 This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
-To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
 This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al.
-(2017) to STVAR models.
-We also consider threshold vector autoregressive models of Tsay (1998) as special cases of STVAR models.
-In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification.
+(2017) to smooth transition vector autoregressive (STVAR) models (e.g., Anderson and Vahid, 1998), which is major class of nonlinear SVAR models.
+We also consider the threshold vector autoregressive model of Tsay (1998) as a special case of the STVAR model.
+The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual shifts between them, governed by the transition weighs.
+In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification.
 Nonetheless, similarly to Lanne et al.
 (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian.
-We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs.
-The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well.
-In line with statistical identification literature, however, external information is required to labelling the identified structural shocks as economic shocks.
-Since the identification can, in our experience, be also weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional information (cf.
+We show that under this condition, when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs.
+The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be extended to other suitable weight functions as well.
+In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks.
+Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes.
+Therefore, since our experience shows that the identification can be weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional identifying information (cf.
 Carriero et al., 2024).
@@
 Moreover, Morioka et al.
-(2021) assume innovations follow a distribution in the exponential family, thereby excluding various alternatives.
+(2021) assume the innovations follow a distribution in the exponential family, thereby excluding various interesting alternatives.
 In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
-Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S.
+Our empirical application studies the macroeconomic effects of climate policy uncertainty (shocks) and considers monthly U.S.
 data from 1987:4 to 2024:12.
@@
 Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable.
-We find that a positive CPU shock decreases production and increases inflation in times both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
+We find that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
 Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@
 Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks.
-Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-phase procedure.
+Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-step procedure.
 Section 5 presents the empirical application and Section 6 concludes.
 Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application.
-Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
+Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
 2 Reduced form STVAR models
@@
 $y_{t}$ $=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu),$ (2.1)
-$\mu_{m,t}$ $=\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
-where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
+$\mu_{m,t}$ $=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
+where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
 The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$.
-Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
+Thus, the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below, are accommodated.
 The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
@@
 See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
-Our framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics.
+Our framework accommodates various types of transition weights employed in empirical macroeconomics.
 To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
 In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2).
-For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$.
+For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constant across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all $i=1,...,p$ and $M=1,...,M$.
 The model then simplifies to
-$y_{t}=\phi_{0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
-The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), so it is more feasible in large scale VAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
+$y_{t}=\phi+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
+The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), making it more feasible for large scale SVAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
 3 Structural STVAR models identified by non-Gaussianity
@@
 Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al.
-(2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian.
+(2017), our identifying restriction is that the shocks are mutually independent and at most one of them is Gaussian.
 Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017).
 We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$.
-In order to parametrize the model directly with the impact matrix, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
-$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
+Hence, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
+$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
 where $e_{t}$ are independent and identically distributed structural errors with identity covariance matrix and a distribution that may depend on the parameter $\nu$.
@@
 To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
-In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
+In practice, the parameter space should be simply restricted in estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
 Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
@@
 The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-If the process is completely in one of the regimes, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of Regime $\tilde{m}$.
-When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends both on the regime-specific covariance matrices and the cross terms.
+If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes as the conditional covariance matrix of Regime $m$.
+When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms.
 Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes.
@@
 3.2 Identification of the shocks by non-Gaussianity
-The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are a mutually independent and at most one of them is Gaussian.
+The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one of them is Gaussian.
 We also assume that $e_{t}$ is an IID sequence with zero mean and identity covariance matrix as is formally stated in the following assumption.
@@
 The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns.
-This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent.
 This lemma, which is partly similar to Proposition 1 in Lanne et al.
 (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
-Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) as well as the weight function parameters are identified.
+Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) and weight function parameters are identified.
 Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
@@
 Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$. If the following conditions hold:
+Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and that the following conditions hold:
   1. (1)
-    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter), and
+    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous (nonrandom), or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter),
   2. (2)
-    $(\phi_{m,0},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n,0},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
-the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-For threshold models (2.3), also the parameters $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-For logistic models (2.4), the parameters $c,\gamma$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
+    $(\phi_{m},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$, and
+  3. (3)
+    there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
+Then, for models with exogenous weights, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+For threshold models (2.3), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+For logistic models (2.4), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,2$, and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ defined in Section 2 is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
 Lemma 2 is proven in Appendix B.2.
@@
 The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1.
-Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, because in some special cases of degenerate data (e.g., if the regime does not switch), there is insufficient variation to distinguish the parameters.
+Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$), because the identification fails in some special cases that comprise a set with Lebesgue measure zero.
+Condition (3) of Lemma 2 ensures that the transition weights exhibit enough variation to distinguish between different parameter values.
+If there are only two regimes, this condition is satisfied if transition weights take at least two distinct values.
 If the impact matrix is time-invariant as in Lanne et al.
@@
 However, when the impact matrix varies in time, two complications arise.
-First, the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, and it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified.
+First, as the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified.
 Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
 To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-To fix the ordering and signs of the columns of $B_{y,t}$, it suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$.
-This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
-It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending on the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others.
-Also, constraints that are identifying in some specifications may not be sufficient for identification in others.
+To fix the ordering and signs of the columns of $B_{y,t}$, such constraints need to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{1},...,B_{M}$ are observationally distinct at some $t$.
+However, constraints that are not restrictive in some specifications may be overidentifying in others, while constraints sufficient in some specifications may not be identifying in others, depending on the transition weights and shock distributions.
 Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
 3.3 Identification of the shocks in structural TVAR models
-Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$.
-Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$.
-That is, we have a structural TVAR model.
-By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained with constraints that fix the ordering and signs of the columns the impact matrix in each regime.
-To state the identification result, a specific shock distribution is assumed here for simplicity.
-Namely, we assume that each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness in the data, nests the Gaussian distribution as a special case, and is discussed in detail in Section 4.111The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
-The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR models.
+Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$, specifically being either exogenous or of the threshold form (2.3).
+Then at every $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$.
+By Lemma 2, $B_{y,t}$ is thereby identified up to ordering and signs of its columns within each regime, and identification requires constraints that fix these orderings and signs for each regime-specific impact matrix $B_{1},...,B_{M}$.
+For simplicity, and to state the identification result under just-identifying constraints, we assume each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness, and nests the Gaussian distribution as a special case.
+The skewed $t$-distribution is described in Section 4 discussing estimation of the model parameters.
+[^1] The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
+The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR model and models with exogenous binary weights.
 Proposition 1.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied, $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$, and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively.
+Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 and the conditions of Lemma 2 satisfied; $\alpha_{m,t}$ either of the threshold form (2.3) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$; and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively.
 Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
     $\lambda_{i}\neq 0$ for all $i=1,...,d$.
-Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
+Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$.
 Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively.
-Any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal).
-If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering corresponding of the columns of $B_{2},...,B_{M}$.
-In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification.
-If Condition (c), in turn, fails for some $i\in\{1,...,d\}$, the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification.
-The result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
-If Condition (b) or (c) is close to break, the identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively.
-Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) are not close to break, the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4).
+Any fixed ordering and signs of the columns of $B_{1}$ can be chosen without loss of generality (e.g., by making the first nonzero entry in each column is positive and ordering these entries in decreasing magnitude, given that none of them are equal).
+If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering of the corresponding columns of $B_{2},...,B_{M}$.
+In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality.
+Similarly, if Condition (c) fails for some $i\in\{1,...,d\}$ (meaning zero skewness), the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality.
+The result of Proposition 1 holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$, because in some special cases that comprise set with a Lebesgue measure zero, changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
+If Condition (b) or (c) is close to break, identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively.
+Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) hold clearly, identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4).
 Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
@@
 3.4 Identification of the shocks in structural STVAR models
-Having established the identification result for structural TVAR models, we now consider the transition weights that exhibit more variation.
-The transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited.
-Because the variation in the transition weights translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns by Lemma 2, the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1.
-Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
+Having established the identification result for structural TVAR models, we now consider transition weights that exhibit more variation.
+In particular, the transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited.
+Under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
 Proposition 2.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed).
+Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed).
 Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
   1. (A)
-    the ordering and signs of the columns of $B_{1}$ are fixed,
+    the ordering and signs of the columns of $B_{1}$ are fixed, and
   2. (B)
-    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero, and
-  3. (C)
-    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values.
-Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
+    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
+Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
 For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
 Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified.
-The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as there may exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
-For logistic models, the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, as special cases of degenerate data may lack sufficient variation to uniquely identify the transition weights (see Lemma 2).
+The result holds only almost everywhere, because the parameters are not identified in some special cases that constitute a set with Lebesgue measure zero (see also Lemma 2).
 Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$.
-If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related skewness parameter.
+If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter.
 Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes.
-If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be more relevant.
-Condition (C) of Proposition 2 states that the transition weights should exhibit at least a certain small degree of variation.
-For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$.
-This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete).
-Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
+If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be relevant.
+Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
 3.5 Labelling the shocks and blended identification
@@
 This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
-Additionally, in line with the statistical identification literature, many economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
 Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes.
 For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs.
-However, as discussed in Sections 3.3 and 3.4, the identification is often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
+However, as discussed in Sections 3.3 and 3.4, identification can be weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
 Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other.
-If the identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
-As our estimation method, discussed in Section 4, can produce a set of local solutions, the possible presence of such local solutions can be easily checked.
+If identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
 To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf.
@@
 Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum.
-To overcome this issue, we propose allowing for unstable estimates, but adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close the boundary of the stability region, and then maximizing the penalized log-likelihood function.
-In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance of the optimization algorithms, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
+To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region.
+In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
 Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024).
@@
 where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise.
-The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is Gamma function.
+The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is the Gamma function.
 Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
@@
 where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3).
-The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{M,0},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters.
-[^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights $\alpha$ is omitted from the parameter vector.
+The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degree-of-freedom and skewness parameters.
+[^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
 The penalized likelihood function is then defined as
@@
 where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space.
-Focusing on avoiding unstable estimates, we define the penalization term as
+To focus on avoiding unstable estimates, we define the penalization term as
 $P(\boldsymbol{\theta})=\kappa Td\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (4.4)
@@
 Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues.
-Our penalization term incorporates the multiplicative coefficient $Td$ in attempt to standardize the strength of the penalization with respect to the number of observations $T$ and variables $d$, as the value of the log-likelihood function (4.2) tends to scale with these quantities.
+Our penalization term incorporates the multiplicative coefficient $Td$ in an attempt to standardize the strength of penalization with respect to the number of observations $T$ and variables $d$, as the magnitude of the log-likelihood typically increases with these quantities.
 In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
-4.3 Three-phase estimation procedure
+4.3 Three-step estimation procedure
 Lanne et al.
-(2017) propose a three-phase estimation procedure for their linear SVAR model identified by non-Gaussianity.
-In the first phase, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent phases to maximize the log-likelihood function.
+(2017) propose a three-step estimation procedure for their linear SVAR model identified by non-Gaussianity.
+In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function.
 The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function.
-To address these challenges, we propose a modified version of the three-phase estimation procedure in which the multimodality is explicitly taken into account.
-Our three-phase estimation procedure operates with the following steps:
+To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
+Our three-step estimation procedure proceeds with the following steps:
   1. 1.
-    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1,0},...,\phi_{M,0},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) as described in Appendix C.
+    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) (see Appendix C for details).
 Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
 (2017) but with the nonlinearity of the model explicitly accounted for.
-Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-phase procedure.
+Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-step procedure.
 Step 2 is comparable to Step 2 of Lanne et al.
-(2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able escape from local maxima.
+(2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima.
 Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
 Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results.
-Since multiple estimation rounds are ran, the estimation procedure produces a set of estimates, possibly corresponding to different local maximums of the (penalized) log-likelihood function.
-If there are multiple solutions with (penalized) log-likelihoods almost equal to the highest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest.
-The proposed three-phase estimation procedure is implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
-To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-phase estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D.
+Since multiple estimation rounds are run, the estimation procedure produces a set of estimates, presumably corresponding to different local maximums of the (penalized) log-likelihood function.
+If there are multiple solutions with (penalized) log-likelihoods almost equal to the greatest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest.
+The proposed three-step estimation procedure is implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
+To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D.
 Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$.
 To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
-According to the results (Tables D.2 and D.3 in Appendix D), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias.
+According to the results (see Tables D.2 and D.3 in Appendix D for details), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias.
 Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models.
-While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for a degrees-of-freedom parameter, and for a certain entry in each impact matrix, explainable by the statistical properties of the true data generating process.
-However, the bias diminishes for all of the parameters as the sample size increases.
+While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for the degrees-of-freedom parameter whose true value is not small, and for the first entry in each impact matrix.
+The upward biases in the scale and not-small degrees-of-freedom parameters can both be explained by weak discrimination between large and very large estimates, as large values already imply nearly discrete regime-switches and near-Gaussian distributions, respectively.
+The upward bias in the first entry of each impact matrix, in turn, could be explained by its sensitivity to extreme values induced by the related particularly heavy tailed skewed $t$-distribution.
+Nonetheless, the bias diminishes for all of the parameters as the sample size increases.
 Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator.
-Overall, the performance of the proposed PML estimator and the three-phase estimation procedure seems reasonable.
+Overall, the performance of the proposed PML estimator and the three-step estimation procedure seems reasonable.
 5 Empirical application
-Our empirical application studies the macroeconomic effects of the climate policy uncertainty (CPU) shock.
+Our empirical application studies the macroeconomic effects of climate policy uncertainty (CPU).
 In the related literature, Fried et al.
 (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output.
-Khalil and Strobel (2023) reach similar conclusions in their DSGE model, attributing the effects to financial institutions’ aversion to uncertainty.
+Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty.
 Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
-We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU).
+We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU).
 Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable.
@@
 Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE).
-The series of the included variables are presented in Figure 1.
+The series of the included variables are presented in Figure 1.333The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, PPI, and the Federal funds rate are retrieved from the Federal reserve bank of St.
+Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
 Figure 1: Monthly U.S.
@@
 From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods.
-The bottom panel shows the transition weights of High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model.
+The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model.
 The shaded areas indicate the U.S.
@@
 We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4).
-They facilitate smooth transitions between the regimes depending the level of the switching variable, enabling an analysis of how the effects of the CPU shock vary depending on the level of EPU by specifying its first lag as the switching variable.
-The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating selection between these two weight functions within estimation.
-We estimate the structural LSTVAR model by PML with the three-phase procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions.
-The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case of the skewed $t$-distribution, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters.
-To avoid overfitting, particularly in the regime prevailing when the level of EPU is high, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
-To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that avoids making restrictive assumptions about the effects of the CPU shock.
+They facilitate smooth transitions between the regimes and enable the analysis of how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable.
+The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating the selection between these two weight functions within estimation.
+We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions.
+The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters.
+To avoid overfitting, particularly in the regime accommodating fewer observations, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
+To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that largely avoids making restrictive assumptions about the effects of the CPU shock.
 We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes.
-Then, to support our identification, we check that the recovered CPU shocks align with several major historical events affecting CPU.
+Then, to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU.
 In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively).
@@
 Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime.
-The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1 where the fitted transition weights of High EPU Regime are presented.
+The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1, where the fitted transition weights of High EPU Regime are presented.
 The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively.
 Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied.
-Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes using the JSR toolbox (Jungers, 2023) in MATLAB.
-The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1.
+Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes.
+[^4] We compute the upper bound using the JSR toolbox (Jungers, 2023) in MATLAB.
+The obtained upper bound is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1.
 Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
@@
 To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives.
-Specifically, for each Regime $\tilde{m}\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{\tilde{m},t}$ is greater than $0.75$, indicating that Regime $\tilde{m}$ is clearly dominant (there are $366$ such histories for Low EPU Regime and $80$ for High EPU Regime).
+Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant (there are $366$ such histories for the Low EPU Regime and $80$ for the High EPU Regime).
 For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B.
@@
 Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead.
-From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of High EPU Regime are depicted in each row.
+From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of the High EPU Regime are depicted in each row.
 The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels.
-The left column shows the responses to conditionally on High EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on High EPU Regime dominating.
+The left column shows the responses to conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating.
 All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points.
 Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
-The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRFs using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf.
+The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf.
 Lanne and Virolainen, 2025).
-The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in Low EPU Regime and decreases in High EPU Regime.
-A possible explanation for the decrease in the EPUI could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
+The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in the Low EPU Regime and decreases in the High EPU Regime.
+A possible explanation for the decrease in the EPUI in the High EPU Regime could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
 Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes.
@@
 (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital.
-Although many of the GIRFs exhibit a short-term increase of IPI in High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
-In both regimes, a positive CPU shock increases consumer prices, aligning with Huang and Punzi (2024) and Khalil and Strobel (2023).
-Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a positive supply shock, whereas according Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup.
-However, we find that the inflationary effects are stronger in High EPU Regime than in Low EPU Regime.
-One possible explanation is that the firms are more risk averse in the periods of high EPU, thereby reacting more sensitively and passing the increased costs or markups onto consumers.
-Producer prices, in turn, mostly decrease in Low EPU Regime but increase in High EPU Regime.
-The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases both regimes, but clearly more so in High EPU Regime.
-Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU.
+Although many of the GIRFs exhibit a short-term increase of IPI in the High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
+In both regimes, a positive CPU shock increases consumer prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023).
+Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup.
+However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime.
+One possible explanation is that the firms are more risk averse in the periods of high EPU, and thereby react more sensitively.
+Producer prices, in turn, mostly decrease in the Low EPU Regime but increase in the High EPU Regime.
+The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime.
+Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of the High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU.
 Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
@@
 Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
-In addition to discussing the identification of the shocks, we discussed the problem of labelling them, and how to utilize blended identification (cf.
-Carriero et al., 2024) to solve the problem of weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices.
-We discussed estimation of the model parameters by the method of penalized maximum likelihood, and proposed a three-phase estimation procedure by adapting the procedure of Lanne et al.
+In addition to establishing the identification of the shocks, we have discussed the problem of labelling them, and employed blended identification (cf.
+Carriero et al., 2024) to address weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices.
+We have proposed estimating the model parameters by the method of penalized maximum likelihood, and introduced a three-step estimation procedure by adapting the approach of Lanne et al.
 (2017) to nonlinear SVAR models.
-By making use of the results of Saikkonen (2008), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model.
-While useful for obtaining a well-behaving estimator, stationarity is not, however, required for the applicability of our identification results.
-The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
-Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S.
-data from 1987:4 to 2024:12.
+Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model.
+The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
+In an empirical application to U.S.
+data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model.
 As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
-We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable.
-We found that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
-Our results are, therefore, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
+To distinguish between periods low and high economic policy uncertainty, we used the first lag of the EPU index (Baker et al., 2016) as the switching variable.
+Our results are in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
+However, we found that while a positive CPU shock decreases production and increases inflation in times of both low and high EPU, its inflationary effects are stronger during the periods of high EPU.
 Appendix A Stationarity and ergodicity
-Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties for the estimator, such as consistency and a conventional limiting distribution.
-Saikkonen (2008) derives a sufficient condition for ergodicity and stationarity of a smooth transition vector error correction (STVEC) model.
+Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator.
+Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model.
 Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008).
 However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes.
-As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model cannot be directly obtained from the model of Saikkonen (2008).
-The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (although the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights).
+As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008).
+The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, but we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights).
 Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
@@
 $\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (A.2)
-The following assumption collects sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
+The following assumption collects the sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
 Assumption A.1.
@@
 Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008).
-Note that this condition is sufficient but not (necessarily) necessary for ergodic stationarity of the process.
+Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the process (to the best of our knowledge, there does not currently exist results presenting a sufficient necessary condition for STVAR models).
 Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example.
-Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied.
-This condition rules out nonrandom exogenous weights, which are perfectly fine in terms of our main identification results discussed in Section 3.
+Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied, ruling out nonrandom exogenous weights (which are accommodated by our identification results).
 The following theorem, which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
@@
 A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere.
-To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a zero Lebesgue measure.
+To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a Lebesgue measure zero.
 Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013).
-This makes it poorly suitable for restricting the parameter space in the numerical maximum likelihood estimation of parameters discussed in Section 4.
-During estimation, this condition would have to be verified repeatedly a very large of times, thus, substantially increasing the computational cost of the estimation, making it extremely tedious.
-Therefore, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
+This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4.
+During estimation, this condition would have to be verified repeatedly a very large of times, making the estimation extremely tedious.
+Hence, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
 Condition A.1.
@@
 It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein).
-Since our estimation method discussed in Section 4 can produce a number of local solutions, the researcher can choose the best appropriate local solution whose stationarity can be verified by Assumption A.1(I).
+After estimation, Assumption A.1(I) can be checked for the solutions of interest.
 Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013).
-The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required.
+The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a tight bound is required.
 The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@
 B.1 Proof of Lemma 1
-Before proving Lemma 1, we establish the following lemma that made use of in the proof:
+Before proving Lemma 1, we establish the following lemma that is made use of in the proof:
 Lemma B.1.
-Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variable $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
+Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
 $P(x_{1},...,x_{k})=0$ (B.1)
@@
 Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible.
-The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
+The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
 Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$.
@@
 The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al.
-(2017) making use of results from independent component analysis (Comon, 1994).
+(2017) building on results from independent component analysis (Comon, 1994).
 First, note that from Equation (3.2), we can rewrite the model as
-$y_{t}-\phi_{t}-\boldsymbol{A}_{t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
-where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
-Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{t}$ and $\tilde{\boldsymbol{A}}_{t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
+$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
+where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$, $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
+Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
 By subtracting this identify from (B.2), we obtain
-$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
+$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
 Taking conditional expectation conditional on $\mathcal{F}_{t-1}$ gives
-$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=0,$ (B.4)
+$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0,$ (B.4)
 as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (B.3), we get
@@
 Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian.
-Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most on column of $C$ may contain more than one nonzero element.
+Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most one column of $C$ may contain more than one nonzero element.
 Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$.
 Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem).
-This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$) so that $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction.
+This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$), so $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction.
 Hence, each column of $C$ has at most one nonzero element.
@@
 To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3).
-Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$.
-That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
-At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{0,m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$.
+Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$, $m=1,...,M$.
+By Condition (3) of Lemma 2, the sets $\mathcal{T}_{m}$ and $\tilde{\mathcal{T}}_{m}$ are nonempty for all $m=1,...,M$.
+That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
+At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$.
 Consequently, Equation (B.4) can be written as
-$\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
-where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$.
-Now, Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$.
+$\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
+where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$ at the time $t$.
+Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$.
 Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true.
-Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t},0}=\phi_{m_{t},0}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$.
-Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so that the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$.
+Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t}}=\phi_{m_{t}}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (as a countable union of measure zero sets has a measure zero).
+It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$.
+Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for a given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$.
 For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches.
 Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$.
-Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-This concludes that the parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-In other words, two distinct parameter sets give at most a measure-zero set of solutions for $(y_{1},...,y_{T})$.
+Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+This concludes that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
 Logistic weights.
 Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$.
-Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$.
-The proof proceed in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then we use these results to show that (3) the intercepts and autoregression matrices are uniquely identified.
-By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{t}=\phi_{t}$ and $\tilde{\boldsymbol{A}}_{t}=\boldsymbol{A}_{t}$ (irrespective of the weight function dynamics).
+Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$.
+The proof proceeds in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then (3) we use these results to show that the intercepts and autoregression matrices are uniquely identified.
+By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$ (irrespective of the weight function dynamics).
 By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
-$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\phi_{2,0})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1,0}-\tilde{\phi}_{2,0})=0$ (B.7)
+$\tilde{\phi}_{1}-\phi_{1}+\alpha_{2,t}(\phi_{1}-\phi_{2})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1}-\tilde{\phi}_{2})=0$ (B.7)
 and
@@
 respectively.
-Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in the (two) variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$.
-Therefore, for given coefficients, Equations (B.7) and (B.8) have at most one solution unless at least $d(1+pd)-1$ scalar equations are identically true.
-All the $d(1+pd)$ scalar equations are identically true only if $\phi_{1,0}=\phi_{2,0}$, $\tilde{\phi}_{1,0}=\tilde{\phi}_{2,0}$, $A_{1,i}=A_{2,i}$, and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$, which violates Condition (2) of Lemma 2.
-If exactly $d(1+pd)-1$ scalar equations are identically true, the coefficients for all but one of the scalar equations are zero, and $\alpha_{2,t}$ can be expressed in the form $\alpha_{2}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some constants $a_{0},a_{1}\in\mathbb{R}$.
-Since any single solution of $\alpha_{2,t},\tilde{\alpha}_{2,t}$ can be expressed in the form $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some $a_{0},a_{1}\in\mathbb{R}$, it follows that the solution to Equations (B.7) and (B.8) must satisfy $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ (for all $t$).
-Making the dependence of the transition weights of $z_{t}$ explicit in the notation, observe that in the identity $\alpha_{2,t}(z_{t})=a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$, the logistic functions $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are both continuous and strictly increasing maps of $z_{t}$ into $(0,1)$ (as $\gamma,\tilde{\gamma}>0$).
-Specifically, $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ increase smoothly from 00 to $1$ in an ”S”-shaped curve as $z_{t}$ goes from $\infty$- ∞ to $\infty$.
-Denote $h(z_{t})\equiv\alpha_{2,t}(z_{t})-a_{0}-a_{1}\tilde{\alpha}_{2,t}(z_{t})$, which is continuous in $z_{t}$ as $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are continuous.
-Suppose $h(z_{t})=0$ is not identically true, and suppose $a_{1}\neq 1$.
-Then, $a_{1}\tilde{\alpha}_{2,t}(z_{t})$ takes values from 00 to $a_{1}$.
-If $a_{1}=0$, then $h(z_{t})=0$ at most in a single point by strict monotonicity of $\alpha_{2,t}(z_{t})$.
-If $a_{1}<0$, then $h(z_{t})$ is strictly increasing, so $h(z_{t})=0$ at most in a single point.
-If $a_{1}\in(0,1)$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})<\tilde{\alpha}_{2,t}(z_{t})$, whereas if $a_{1}>1$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})>\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}\in\mathbb{R}$.
-In either scenario, $h(z_{t})$ can cross zero only a finite number of times, as otherwise repeated crossing would force repeated sign changes, clearly contradicting the smooth logistic shapes that are “similar” but scaled.
-On the other hand, $h(z_{t})=0$ cannot hold on any set of positive Lebesgue measure because $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are strictly increasing logistic functions, and two such functions that coincide on an interval must coincide on $\mathbb{R}$, thereby forcing $a_{0}=0$ and $a_{1}=1$.
-On the other hand, if $a_{1}=1$ and $a_{0}\neq 0$, then $\alpha_{2,t}(z_{t})$ and $a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$ differ by constant shifting, and they can intersect in at most finitely many points.
-Consequently, $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ only in sets of Lebesgue measure zero in $z_{t}\in\mathbb{R}$, implying that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ (and thereby also $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$) almost everywhere in $z_{t}\in\mathbb{R}$.
-Thus, also $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ and $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$ for all $t$ almost everywhere in $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in two variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$ (interpreting the AR parameters as coefficients here).
+Suppose first that all of them are identically true, i.e., that all the coefficients are zero.
+In (B.7), this requires that $\tilde{\phi}_{1}=\phi_{1}$, $\phi_{1}=\phi_{2}$, and $\tilde{\phi}_{1}=\tilde{\phi}_{2}$.
+Similarly in (B.8), zero coefficients require that $\tilde{A}_{1,i}=A_{1,i}$, $A_{1,i}=A_{2,i}$ and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$.
+But as $\phi_{1}=\phi_{2}$ and $A_{1,i}=A_{2,i}$ for all $i=1,...,p$ violates Condition (2) of Lemma 2, it must be that not all scalar equations in (B.7) and (B.8) are identically true.
+For each scalar equation that is not identically true, the coefficients are not all zero, and each of these scalar equations can be expressed in the form $a_{2,j}\alpha_{2,t}=a_{0,j}+a_{3,j}\tilde{\alpha}_{2,t}$ for some $a_{0,j},a_{2,j},a_{3,j}\in\mathbb{R}$ for the $j$th not identically true equation.
+If $a_{2,j}=0$, then $\tilde{\alpha}_{2,t}$ is the constant $-a_{0,j}/a_{3,j}$ for all $t$, which violates Condition (3) of Lemma 2.
+Hence, $a_{2,j}\neq 0$ (in the not identically true equations), implying that $\alpha_{2,t}=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}$, where $a_{1,j}\equiv a_{3,j}a_{2,j}^{-1}$.
+Thus, the solution to Equations (B.7) and (B.8) must satisfy this relationship (for all $j$).
+Making the dependence of the transition weights of $z_{t}$ explicit in the notation, suppose $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds in a set of positive Lebesque measure in $z_{t}\in\mathbb{R}$.
+Since $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are logistic, $\alpha_{2,t}(z_{t})$ and $a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ are real analytic functions.
+Thus, by the Identity theorem, we have that $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}$.
+Since this must hold for all $z_{t}$, it follows from $\underset{z_{t}\rightarrow-\infty}{\lim}\alpha_{2,t}(z_{t})=0$ and $\underset{z_{t}\rightarrow-\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}$ that $a_{0,j}=0$.
+Similarly, assuming $a_{1,j}\neq 0$, it follows from $a_{0,j}=0$, $\underset{z_{t}\rightarrow+\infty}{\lim}\alpha_{2,t}(z_{t})=1$, and $\underset{z_{t}\rightarrow+\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}+a_{1,j}$ that $a_{1,j}=1$.
+If $a_{1,j}=0$, then clearly $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds only in the limit $z_{t}\rightarrow-\infty$, and thereby not in a set of positive Lebesgue measure.
+Thus, it must be that $a_{0,j}=0$ and $a_{1,j}=1$ in sets of positive Lebesgue measure, concluding that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ almost everywhere in $z_{t}\in\mathbb{R}$.
 Next, we make use of the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ to show that $\tilde{c}=c$ and $\tilde{\gamma}=\gamma$.
@@
 If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$.
-Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $z_{t}\in\mathbb{R}$.
-We have established that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$, $\tilde{\gamma}=\gamma$, and $\tilde{c}=c$, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
-Consequently, Equations (B.7) and (B.8) simplify to
-$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\tilde{\phi}_{1,0})+\alpha_{2,t}(\tilde{\phi}_{2,0}-\phi_{2,0})=0$ (B.10)
+Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
+We have shown that for each $t$, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
+If the $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$ (rather than exogenous), we have that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ (as a countable union of measure zero sets has a measure zero).
+Since this holds for each $t$, combining it for $t=1,....,T$ yields that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+If $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$, this result holds almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+Before showing that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are identified, we present the following lemma that is used in the proof:
+Lemma B.2.
+Consider the $T$ linear equations in $M$ variables
+$\alpha_{1,t}x_{1}+...+\alpha_{M,t}x_{M}=c_{t},\quad t=1,...,T,$ (B.10)
+where $x_{1},...,x_{M}$ are scalar variables, and $\alpha_{1,t},...,\alpha_{M,t}$ and $c_{t}$ are scalar constants.
+Suppose there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
+Then, there exists at most one solution to the system of linear equations (B.10).
+Proof.
+By assumption, there exists $M$ indices $t_{1},...,t_{M}$ such that the coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
+Now, the rows of the $(M\times M)$ matrix
+$A^{\alpha}\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\\ \end{bmatrix}$ (B.11)
+are linearly independent, and thus $A^{\alpha}$ is invertible.
+The $M$ linear equations corresponding to these indices form the system $A^{\alpha}x=c$, where $x=(x_{1},...,x_{M})$ and $c=(c_{t_{1}},...,c_{t_{M}})$.
+Because $A^{\alpha}$ is invertible, this system has exactly one solution.
+Now, any solution $x$ of the full system (B.10) for $t=1,....,T$ must satisfy the above subsystem, which has exactly one solution. Therefore, the full system (B.10) has at most one solution. ∎
+We have established that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ (and $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ for all $t$) almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+Using these identities, Equations (B.7) and (B.8) simplify to
+$(1-\alpha_{2,t})\tilde{\phi}_{1}+\alpha_{2,t}\tilde{\phi}_{2}=(1-\alpha_{2,t})\phi_{1}+\alpha_{2,t}\phi_{2}$ (B.12)
 and
-$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-\tilde{A}_{1,i})+\alpha_{2,t}(\tilde{A}_{2,i}-A_{2,i})=0,\ \ i=1,...,p,$ (B.11)
+$(1-\alpha_{2,t})\tilde{A}_{1,i}+\alpha_{2,t}\tilde{A}_{2,i}=(1-\alpha_{2,t})A_{1,i}+\alpha_{2,t}A_{2,i},\ \ i=1,...,p,$ (B.13)
 respectively.
-Treating and $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$ as fixed constants, Equations (B.10) and (B.11) define a system of $d(1+pd)$ linear equations and $2d(1+pd)$ variables ($\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$).
-Hence, there exists at most one solution to this system linear equations.
-Since one solution is $\tilde{\phi}_{m,0}=\phi_{m,0},\tilde{A}_{m,1}=A_{m,1},...,\tilde{A}_{m,p}=A_{m,p}$, $m=1,2$, this is the unique solution.
-This concludes the proof that the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$ as well as the parameter $r_{1},...,r_{m-1}$ models with threshold and the parameters $c,\gamma$ for models with logistics transition weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as ($\mathcal{F}_{t-1}$-measurable or nonrandom) transition weights $\alpha_{2,t}$ as coefficients here, at each $t$, each scalar equation in (B.12)-(B.13) is a linear equation in two variables given by the corresponding entries of $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$.
+Each scalar equation in (B.12)-(B.13) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and two variables.
+By Condition (3) of Lemma 2 (with $M=2$), there exists two indices $t_{1},t_{2}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},\alpha_{2,t_{i}})$, or equally $(1-\alpha_{2,t_{i}},\alpha_{2,t_{i}})$, $i=1,2$, are linearly independent.
+Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td(1+pd)$ linear equations (B.12)-(B.13), $t=1,....,T$, in $2d(1+pd)$ variables.
+Since one solution is $\tilde{\phi}_{m}=\phi_{m},\tilde{A}_{m,i}=A_{m,i}$, $i=1,...,p$, $m=1,2$, this is the unique solution.
+This concludes the proof that the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
 Exogenous nonrandom weights.
 Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom).
-Then, arguments similar to those used for logistic weights, but applied for any finite number of regimes, show that the parameters $\phi_{m,0},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.∎
+If the transition weights are binary, i.e., $\alpha_{m,t}\in\{0,1\}$ for all $m=1,...,M$ and $t=1,...,T$, then arguments similar to those used for threshold weights can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+Next, suppose the transition weights are not binary.
+By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$.
+Substituting the exogenous weights ($\alpha_{m,t}=\tilde{\alpha}_{m,t}$ is given for all $m=1,...,M$ and $t=1,...,T$) to these equations yields
+$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})$ $=0,\quad\text{and}$ (B.14)
+$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{A}_{m,i}-A_{m,i})$ $=0,\quad i=1,...,p,$ (B.15)
+respectively.
+Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as transition weights $\alpha_{1,t},...,\alpha_{M,t}$ as constants, each scalar equations in (B.14)-(B.15) defined together for all $t=1,...,T$ defines a system of $T$ linear equations in $M$ variables.
+Therefore, arguments similar to those used for logistic weights but applied for any finite number of regimes regimes, making use of Lemma B.2, can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.∎
 B.3 Proof of Proposition 1
-We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ would lead to model that is not observationally equivalent with the original one.
+We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that models before and after reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ are not observationally equivalent.
 Specifically, we say that two models are observational equivalent if and only if their joint densities are equal.
@@
 Proof.
-Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$.
+Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$, which is not identically zero by $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$.
 It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable.
-It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is an one-dimensional manifold locally in a neighbourhood of $x^{0}$.
+It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is a one-dimensional manifold locally in a neighbourhood of $x^{0}$.
 Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds.
@@
 Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model.
-Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$.
+Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{1},...,y_{T})$.
 Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
@@
 $f_{y}(y_{t}-\mu_{y,t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i}),$ (B.22)
-where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices (identified by Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
+where $\mathcal{T}_{m}$, $m=1,...,M$, is the set of time indices (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and nonempty by Condition (3) of Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
 Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$.
@@
 Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$.
-Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
-$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.19)
+Therefore, as $\mathcal{T}_{1},...,\mathcal{T}_{M}$ and $\mu_{y,t}$, $t=1,...,T$ are identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ by Lemma 2 and the shocks and thereby $y_{t}$ follow continuous distributions, we have
+$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.23)
 and thus also
 $f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\boldsymbol{\tau}_{2}),...,B_{M}(\boldsymbol{\tau}_{M}))\neq f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\tilde{\boldsymbol{\tau}}_{2})...,B_{M}(\tilde{\boldsymbol{\tau}}_{M}))$ (B.24)
-almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$.
-Since this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.
+almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$.
+Since this result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.
 By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$.
@@
 Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed.
-Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$, i.e., to models that are not observationally equivalent.
-Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.∎
+Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different joint densities of the model almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, i.e., to models that are not observationally equivalent.
+Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.∎
 B.4 Proof of Proposition 2
-To proof consists of two parts.
-First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix.
+The proof consists of two parts.
+First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix.
 In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$.
 Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified.
-For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$ (where $z_{t}$ is the switching variable), whereas for models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
-By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+For models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
+By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},$ $y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
 With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given.
-In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
+In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
 Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively.
@@
 We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
+Suppose $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
 At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
-$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.21)
+$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.25)
 for some $(d\times d)$ diagonal matrix $D_{t}$ with ±$\pm 1$ diagonal elements and some $(d\times d)$ permutation matrix $P_{t}$ (that may vary in time). Rearranging the above equation gives
@@
 $\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))=0.$ (B.27)
-By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
+By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
 At each such time period $t$, Equation (B.27) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other.
@@
 Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.26) as $C_{t}\in\mathbb{R}^{d\times d}$.
-By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
+By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
 At each such time period $t$, the right side of Equation (B.26) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros.
@@
 Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.27) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
-Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
+This concludes that, almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, the matrix $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$
+Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
 Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$.
 By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$).
-Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A)), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$.
-Hence, reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ would lead to a model that is not observationally equivalent with the original one.
-The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A), implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$).
+Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A) of Proposition 2), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$.
+Hence, models before and after reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ are not be observationally equivalent.
+The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A) of Proposition 2, implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$).
 Thus, $B_{y,t}$ is uniquely identified at all $t$.
@@
 At each $t$, Equation (B.29) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$.
-The weights $\alpha_{m,t}$ are treated as constant multipliers here, as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here.
-If $\alpha_{m,t}=0$ for some regimes $m$, the corresponding variables in $\tilde{B}_{m}$ drop out of the system of equations, however.
-Under Condition (C), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values (plus one accounting for transition weights that equal to zero).
-Therefore, Equation (B.25) defined together for all $t$ defines a system of linear equations involving $Md^{2}$ variables and at least $Md^{2}$ linearly independent equations.
-Moreover, as $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values, each variable $\tilde{b}_{ij}^{(m)}$ appears in at least $M$ linearly independent the equations.
-Thus, there exists at most one solution to this system of linear equations.
-Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified.
-As already stated, the transition weights weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$ for logistic models and almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ for models with nonrandom exogenous weight, concluding the proof.∎
+The weights $\alpha_{m,t}$ are treated as constant coefficients here (some of which may be zero), as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here.
+Each scalar equation in (B.29) is a linear equation in $M$ variables.
+Thus, each scalar equation in (B.29) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and $M$ variables.
+By Condition (3) of Lemma 2 (assumed in Proposition 2), there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
+Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td^{2}$ linear equations (B.29), $t=1,....,T$, in $Md^{2}$ variables.
+Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified.
+For exogenous models, the identification of the transition weights is given.
+For logistic models, by Lemma 2, the transition weights weights are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$, concluding the proof.∎
 B.5 Proof of Theorem A.1
@@
 The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
-Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows.
+Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows ($z_{t}$ of Saikkonen (2008) should not be confused with the switching variable of this paper).
 First, we assume that the number of unit roots in the model is zero, so $n=r$ in Saikkonen (2008).
@@
 Using the notation of this paper, the first term on the right side of Equation (17) in Saikkonen (2008) then reduces to $\sum_{m=1}^{M}\alpha_{m,t}\sum_{i=1}^{p}A_{m,i}y_{t-i}$.
-Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ (with the notation of this paper).
+Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ (with the notation of this paper).
 That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), reduce to the first term on the right hand side of Equation (3.2) defining our structural STVAR model.
@@
 358) is satisfied.
-To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$, where $j$ is a positive integer, is bounded from above by a finite real number.
+To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$ is bounded from above by a finite real number, where $j$ is a positive integer, $\mathcal{A}=\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\}$, and $\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}$ as well as the set $\mathcal{A}^{j}$ are defined in Appendix A.
 Similarly to Saikkonen (2008, p. 313), clearly,
@@
 If the transition weights are exogenous, we simply treat them as known constants.
-The NLS estimates of the parameters $\phi_{m,0},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
-If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ they depend on needs to be estimated.
+The NLS estimates of the parameters $\phi_{m},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
+If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ (they depend on) needs to be estimated.
 Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only.
-However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that not able escape from local maxima are not sufficient.
-Since we only need approximate estimates to serve as starting values for Phase 2 and 3 estimation in the three-phase procedure described in Appendix 4.3, the following procedure is adopted.
+However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are not sufficient.
+Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Appendix 4.3, the following procedure is adopted.
 First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model.
-As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model and its parameters are not identified.
+As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model at all.
 Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
-One way to asses the contribution of each regime is by summing over the transition weights for each regime, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$.
+One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$.
 In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime.
@@
 We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough.
-[^4] Our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
+[^6] We acknowledge that our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
 Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range.
@@
 With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$.
-Hence, it is advisable to select a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time.
+Hence, we recommend selecting a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time.
 Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$.
@@
 Denote this subset as $\mathcal{W}$.
-Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1,0},....,\phi_{M,0}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$.
+Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1},....,\phi_{M}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$.
 Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
@@
 where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space.
-Here we are interested only in avoiding estimates that do not satisfy the usual stability condition for all the regimes.
-Therefore, we define the penalty term as
+To focus on avoiding estimates that do not satisfy the usual stability condition for all regimes, we define the penalty term as
 $P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (C.3)
@@
 The penalty term (C.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
-The penalized NLS estimates of $\phi_{m,0},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2).
+The penalized NLS estimates of $\phi_{m},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2).
 In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
 Appendix D A Monte Carlo study
-This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-phase estimation procedure discussed in Section 4.
+This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4.
 We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable.
@@
 In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region.
-To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus).
-The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specification (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
+To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus).
+The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
 Regime 1 Regime 2
-$\phi_{1,0}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2,0}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
+$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
 […]
@@
 […]
-Table D.1: The parameters values of $\phi_{1,0},\phi_{2,0},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study.
-Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m,0}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
+Table D.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study.
+Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
 For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$.
-Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-phase estimation procedure based on the penalized log-likelihood function, proposed in Section 4.
+Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4.
 After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values.
@@
 $T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
-TVAR 1 $\phi_{1,0}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+TVAR 1 $\phi_{1}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $0.00$ $(0.07)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$\phi_{2,0}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
+$\phi_{2}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
 $0.01$ $(0.28)$ $0.00$ $(0.19)$ $0.00$ $(0.13)$ $0.00$ $(0.09)$ $0.00$ $(0.04)$
@@
 $\lambda_{2}$ $0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-TVAR 2 $\phi_{1,0}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
+TVAR 2 $\phi_{1}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
 $-0.05$ $(0.14)$ $-0.04$ $(0.11)$ $-0.02$ $(0.07)$ $-0.02$ $(0.04)$ $-0.01$ $(0.02)$
-$\phi_{2,0}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
+$\phi_{2}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
 $-0.01$ $(0.30)$ $0.00$ $(0.18)$ $0.01$ $(0.12)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$
@@
 $\lambda_{2}$ $-0.01$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4.
+Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4.
 The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
@@
 $T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
-LSTVAR 1 $\phi_{1,0}$ $0.01$ $(0.17)$ $0.01$ $(0.08)$ $0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
+LSTVAR 1 $\phi_{1}$ $0.01$ $(0.17)$ $0.01$ $(0.08)$ $0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
 $0.00$ $(0.25)$ $0.00$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
-$\phi_{2,0}$ $-0.01$ $(0.17)$ $-0.01$ $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
+$\phi_{2}$ $-0.01$ $(0.17)$ $-0.01$ $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
 $-0.02$ $(0.39)$ $-0.02$ $(0.24)$ $-0.01$ $(0.15)$ $0.00$ $(0.11)$ $0.00$ $(0.05)$
@@
 $\lambda_{2}$ $0.02$ $(0.10)$ $0.01$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-LSTVAR 2 $\phi_{1,0}$ $-0.01$ $(0.18)$ $-0.01$ $(0.11)$ $-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
+LSTVAR 2 $\phi_{1}$ $-0.01$ $(0.18)$ $-0.01$ $(0.11)$ $-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
 $-0.10$ $(0.19)$ $-0.06$ $(0.13)$ $-0.03$ $(0.08)$ $-0.02$ $(0.06)$ $-0.02$ $(0.02)$
-$\phi_{2,0}$ $0.07$ $(0.67)$ $0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
+$\phi_{2}$ $0.07$ $(0.67)$ $0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
 $0.12$ $(1.18)$ $0.02$ $(0.19)$ $0.01$ $(0.13)$ $0.01$ $(0.08)$ $0.01$ $(0.04)$
@@
 $\lambda_{2}$ $0.00$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4.
+Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4.
 The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
@@
 This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution.
-Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values, and hence, possibly very large degrees-of-freedom parameter values drawn by random in Phase 2 estimation may remain large through Phase 3 estimation.
-Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the upward bias should not be a concern (for example, if only the Monte Carlo repetitions in which the estimate of $\nu_{2}$ is less than $100$ are considered, then for the specification TVAR 1, the estimated bias reduces to $5.37$ and the standard deviation to $17.63$ when $T=250$).
+Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values.
+Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation.
+Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the large upward bias should not be a concern.
 Second, in the LSTVAR models, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small.
@@
 The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
-Third, somewhat surprisingly, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small.
+Third, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small.
 A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters.
-In particular, the estimates may be sensitive to extreme values drawn for the skewed $t$ shocks, potentially amplifying small-sample bias.
+In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.
 In general, the small-sample bias of the PML estimator seems to be slightly larger the specifications where the true AR matrices are in the penalization region (TVAR 2 and LSTVAR 2) than the specifications where the true AR matrices lie well inside the stability region (TVAR 1 and LSTVAR 1).

### v5

@@
-Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models
-Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models
+Identification by non-Gaussianity in structural smooth transition vector autoregressive models
+Identification by non-Gaussianity in structural smooth transition vector autoregressive models
 Savi Virolainen
 University of Helsinki
-Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian.
-We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes.
-We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model.
-The introduced methods are implemented to the accompanying R package sstvars.
-Our empirical application studies the effects of the climate policy uncertainty shock on the U.S.
-macroeconomy.
-In a structural logistic smooth transition vector autoregressive model with two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
-Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
+We show that structural smooth transition vector autoregressive models are statistically identified if the shocks are mutually independent and at most one of them is Gaussian.
+This extends a known identification result for linear structural vector autoregressions to a time-varying impact matrix.
+We also propose an estimation method, show how a blended identification strategy can be adopted to address weak identification, and establish a sufficient condition for ergodic stationarity.
+The introduced methods are implemented in the accompanying R package sstvars.
+Our empirical application finds that a positive climate policy uncertainty shock reduces production and raises inflation under both low and high economic policy uncertainty, but its effects, particularly on inflation, are stronger during the latter.
+Keywords: smooth transition vector autoregression, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR, climate policy uncertainty
 The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P.
@@
 Box 17, FI–00014 University of Helsinki, Finland; e-mail: savi.virolainen@helsinki.fi.
-ORCiD ID: $0$-0002-5075-6821.
+ORCiD ID: 0000-0002-5075-6821.The author does not have conflicts of interest to declare.
 1 Introduction
-Linear vector autoregressive (VAR) models are a standard tool in empirical macroeconomics due to their ability to effectively capture constant dynamic relationships of the included variables.
-Macroeconomic systems may, however, exhibit variation in their dynamics, induced by wars, crises, policy shifts, or business cycle fluctuations, for example.
-Such structural changes cannot be accommodated by linear VAR models, and therefore, nonlinear VAR models that are able capture such dynamics are often employed.
-Nonlinear structural VAR (SVAR) models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock (Kilian and Lütkepohl, 2017, Chapter 18).
-To estimate these effects, the economic shocks need to be identified.
-Conventional identification methods typically employed in nonlinear SVAR models rely on restrictive assumptions, such as zero contemporaneous interactions among some of the variables (see Kilian and Lütkepohl, 2017, Section 18.2.2).
-Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification.
-To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
-There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others).
-To the best our knowledge, this paper is the first to study identification by non-Gaussianity in nonlinear SVAR models.
-Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions.
-However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017).
-This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
-This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al.
-(2017) to smooth transition vector autoregressive (STVAR) models (e.g., Anderson and Vahid, 1998), which is major class of nonlinear SVAR models.
-We also consider the threshold vector autoregressive model of Tsay (1998) as a special case of the STVAR model.
-The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual shifts between them, governed by the transition weighs.
+Linear structural vector autoregressive (SVAR) models have become a standard tool in empirical macroeconomics, as they can effectively capture constant dynamic relationships among the variables and facilitate tracing out the causal effects of economic shocks.
+Macroeconomic systems may, however, exhibit variation in their dynamics, induced by crises, policy shifts, or business cycle fluctuations, for example.
+Hence, also the effects of the shocks may depend on the state of the economy.
+Such features cannot be accommodated by linear SVAR models, and therefore nonlinear SVAR models that are able to capture such features are often employed (see, e.g., Kilian and Lütkepohl, 2017, Chapter 18).
+In nonlinear SVAR models, the structural shocks are typically identified with conventional methods that impose economically interpretable but restrictive assumptions, such as zero contemporaneous interactions among some of the variables, which are not always plausible in practice.
+To address this issue, statistical identification methods relying on the statistical properties of the data can be used (see Kilian and Lütkepohl, 2017, Chapter 14).
+There are two main branches in the statistical identification literature: identification by heteroskedasticity (Rigobon, 2003, Lanne and Lütkepohl, 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others).
+To the best of our knowledge, this paper is the first to examine identification by non-Gaussianity in nonlinear SVAR models.
+Under certain statistical conditions, both types of statistical information typically suffice to identify the shocks in a linear SVAR model.
+As pointed out by Lütkepohl and Netšunajev (2017), among others, identification by heteroskedasticity without additional restrictions, however, has the major drawback in nonlinear SVAR models that it imposes time-invariant (relative) impact effects of the shocks.
+We show that identification by non-Gaussianity, in turn, facilitates recovering the shocks without such undesirable restrictions.
+This paper contributes to the literature on identification by non-Gaussianity by extending the framework of Lanne et al.
+(2017) to structural smooth transition vector autoregressive (STVAR) models, which is a major class of nonlinear SVAR models (see, e.g., Hubrich and Teräsvirta, 2013).
+The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual as well as abrupt shifts between them, governed by the transition weights.
 In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification.
-Nonetheless, similarly to Lanne et al.
+Nevertheless, similarly to Lanne et al.
 (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian.
-We show that under this condition, when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs.
-The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be extended to other suitable weight functions as well.
+We show that under this condition, the shocks are readily identified up to ordering and signs when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes.
+While we focus on exogenous, logistic, and threshold transition weights, our results can be extended to other suitable weight functions as well.
 In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks.
 Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes.
-Therefore, since our experience shows that the identification can be weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional identifying information (cf.
+Moreover, our experience shows that identification is often weak in the sense that there are often multiple local solutions with a fit close to the global solution.
+We find such local solutions to be often largely quite similar to each other, with the differences frequently but not exclusively related to different ordering and signs of the columns of the regime-specific impact matrices.
+Nonetheless, since different local solutions may produce different results in structural analysis, we recommend addressing weak identification by adopting a blended identification strategy that combines identification by non-Gaussianity with supplementary identifying information (cf.
 Carriero et al., 2024).
-Developments related to ours include the new framework of Morioka et al.
-(2021) called independent innovation analysis (IIA).
-The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form.
-Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations.
-While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis.
-Moreover, Morioka et al.
-(2021) assume the innovations follow a distribution in the exponential family, thereby excluding various interesting alternatives.
-In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
-Our empirical application studies the macroeconomic effects of climate policy uncertainty (shocks) and considers monthly U.S.
+Our methods are demonstrated in an empirical application to the macroeconomic effects of climate policy uncertainty (shocks) that considers monthly U.S.
 data from 1987:4 to 2024:12.
-Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU.
+Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on topics related to CPU.
 We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU).
 Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable.
-We find that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU.
-Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
+We find that a positive CPU shock reduces production and raises inflation in times of both low and high EPU, but its effects, particularly on inflation, are stronger in the periods of high EPU.
+Our results are, hence, in line with the previous literature suggesting that a positive CPU shock reduces production and raises inflation (Khalil and Strobel, 2023 and Huang and Punzi, 2024), while Fried et al.
+(2022) found that it decreases production.
 The rest of this paper is organized as follows.
-Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models.
-Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks.
+Section 2 presents the framework of reduced form STVAR models.
+Section 3 discusses identification of the shocks in structural STVAR models, presents our identification results, and discusses the problem of labeling the shocks as well as how to address weak identification by adopting a blended identification strategy.
 Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood-based estimator using a three-step procedure.
 Section 5 presents the empirical application and Section 6 concludes.
-Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application.
-Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
+Further details can be found in the appendices, and the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
 2 Reduced form STVAR models
-Let $y_{t}$, $t=1,2,...$, be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$.
+Let $y_{t}$, $t=1,2,$..., be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$ (i.e., $\mathcal{F}_{t-1}$ contains the information about the history of the process).
 We consider STVAR models with $M$ regimes and autoregressive order $p$ assumed to satisfy
@@
 $\mu_{m,t}$ $=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
-where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
+where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, i=1,,pi=1,...,p, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
 The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$.
-Thus, the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below, are accommodated.
-The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
+Thus, we accommodate the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below.
+The transition weights express the proportions of the regimes the process is in at each point of time and determine how the process shifts between them.
 It is easy to see that, conditional on $\mathcal{F}_{t-1}$, the conditional mean of the above-described process is $\mu_{y,t}\equiv E[y_{t}|\mathcal{F}_{t-1}]=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$, a weighted sum the regime-specific means $\mu_{m,t}$ with the weights given by the transition weights $\alpha_{m,t}$.
-The conditional covariance matrix is $\Omega_{y,t}$, but discussion on its specification is postponed to Section 3.
-See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
-Our framework accommodates various types of transition weights employed in empirical macroeconomics.
-To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
-$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right.$ (2.3)
-where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable.
-At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables, for example.
-Another popular specification is to assume two regimes ($M=2$) and logistic transition weights.
+The discussion on the specification of the conditional covariance matrix $\Omega_{y,t}$ is postponed to Section 3.
+The linear vector autoregressive (VAR) model is obtained as a special case by assuming constant autoregressive dynamics across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all i=1,,pi=1,...,p and m=1,,Mm=1,...,M.
+[^1] See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
+A popular specification is obtained by assuming two regimes ($M=2$) and logistic transition weights.
 In the logistic STVAR model (Anderson and Vahid, 1998), the transition weights vary according to a logistic function as
 $\alpha_{1,t}=1-\alpha_{2,t}\ \ \text{and}\ \ \alpha_{2,t}=[1+\exp\{-\gamma(z_{t}-c)\}]^{-1},$ (2.3)
-where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter.
+where $c\in\mathbb{R}$ is the location parameter, $\gamma>0$ is the scale parameter, and $z_{t}$ is the switching variable, which we assume to be a lagged endogenous variable up to the lag $p$ (that is, $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$).
 The location parameter $c$ determines the midpoint of the transition function, i.e., the value of the switching variable when the weights are equal.
-The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
-In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2).
-For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constant across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all $i=1,...,p$ and $M=1,...,M$.
-The model then simplifies to
-$y_{t}=\phi+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
-The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), making it more feasible for large scale SVAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
+The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (the smaller $\gamma$ is, the smoother the transition is), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
+In the special case when $\gamma\rightarrow\infty$, regime-switches are discrete and the logistic weights reduce to the threshold weights of Tsay (1998) (see Appendix A for details).
 3 Structural STVAR models identified by non-Gaussianity
@@
 A structural STVAR model is obtained from the reduced form model defined in Section 2 by identifying the serially and mutually uncorrelated structural shocks $e_{t}=(e_{1t},...,e_{dt})$ $(d\times 1)$ from the reduced form innovations $u_{t}$.
-Specifically, the structural shocks are recovered from the reduced form shocks with the transformation
+Specifically, the structural shocks are recovered from the reduced form innovations with the transformation
 $e_{t}=B_{y,t}^{-1}u_{t},$ (3.1)
 where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$.
-In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$.
-The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$.
-Various types of restrictions that recover the structural shocks have been proposed in the literature, many of which are discussed in Kilian and Lütkepohl (2017), for example.
-Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al.
-(2017), our identifying restriction is that the shocks are mutually independent and at most one of them is Gaussian.
-Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017).
+In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$.
+However, $B_{y,t}$ is not generally uniquely identified without further restrictions, as many such decompositions exist.
+Various identification methods have been proposed to resolve this issue (see, e.g., Kilian and Lütkepohl, 2017).
+Lanne et al.
+(2017), among others, have shown that if the shocks are mutually independent and at most one of them is Gaussian, a static impact matrix is readily identified (up to ordering and signs of its columns) without additional restrictions.
 We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$.
@@
 To incorporate time-variation in the impact matrix, we specify its functional form.
-A natural specification for STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
+A natural specification for structural STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
 $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m},$ (3.3)
@@
 Also the impact matrix $B_{y,t}$ needs to be invertible to ensure positive definiteness of the conditional covariance matrix, which does not automatically follow from the invertibility of $B_{1},...,B_{M}$.
-Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix B.1).
+Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix C.1).
 Lemma 1.
-Suppose $B_{1},...,B_{M}$ are invertible $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and $m=1,...,M$.
-Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero.
-This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$).
-To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
-In practice, the parameter space should be simply restricted in estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
-Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
+Suppose $B_{1},...,B_{M}$ are $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and m=1,,Mm=1,...,M.
+Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+The result of Lemma 1 holds almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has Lebesgue measure zero.
+In other words, invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$).
+To support intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
+Assuming an impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
 $\Omega_{y,t}=\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}^{2}\Omega_{m}+\sum_{m=1}^{M}\sum_{n=1,n\neq m}^{M}\alpha_{m,t}\alpha_{n,t}\Omega_{m,n},$ (3.4)
 where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$.
-The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes as the conditional covariance matrix of Regime $m$.
+The conditional covariance matrix of $y_{t}$ can thus be described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M.
+If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes the conditional covariance matrix of Regime $m$.
 When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms.
 Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes.
-Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.
+Nevertheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification as is shown in the next section.
 3.2 Identification of the shocks by non-Gaussianity
@@
   1. (i)
-    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, follows a continuous distribution with zero mean and unit variance.
+    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, i=1,,di=1,...,d, follows a continuous distribution with zero mean, unit variance, and a density that is strictly positive almost everywhere in $\mathbb{R}$.
   2. (ii)
-    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one them has a Gaussian marginal distribution.
-Note that the assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$.
-Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
-The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns.
-This lemma, which is partly similar to Proposition 1 in Lanne et al.
-(2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified.
-Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) and weight function parameters are identified.
+    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one of them has a Gaussian marginal distribution.
+The assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$.
+Also, due to the time-variation of the impact matrix, Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
+The assumption of strictly positive density on sets of positive (Lebesgue) measure mainly rules out bounded shock distributions.
+It is helpful to distinguish two layers of identification.
+First, for any fixed $t$ and conditional on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ in (3.2) is identified up to column permutations and sign changes by independent component analysis (ICA) under Assumption 1 (cf.
+Lanne et al., 2017).
+Second, under the structural parametrization $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ in (3.3), identification of the regime-specific matrices $B_{1},...,B_{M}$ (up to column ordering and signs) additionally requires identification of the transition weight parameters, or in the case of exogenous weights, sufficient variation in the known weights.
+The next lemma, partly similar to Proposition 1 in Lanne et al.
+(2017), therefore formalizes the per-$t$ ICA result, and, in the case of logistic weights (2.3) establishes identification of the weight function parameters (and, for completeness, identification of the AR parameters for logistic and exogenous weights).
 Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
+Identification of $B_{1},...,B_{M}$ then follows from these ingredients.
 Lemma 2.
 Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied.
-Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and that the following conditions hold:
-  1. (1)
-    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous (nonrandom), or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter),
+Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Moreover, suppose the transition weights are either exogenous (nonrandom) or follow the logistic process defined in (2.3) (with $M=2$ assumed), and that the following conditions hold:
   1. (1)
@@
   2. (2)
-    there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
-Then, for models with exogenous weights, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-For threshold models (2.3), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-For logistic models (2.4), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,2$, and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ defined in Section 2 is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
-Lemma 2 is proven in Appendix B.2.
-We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1.
-In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model.
-The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1.
-Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$), because the identification fails in some special cases that comprise a set with Lebesgue measure zero.
-Condition (3) of Lemma 2 ensures that the transition weights exhibit enough variation to distinguish between different parameter values.
-If there are only two regimes, this condition is satisfied if transition weights take at least two distinct values.
-If the impact matrix is time-invariant as in Lanne et al.
-(2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs.
-Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017).
-However, when the impact matrix varies in time, two complications arise.
-First, as the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified.
-Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
-To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$.
-To fix the ordering and signs of the columns of $B_{y,t}$, such constraints need to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{1},...,B_{M}$ are observationally distinct at some $t$.
-However, constraints that are not restrictive in some specifications may be overidentifying in others, while constraints sufficient in some specifications may not be identifying in others, depending on the transition weights and shock distributions.
-Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
-3.3 Identification of the shocks in structural TVAR models
-Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$, specifically being either exogenous or of the threshold form (2.3).
-Then at every $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$.
-By Lemma 2, $B_{y,t}$ is thereby identified up to ordering and signs of its columns within each regime, and identification requires constraints that fix these orderings and signs for each regime-specific impact matrix $B_{1},...,B_{M}$.
-For simplicity, and to state the identification result under just-identifying constraints, we assume each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness, and nests the Gaussian distribution as a special case.
-The skewed $t$-distribution is described in Section 4 discussing estimation of the model parameters.
-[^1] The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
-The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR model and models with exogenous binary weights.
+    (for exogenous weights) there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, i=1,,Mi=1,...,M, are linearly independent.
+Then, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, m=1,,Mm=1,...,M, and for logistic models $c$ and $\gamma$, are uniquely identified.
+Lemma 2 is proven in Appendix C.2.
+Condition (1) implies distinguishable regimes, whereas Condition (2) guarantees a certain small amount of variation in exogenous weights.
+Lemma 2 concludes that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1.
+That is, at each tt‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model.
+It follows that if the impact matrix is time-invariant as in Lanne et al.
+(2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, the structural shocks are identified up to ordering and signs.
+However, when the impact matrix varies over time, two complications arise.
+First, because $B_{y,t}$ (3.3) is identified only up to column ordering and signs at each $t$, its unique identification requires constraints on the regime-specific impact matrices $B_{1},...,B_{M}$ such that any reordering or sign changes in their columns lead to observationally distinct $B_{y,t}$ at some $t$.
+Second, since $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in its functional form are identified.
+Also, in addition to $B_{1},...,B_{M}$, the impact matrix $B_{y,t}$ depends on the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, which must therefore be identified as well.
+To that end, we assume either the logistic or exogenous transition weights considered in Lemma 2.
+The following proposition, proven in Appendix C.3, establishes unique identification of $B_{1},...,B_{M}$.
 Proposition 1.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 and the conditions of Lemma 2 satisfied; $\alpha_{m,t}$ either of the threshold form (2.3) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$; and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively.
-Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
-  1. (a)
-    the ordering and signs of the columns of $B_{1}$ are fixed,
-  2. (b)
-    $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j\in\{1,...,d\}$,
-  3. (c)
-    $\lambda_{i}\neq 0$ for all $i=1,...,d$.
-Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$.
-Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively.
-Any fixed ordering and signs of the columns of $B_{1}$ can be chosen without loss of generality (e.g., by making the first nonzero entry in each column is positive and ordering these entries in decreasing magnitude, given that none of them are equal).
-If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering of the corresponding columns of $B_{2},...,B_{M}$.
-In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality.
-Similarly, if Condition (c) fails for some $i\in\{1,...,d\}$ (meaning zero skewness), the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality.
-The result of Proposition 1 holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$, because in some special cases that comprise set with a Lebesgue measure zero, changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
-If Condition (b) or (c) is close to break, identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively.
-Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) hold clearly, identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4).
-Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$.
-As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
-3.4 Identification of the shocks in structural STVAR models
-Having established the identification result for structural TVAR models, we now consider transition weights that exhibit more variation.
-In particular, the transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited.
-Under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
-Proposition 2.
-Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed).
-Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
+Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and suppose the transition weights are either exogenous (nonrandom) or follow the logistic process defined in (2.3) with $\gamma\in(0,\infty)$ (and $M=2$ assumed).
+Moreover, suppose the following conditions hold:
   1. (a)
@@
   2. (b)
-    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
-Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
-For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+    (for exogenous weights) the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
+Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
 Proposition 1 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified.
-The result holds only almost everywhere, because the parameters are not identified in some special cases that constitute a set with Lebesgue measure zero (see also Lemma 2).
-Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$.
-If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter.
-Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes.
-If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be relevant.
-Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
-3.5 Labelling the shocks and blended identification
+The result holds for almost every $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ because the identification may fail in some special cases, but this set has Lebesgue measure zero.
+[^2] Beyond possible singularity of the impact matrix $B_{y,t}$ (see Lemma 2), the failure of the identification in a measure zero set only concerns the possibility that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$.
+Condition (a) of Proposition 1 states that the ordering and signs of the columns of $B_{1}$ should be fixed (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal), which fixes the ordering and signs of the columns of $B_{y,t}$ for all $t$.
+If the shocks follow, for example, the skewed $t$-distributions defined in Equation (4.1) (in Section 4 discussing estimation), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality.
+This is because reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter.
+Condition (b) states that exogenous transition weights should, for some $t$, be strictly positive for all the regimes, which is used to establish that Condition (a) fixes the ordering and signs of the columns of $B_{y,t}$ for all $t$.
+For logistic weights, this is achieved via the assumption $\gamma<\infty$, guaranteeing that the regime-switches are not discrete.
+[^3] To accommodate discrete regime-switches, which are obtained as a special case of the logistic weights (2.3) with the smoothness parameter tending to infinity, we establish the identification of the shocks in the threshold VAR model of Tsay (1998) in Appendix A.
+Somewhat surprisingly, our experience shows that identification appears to be often weak in the sense that there are multiple local maxima of the (penalized) log-likelihood function (discussed in Section 4) with (penalized) log-likelihoods close to each other and to the global maximum.
+The parameter values related to each such local maximum seem to be often largely quite similar to each other, but have some differences, frequently but not exclusively related to different ordering and signs of the columns of $B_{2},...,B_{M}$.
+Since such different local solutions may produce different results in structural analysis, this weak identification should be appropriately addressed, for instance, by combining Proposition 1 with supplementary identifying information as discussed in the next section.
+3.3 Labeling the shocks and blended identification
 As in the linear SVAR model of Lanne et al.
-(2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information.
-Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables.
-This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
-Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes.
-For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs.
-However, as discussed in Sections 3.3 and 3.4, identification can be weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
-Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other.
-If identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
-To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf.
+(2017), the statistically identified structural shocks do not necessarily have economic interpretations, and labeling them as economic shocks requires external information.
+However, labeling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock must be associated with the same column of the impact matrix $B_{m}$ in all regimes.
+For example, if a positive supply shock should increase output and decrease prices on impact in all regimes, labeling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs.
+Moreover, as discussed in Section 3.2, our experience shows that identification is often weak in the sense that there are multiple local solutions with fit close to the global optimum.
+While we find such local solutions to be often largely quite similar to each other, they may produce different impulse response functions even when the shock of interest is associated with the same column of all $B_{1},...,B_{M}$.
+To formally address weak identification and the problem of labeling the shocks, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with supplementary identifying information (cf.
 Carriero et al., 2024).
-As a concrete example, if a supply shock should increase output and decrease prices at impact in all regimes, selecting the best local solution where the same column in all $B_{1},...,B_{M}$ satisfies such signs amounts to combining identification by non-Gaussianity with the appropriate short-run sign restrictions.
-In addition to short-run sign restrictions, also other types of information can be exploited in blended identification, including narrative restrictions and testable zero impact effect restrictions, for instance.
+Specifically, we propose imposing overidentifying restrictions that are sufficient to yield a unique local solution (among the ones with fit close to the global solution) and facilitate labeling the shocks of interest.
+Since different local solutions may yield different results, the restrictions should be economically reasonable and serve to exclude less plausible alternative solutions.
+As a simple example, in a bivariate system of output and prices, it may be reasonable to assume that for each variable, one of the shocks has greater impact effect on that variable than the other shock, and that this effect has the same sign across regimes.
+If one of the shocks is to be interpreted as a demand shock, it might be useful to further assume that the shock with the largest effect on output also moves prices in the same direction in all regimes.
+Similarly, if the other shock is intended to represent a supply shock, it could be useful to assume that it moves output and prices in the opposite directions in all regimes.444 Also other forms of information can be incorporated into the blended identification strategy, such as narrative restrictions and zero impact effect restrictions (the latter of which are testable due to statistical identification), for instance.
+See our empirical application in Section 5 for an example that involves more variables.
+As a practical consideration, note that our estimation method, described in Section 4 and implemented to the accompanying R package sstvars Virolainen (2025a), produces a set of alternative local solutions.
+By filtering these local solutions based on whether they satisfy the imposed overidentifying restrictions, it is straightforward to determine which restrictions are sufficient in the considered specification to exclude all but one of the local maximums (among the ones with a fit close to the presumed global optimum).
+Finally, to label the shocks of interest, the imposed restrictions can often be used to uniquely associate each of them to the same column of the impact matrix across all regimes.
 4 Estimation
 4.1 Maximum likelihood estimation and ergodic stationarity
-The parameters of the structural STVAR models discussed in Section 3 can be estimated by the method of maximum likelihood (ML), which amounts to maximizing the log-likelihood function over the parameter space.
+The parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of maximum likelihood (ML).
 To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary.
-Based on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one.
-But since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in estimation.
-Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition.
-After estimation, the sufficient condition can be checked for the solutions of interest.
-Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics.
-Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum.
+Building on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), it can be shown that when the transition weights $\alpha_{m,t}$ are logistic (2.3) or of the threshold form, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one (see Theorem B.1 in Appendix B).555 For logistic models (with $M=2$ assumed), we additionally require that the matrix $B_{1}^{-1}B_{2}$ has no negative real eigenvalues to rule out singular convex combinations of $B_{1}$ and $B_{2}$ (see Appendices B and C.5 for details).
+See Appendix A for the definition of the threshold weights.
+Because this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not particularly useful for restricting the parameter space during estimation.
+Therefore, we instead impose the usual stability condition for each regime (Condition B.1 in Appendix B), as it is a necessary condition for the sufficient one.
+The sufficient condition can then be checked after estimation for the solutions of interest.
+Maximizing the log-likelihood function can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics.
+[^6] The log-likelihood function is presented in Section 4.2 for shocks that follow skewed $t$-distributions.
+Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly.
 To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region.
-In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
-Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024).
-They impose nonnegativity constraints in the autoregressive conditional heteroscedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space.
+In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the algorithm back to the stability region.
+[^7] Penalized likelihood-based estimation has been previously applied to time series models by Nielsen and Rahbek (2024).
+They impose nonnegativity constraints in the autoregressive conditional heteroskedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space.
 Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
@@
 Before introducing the penalized log-likelihood function, the log-likelihood function is presented.
-We assume that each shock $e_{it}$, $i=1,...,d$, follow the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$.
+We assume that each shock $e_{it}$, i=1,,di=1,...,d, follows the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$.
+The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the plausibility of the identifying Assumption 1 can be (informally) assessed based on the estimates of these parameters.
 The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
@@
 $L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (4.2)
-where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3).
-The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degree-of-freedom and skewness parameters.
-[^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
-The penalized likelihood function is then defined as
+where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is the conditional mean defined in Section 2 and $B_{y,t}$ is the impact matrix defined in Equation (3.3).
+The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),$..., $\text{vec}(A_{m,p}))$, m=1,,Mm=1,...,M, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters.
+With logistic weights (2.3), $\alpha=(c,\gamma)$, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
+The penalized log-likelihood function is then defined as
 $PL(\boldsymbol{\theta})=L(\boldsymbol{\theta})-P(\boldsymbol{\theta}),$ (4.3)
@@
 In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function.
-The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function.
+The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a large number of modes to the (penalized) log-likelihood function.
 To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
@@
   1. 1.
-    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) (see Appendix C for details).
+    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by (penalized) nonlinear least squares (NLS) (see Appendix D for details).
 Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
 (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima.
-Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
-Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results.
-Since multiple estimation rounds are run, the estimation procedure produces a set of estimates, presumably corresponding to different local maximums of the (penalized) log-likelihood function.
-If there are multiple solutions with (penalized) log-likelihoods almost equal to the greatest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest.
-The proposed three-step estimation procedure is implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
-To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D.
-Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$.
-To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
-According to the results (see Tables D.2 and D.3 in Appendix D for details), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias.
-Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models.
-While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for the degrees-of-freedom parameter whose true value is not small, and for the first entry in each impact matrix.
-The upward biases in the scale and not-small degrees-of-freedom parameters can both be explained by weak discrimination between large and very large estimates, as large values already imply nearly discrete regime-switches and near-Gaussian distributions, respectively.
-The upward bias in the first entry of each impact matrix, in turn, could be explained by its sensitivity to extreme values induced by the related particularly heavy tailed skewed $t$-distribution.
-Nonetheless, the bias diminishes for all of the parameters as the sample size increases.
-Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator.
-Overall, the performance of the proposed PML estimator and the three-step estimation procedure seems reasonable.
+Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps.
+Due to the high multimodality of the (penalized) log-likelihood function, we recommend running Steps 2 and 3 a large number of times to improve the reliability of the results.
+Since a large number of estimation rounds are run, the estimation procedure produces a set of estimates, some of which presumably corresponding to different local maximums of the (penalized) log-likelihood function.
+If there are multiple solutions with (penalized) log-likelihoods close to the greatest found (penalized) log-likelihood, the blended identification strategy discussed in Section 3.3 can be employed to address weak identification.
+To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix E.
+Since estimation is computationally demanding, we consider the simple bivariate (structural) LSTVAR model with $p=1$ and $M=2$.
+To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both of LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
+According to the results (see Table E.2 in Appendix E for details), the estimation accuracy is slightly better when the AR matrices are inside the stability region.
+Both specifications exhibit some bias in small samples.
+However, since both the bias and the standard deviations of the estimates diminish as the sample size increases, our results align with (possible) consistency of the PML estimator.
 5 Empirical application
@@
 Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty.
-Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
-We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU).
+Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, which reduces output and raises inflation.
+These findings are supported by their recursively identified Bayesian SVAR model.
+We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU), as the level of uncertainty may affect the behavior of economic agents.
 Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable.
-This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
+This approach allows us to capture potential state-dependent effects, providing more detailed results than linear SVAR analysis.
 5.1 Data
 We consider a monthly U.S.
-dataset consisting of six variables and covering the time period from 1987:4 to 2024:12.
+dataset consisting of five variables and covering the time period from 1987:4 to 2024:12.
 Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
+In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) of Baker et al.
+(2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
+As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences.
+For measuring the price level, we use the log of consumer price index (CPI), likewise detrended by taking first differences.
+Finally, the monetary policy stance is measured with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE).
+The series of the included variables are presented in Figure 1.888The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, and the Federal funds rate are retrieved from the Federal Reserve Bank of St.
+Louis database and the Wu and Xia (2016) shadow rate from the Federal Reserve Bank of Atlanta’s website.
 The availability of the CPUI data determines the beginning of our sample period.
-In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
-As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences.
-For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences.
-Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE).
-The series of the included variables are presented in Figure 1.333The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, PPI, and the Federal funds rate are retrieved from the Federal reserve bank of St.
-Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
 Figure 1: Monthly U.S.
 time series covering the period from 1987:4 to 2024:12.
-From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods.
+From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index and consumer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods.
 The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model.
@@
 5.2 The structural LSTVAR model
-In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified.
-We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4).
-They facilitate smooth transitions between the regimes and enable the analysis of how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable.
-The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating the selection between these two weight functions within estimation.
-We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions.
-The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters.
-To avoid overfitting, particularly in the regime accommodating fewer observations, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
-To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that largely avoids making restrictive assumptions about the effects of the CPU shock.
-We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes.
-Then, to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU.
-In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively).
-The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition.
-Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
-The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU.
-Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime.
-The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1, where the fitted transition weights of High EPU Regime are presented.
-The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively.
-Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied.
-Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes.
-[^4] We compute the upper bound using the JSR toolbox (Jungers, 2023) in MATLAB.
-The obtained upper bound is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1.
-Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
-5.3 Generalized impulse response functions
-The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, which makes the conventional way of calculating impulse responses unsuitable.
-Therefore, we consider the general impulse response function (GIRF) (Koop et al., 1996) that accommodates such features and is defined as
+The structural STVAR model is specified from (3.2)-(3.3) by defining the transition weights and distributions of the shocks.
+We employ the logistic transition weights with two regimes given in Equation (2.3).
+They facilitate smooth transitions between the regimes and enable analyzing how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable.
+We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, and the shocks are assumed to follow skewed $t$-distributions.
+The autoregressive order $p=2$ is selected based on the Hannan-Quinn information criterion.
+The relatively low autoregressive order avoids overfitting, particularly in the regime accommodating fewer observations.
+To deal with the problem of weak identification discussed in Section 3.2 and 3.3, we implement a blended identification strategy that imposes the following overidentifying restrictions on the impact matrices $B_{1}$ and $B_{2}$.
+First, for each variable, one of the shocks has a greater impact effect on that variable in both regimes than the other shocks, this shock is unique to that variable, and the impact effect has the same sign in both regimes.
+Second, the shock that has the greatest impact effect on output moves inflation in the same direction in both regimes (consistent with a demand shock).
+Third, the shock that has the greatest impact effect on inflation moves output in the opposite direction in both regimes (consistent with a supply shock).
+Fourth, the shock that has the greatest impact effect on interest rate moves output in the opposite direction in both regimes (consistent with a monetary policy shock).
+The shock that has the greatest impact effect on CPUI is deemed as the CPU shock, and to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU.
+In particular, we find that President Trump’s election is accompanied by large positive CPU shocks in November 2016 and 2024.
+Moreover, there is a large negative CPU shock in January 2021, when President Biden signed executive orders to re-enter the Paris Agreement and launch a broad federal climate strategy.
+The location parameter estimate $\hat{c}=16.16$, roughly implying that Regime 1 dominates when the level of EPU is not high, while Regime 2 prevails in times of high EPU.
+Hence, we label Regime 1 as the Low EPU Regime and Regime 2 as the High EPU Regime.
+The scale parameter estimate $\hat{\gamma}=64.23$, indicating that transitions between the regimes are quite fast, which can be observed also from the bottom panel of Figure 1, where the evolution of the fitted transition weight of the High EPU Regime is depicted.
+The estimated vectors of the degrees-of-freedom and skewness parameters are $(2.77,2.71,3.00,3.15,2.03)$ and (0.38,0.41,0.16,0.15,0.11)(0.38,0.41,-0.16,0.15,0.11), respectively, in the order of the variables CPUI, EPUI, IPI, CPI, and RATE.
+Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is plausible.
+Then, we check whether our model satisfies the stationarity condition by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes (see Theorem B.1 in Appendix B).
+The obtained upper bound is strictly less than one ($0.98$), implying that our model is ergodic stationary.
+[^9] In addition, we checked that the estimated $B_{1}^{-1}B_{2}$ does not have real negative eigenvalues, so the additional condition mentioned in Footnote 5 for logistic weights is satisfied.
+Finally, based on graphical residual diagnostics, the overall adequacy of our model seems reasonable (for details, see Appendix F).
+5.3 Impulse response analysis
+The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, making the conventional way of calculating impulse responses unsuitable.
+Therefore, we consider the generalized impulse response function (GIRF) (Koop et al., 1996) that accommodates such features, defined as
 $\text{GIRF}(h,\delta_{i},\mathcal{F}_{t-1})=\text{E}[y_{t+h}|\delta_{i},\mathcal{F}_{t-1}]-\text{E}[y_{t+h}|\mathcal{F}_{t-1}],$ (5.1)
@@
 The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random.
-The GIRFs also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
-Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$.
-To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives.
-Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant (there are $366$ such histories for the Low EPU Regime and $80$ for the High EPU Regime).
+Since our model has the $p$-step Markov property, the conditioning set $\mathcal{F}_{t-1}$ can be replaced by the lag vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
+The GIRF also facilitates tracing out the effects of the shocks on the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
+To study the state-dependent effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives.
+Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant.
 For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B.
 Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
-Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead.
-From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of the High EPU Regime are depicted in each row.
-The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels.
-The left column shows the responses to conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating.
-All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points.
+Figure 2: Generalized impulse response functions to the identified CPU shock h=0,1,.,36h=0,1,....,36 months ahead.
+From top to bottom, the responses of CPUI, EPUI, IPI, CPI, RATE, and the transition weights of the High EPU Regime are depicted in each row.
+The responses of IPI and CPI are accumulated to ($100\times$)log-levels.
+The left column shows the responses conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating.
+All the GIRFs have been scaled so that the instantaneous increase in CPUI is 55 points.
 Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
-The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf.
+Figure 2 illustrates the distribution of the GIRFs in each regime in a so-called ”shotgun plot”, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf.
 Lanne and Virolainen, 2025).
-The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in the Low EPU Regime and decreases in the High EPU Regime.
-A possible explanation for the decrease in the EPUI in the High EPU Regime could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
+The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs.
+EPUI increases in both regimes, but response is stronger in the High EPU Regime.
 Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes.
-Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al.
-(2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital.
-Although many of the GIRFs exhibit a short-term increase of IPI in the High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
-In both regimes, a positive CPU shock increases consumer prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023).
-Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup.
+Huang and Punzi (2024) attribute the decline in output to firms postponing investment, whereas Fried et al.
+(2022) and Khalil and Strobel (2023) emphasize reallocation of capital away from carbon-intensive sectors.
+We find that the contraction is slightly stronger in the High EPU Regime than in the Low EPU Regime, possibly reflecting greater risk aversion among firms and financial intermediaries during periods of elevated economic policy uncertainty.
+In both regimes, a positive CPU shock increases prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023).
+Huang and Punzi (2024) find that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices preemptively to lower the probability of being stuck with a negative markup.
 However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime.
-One possible explanation is that the firms are more risk averse in the periods of high EPU, and thereby react more sensitively.
-Producer prices, in turn, mostly decrease in the Low EPU Regime but increase in the High EPU Regime.
+One possible explanation is that the firms might be more risk averse in the periods of high EPU and thereby react more sensitively.
 The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime.
-Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of the High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU.
-Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
+Finally, consistent with the GIRFs of EPUI, a positive CPU shock increases the transition weights of the High EPU Regime, more so in the High EPU Regime, where there is first a short-term decline in the weights.
+In other words, a positive CPU shock generally drives the economy towards the High EPU Regime.
+Overall, our results suggest that a positive CPU shock reduces production and raises inflation in both regimes, but its effects, particularly on inflation, are stronger under high economic policy uncertainty.
 6 Conclusion
-Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017).
-We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes.
-The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well.
-Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
-In addition to establishing the identification of the shocks, we have discussed the problem of labelling them, and employed blended identification (cf.
-Carriero et al., 2024) to address weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices.
-We have proposed estimating the model parameters by the method of penalized maximum likelihood, and introduced a three-step estimation procedure by adapting the approach of Lanne et al.
+Linear structural vector autoregressive models can be identified statistically by non-Gaussianity, provided that the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017).
+This paper is the first to extend identification by non-Gaussianity to nonlinear SVAR models.
+Specifically, we have shown that structural smooth transition vector autoregressive models, featuring a time-varying impact matrix defined as a weighted sum of regime-specific impact matrices, are likewise identified under these assumptions.
+While we have focused on exogenous, logistic, and threshold transition weights typically used in macroeconomic applications, our results can be extended to other suitable weight functions as well.
+In addition to establishing the identification of the shocks, we have discussed the problem of labeling them.
+We have also employed blended identification (cf.
+Carriero et al., 2024) to address weak identification, which in our experience is often but not exclusively related to the ordering and signs of the columns of the regime-specific impact matrices.
+We have proposed estimating the model parameters by the method of penalized maximum likelihood, and we introduced a three-step estimation procedure by adapting the approach of Lanne et al.
 (2017) to nonlinear SVAR models.
 Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model.
-The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
+The introduced methods are implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
 In an empirical application to U.S.
 data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model.
-As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
-To distinguish between periods low and high economic policy uncertainty, we used the first lag of the EPU index (Baker et al., 2016) as the switching variable.
-Our results are in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
-However, we found that while a positive CPU shock decreases production and increases inflation in times of both low and high EPU, its inflationary effects are stronger during the periods of high EPU.
-Appendix A Stationarity and ergodicity
-Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator.
+As a measure of climate policy uncertainty, we employ the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics.
+To distinguish between periods of low and high economic policy uncertainty, we use the first lag of the EPU index (Baker et al., 2016) as the switching variable.
+Our results are in line with the previous literature suggesting that a positive CPU shock reduces production and raises inflation (Khalil and Strobel, 2023 and Huang and Punzi, 2024), while Fried et al.
+(2022) found that it decreases production.
+However, we found that while a positive CPU shock reduces production and raises inflation in times of both low and high EPU, its effects, particularly on inflation, are stronger during the periods of high EPU.
+Appendix A Identification of the shocks in structural TVAR models
+The threshold VAR (TVAR) model of Tsay (1998) is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
+$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right$. (A.1)
+where $r_{m}\in\mathbb{R}$, m=1,,M1m=1,...,M-1, are the threshold parameters that satisfy $\infty<r_{1}<...<r_{M-1}<\infty$, $r_{0}\equiv$-$\infty$, $r_{M}\equiv\infty$, and $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$ is the switching variable, which we assume the be a lagged endogenous variable (up to the lag $p$).
+At each tt‚ the model defined in Equations (2.1), (2.2) and (A.1) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$.
+Notably, the TVAR model can be obtained as special case of the LSTVAR model when the smoothing parameter tends to infinity, i.e., $\gamma=\infty$ in (2.3).
+Also, exogenous transition weights may sometimes be binary or close-to-binary.
+Therefore, it is useful to establish the identification result also in such special cases of the STVAR model.
+Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and m=1,,Mm=1,...,M, specifically, either exogenous or of the threshold form (A.1).
+Then, at each $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$.
+If the allocation of the time periods to the regimes is uniquely identified, it then follows from Lemma 2 that $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns.
+Thus, to obtain identification up to ordering and signs, identification of the threshold and AR parameters is established in the following proposition, which is proven in Appendix C.4.
+Proposition A.1.
+Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 satisfied; and $\alpha_{m,t}$ either of the threshold form (A.1) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and m=1,,Mm=1,...,M.
+Suppose the following conditions hold:
+  1. (A)
+    $(\phi_{m},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
+  2. (B)
+    (for exogenous weights) there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, i=1,,Mi=1,...,M, are linearly independent.
+Then, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, m=1,,Mm=1,...,M, and (for threshold models) $r_{1},...,r_{M-1}$ are uniquely identified.
+Moreover, $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns.
+Condition (A) of Proposition A.1 implies distinguishable regimes, whereas Condition (B) guarantees (for exogenous weights) that each regime prevails in at least one time period.
+Under these conditions, the AR parameters and the partition of the time periods into the regimes are uniquely identified.
+Moreover, the regime-specific impact matrices $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns.
+[^10] The identification is for every invertible $B_{1},...,B_{M}$ and not just for almost every, as $B_{y,t}=B_{m}$ for some $m$ for all $t$, implying that $B_{y,t}$ is invertible for all invertible $B_{m}$.
+Thus, the result of Lemma 2 applies for all invertible $B_{1},...,B_{M}$.
+By Proposition A.1, the shocks are identified by fixing the ordering and signs of the columns of $B_{1},...,B_{M}$.
+In general, fixing them in one regime does not necessarily determine them in the others, and whether such constraints are purely normalizations or overidentifying depends on the distributions of the shocks.
+In practice, we recommend fixing the ordering and signs in all $B_{1},...,B_{M}$ as a part of blended identification in Section 3.3, which combines non-Gaussianity with supplementary information to address weak identification discussed in Sections 3.2 and 3.3.
+Appendix B Stationarity and ergodicity
+Establishing ergodic stationarity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator.
 Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model.
@@
 As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008).
-The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, but we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights).
-Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
+Nonetheless, we show in this section that the stationarity condition of Saikkonen (2008) applies to our model as well, given that the transition weights are either logistic or of the threshold form, with an extra assumption required for the logistic weights.
 The sufficient stationarity condition is expressed in terms of the joint spectral radius (JSR) of certain matrices. The JSR of a finite set of square matrices $\mathcal{A}$ is defined by
-$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (A.1)
+$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (B.1)
 where $\mathcal{A}^{j}=\{A_{1}A_{2}...A_{j}:A_{i}\in\mathcal{A}\}$ and $\rho(A)$ is the spectral radius of the square matrix $A$.
 Consider the companion form AR matrices of the regimes defined as
-$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (A.2)
+$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (B.2)
 The following assumption collects the sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
-Assumption A.1.
+Assumption B.1.
 Suppose the following conditions hold:
@@
   3. (III)
-    The transition weight $\alpha_{m,t}$ is an $\mathcal{F}_{t-1}$-measurable function of $\{y_{t-j},j=1,...,p\}$ for all $t$ and $m=1,...,M$.
+    The transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, follow either the threshold (A.1) or logistic (2.3) process with an endogenous switching variable $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$, and
+  4. (IV)
+    for logistic weights (with $M=2$ assumed), the matrix $B_{1}^{-1}B_{2}$ has no negative real eigenvalues.
 Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (19) of Saikkonen (2008).
-Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the process (to the best of our knowledge, there does not currently exist results presenting a sufficient necessary condition for STVAR models).
+Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the model.
 Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example.
-Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied, ruling out nonrandom exogenous weights (which are accommodated by our identification results).
-The following theorem, which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
-Theorem A.1.
-Consider the STVAR process $y_{t}$ defined in Equations (3.2) and (3.3), and suppose Assumptions 1 and A.1 are satisfied.
-Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the process $\boldsymbol{y}_{t}=(y_{t},...,y_{t-p+1})$ is a $(1+||x||^{2})$-geometrically ergodic Markov chain.
-Thus, there exists a choice of initial values $y_{-p+1},...,y_{0}$ such that the process $y_{t}$ is strictly stationary, second-order stationary, and $\beta$-mixing with geometrically decaying mixing numbers.
-Theorem A.1 is proven in Appendix A.1.
-A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere.
-To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a Lebesgue measure zero.
-Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013).
+Condition (III), in turn, restricts the form of the regime weights to either the standard threshold or logistic processes with endogenous switching, ensuring piecewise constancy or continuity of the weights.
+Finally, Condition (IV) rules out singular convex combinations of the regime-specific impact matrices in the two-regime logistic case, which guarantees positive definiteness of the conditional covariance matrix for all values of the transition weights.
+In the threshold case, positive definiteness is automatically guaranteed since the impact matrix coincides with one of the regime-specific matrices, each assumed invertible.
+The following theorem (proven in Appendix C.5), which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
+Theorem B.1.
+[…]
+Assumption B.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption B.1(I) is computationally demanding to verify in practice (see, e.g., Chang and Blondel, 2013).
 This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4.
-During estimation, this condition would have to be verified repeatedly a very large of times, making the estimation extremely tedious.
-Hence, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
-Condition A.1.
+During estimation, this condition would have to be verified repeatedly a very large number of times, making the estimation extremely tedious.
+Hence, it is useful to make use of the following condition in the estimation, which is necessary for Assumption B.1(I).
+Condition B.1.
 $\max\{\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M})\}<1$,
-where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, $m=1,...,M$.
-Condition A.1 states that the usual stability condition is satisfied by each of the regimes.
-It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein).
-After estimation, Assumption A.1(I) can be checked for the solutions of interest.
-Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013).
-The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a tight bound is required.
+where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, m=1,,Mm=1,...,M.
+Condition B.1 states that the usual stability condition is satisfied by each of the regimes.
+It is necessary for Assumption B.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein).
+After estimation, Assumption B.1(I) can be checked for the solutions of interest.
+Several methods have been proposed for computing bounds for the JSR, many of which are discussed in Chang and Blondel (2013).
+The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a relatively tight bound is required.
 The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
-Appendix B Proofs
-B.1 Proof of Lemma 1
+Appendix C Proofs
+C.1 Proof of Lemma 1
 Before proving Lemma 1, we establish the following lemma that is made use of in the proof:
-Lemma B.1.
-Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
-$P(x_{1},...,x_{k})=0$ (B.1)
-as $\mathcal{X}\equiv\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$.
-Then, the set $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$, unless Equation (B.1) is identically true.
+Lemma C.1.
+Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, i=1,,ki=1,...,k, that is not identically zero.
+Then, the set $\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$ has Lebesgue measure zero in $\mathbb{R}^{k}$.
 Proof.
-If $k=1$, the proof is trivial.
-Suppose $k>1$.
-It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},...,x_{k,0})\in\mathbb{R}^{k}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, the set $\mathcal{X}$ is an $k-1$-dimensional manifold locally in a neighbourhood of $x^{0}$.
-Because this argument applies at all points $x^{0}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, it follows that the set $\mathcal{X}$ is an $k-1$-dimensional manifold in regions where this condition holds.
-Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{k}$ in the region where the gradient of $P(x_{1},...,x_{k})$ is nonzero.
-On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P(x_{1},...,x_{k})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{k}$.
-Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$.
+Proof by induction on $k$.
+For $k=1$, the claim clearly holds, as a nontrivial univariate polynomial has finitely many zeros.
+For the induction step, suppose the set $\{(x_{1},...,x_{k-1})\in\mathbb{R}^{k-1}|P(x_{1},...,x_{k-1})=0\}$ has Lebesgue measure zero in $\mathbb{R}^{k-1}$.
+Write a point $x\in\mathbb{R}^{k}$ as $x=(x^{*},z)$ where $x^{*}=(x_{1},...,x_{k-1})\in\mathbb{R}^{k-1}$ and $z\in\mathbb{R}$ is the last component.
+Then, we can view $P(x^{*},z)$ as a polynomial in $z$, whose coefficients are polynomials in $x^{*}$:
+$P(x^{*},z)=\sum_{i=0}^{s}a_{i}(x^{*})z^{i},$ (C.1)
+where $s$ is the degree of $P$ in the variable $z$ and each coefficient $a_{i}(x^{*})$ is itself a polynomial in k1k-1 variables.
+Define $\mathcal{E}\equiv\{x^{*}\in\mathbb{R}^{k-1}:a_{0}(x^{*})=a_{1}(x^{*})=\cdots=a_{s}(x^{*})=0\}$.
+If $x^{*}\in\mathcal{E}$, then all coefficients vanish, so $P(x^{*},z)=0$ for all $z$.
+If $x^{*}\notin\mathcal{E}$, then $P(x^{*},z)$ is a nontrivial univariate polynomial in $z$.
+Because $P$ is not identically zero, at least one coefficient $a_{i}(x^{*})$, $i\in\{1,...,s\}$, is nontrivial.
+Then, by the induction hypothesis, the zero set of this nontrivial $a_{i}(x^{*})$ (and hence $\mathcal{E}\subset\{x^{*}:a_{i}(x^{*})=0\}$) has measure zero in $\mathbb{R}^{k-1}$.
+Also, for each fixed $x^{*}\in\mathbb{R}^{k-1}$, define the set $\mathcal{S}(x^{*})\equiv\{z\in\mathbb{R}:P(x^{*},z)=0\}$.
+If $x^{*}\notin\mathcal{E}$, then $P(x^{*},z)$ is a nontrivial univariate polynomial, so $\mathcal{S}(x^{*})$ is a finite set of points in $\mathbb{R}$ and thereby has measure zero.
+If $x^{*}\in\mathcal{E}$, then $\mathcal{S}(x^{*})=\mathbb{R}$.
+Now, the zero set of $P$ is $\mathcal{Z}\equiv\{(x^{*},z)\in\mathbb{R}^{k}:P(x^{*},z)=0\}$. Denoting by $\lambda_{k}(\cdot)$ the $k$-dimensional Lebesgue measure, we have by Fubini’s theorem:
+$\lambda_{k}(\mathcal{Z})=\int_{\mathbb{R}^{k-1}\times\mathbb{R}}\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}dx^{*}dz=\int_{\mathbb{R}^{k-1}}\left(\int_{\mathbb{R}}\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}dz\right)dx^{*},$ (C.2)
+where $\mathbbm{1}\{(x^{*},z)\in\mathcal{Z}\}$ is an indicator function that takes the value one if $(x^{*},z)\in\mathcal{Z}$ and the value zero otherwise.
+The inner integral on the right side of (C.2) is the one-dimensional Lebesgue measure of the slice $\{z:(x^{*},z)\in\mathcal{Z}\}=S(x^{*})$, so $\lambda_{k}(\mathcal{Z})=\int_{\mathbb{R}^{k-1}}\lambda_{1}(S(x^{*}))dx^{*}$.
+If $x^{*}\notin\mathcal{E}$, the slice $S(x^{*})$ is finite, implying that $\lambda_{1}(S(x^{*}))=0$.
+On the other hand, if $x^{*}\in\mathcal{E}$, the slice equals $\mathbb{R}$, implying that $\lambda_{1}(S(x^{*}))=\infty$.
+However, since $\mathcal{E}$ has measure zero in $\mathbb{R}^{k-1}$, it follows that $\lambda_{1}(S(x^{*}))$ is zero almost everywhere, and hence, $\lambda_{k}(\mathcal{Z})=0$, concluding the proof.
 ∎
-To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$.
-Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible.
-The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
+To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+First, note that for all $\alpha_{1,t},...,\alpha_{M,t}$ there exists (invertible) matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible, e.g., $B_{1}=...=B_{M}=I_{d}$, so $B_{y,t}=I_{d}$ (as $\sum_{m=1}^{M}\alpha_{m,t}=1$ for all $t$).
+The set of points $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ for which $B_{y,t}$ is singular is $S_{t}\equiv\{([B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$.
 Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$.
@@
 Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero.
-It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
-Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$ for each $t$.
-Hence, the union $\cup_{t=1}^{\infty}S_{t}$ is a union of a countable number of measure zero sets and thus itself has a Lebesgue measure zero.
-It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
-B.2 Proof of Lemma 2
+It then follows from Lemma C.1 that $S_{t}$ has Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
+Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has Lebesgue measure zero in $\mathbb{R}^{d\times dM}$ for each $t$.
+Hence, the union $\cup_{t\in\mathbb{N}}S_{t}$ is a union of a countable number of measure zero sets and thus itself has Lebesgue measure zero.
+It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
+C.2 Proof of Lemma 2
 The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al.
@@
 First, note that from Equation (3.2), we can rewrite the model as
-$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
-where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$, $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
-Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
-By subtracting this identify from (B.2), we obtain
-$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
+$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (C.3)
+where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$; $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$; $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$, i=1,,pi=1,...,p, are the time-varying AR matrices; and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$.
+Then, suppose Equation (C.3) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, m=1,,Mm=1,...,M, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$.
+By subtracting this identify from (C.3), we obtain
+$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (C.4)
 Taking conditional expectation conditional on $\mathcal{F}_{t-1}$ gives
-$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0,$ (B.4)
-as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (B.3), we get
-$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.5)
-Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable.
+$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0$ (C.5)
+almost surely, as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$).
+By substituting this back to (C.4), we get
+$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (C.6)
+Since $B_{y,t}$ is invertible almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable.
 By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian.
@@
 Hence, each column of $C$ has at most one nonzero element.
-Now, from the invertibility of $C$ it follows that each column of $C$ as exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element.
+Now, from the invertibility of $C$ it follows that each column of $C$ has exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element.
 Therefore, there exists a permutation matrix $P$ and a diagonal matrix $D=\text{diag}(d_{1},...,d_{d})$ with nonzero diagonal elements such that $C=DP$.
 Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.
-To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$.
+To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or 1-1.
 By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$.
 By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$.
-That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$.
-Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
-Threshold weights.
-To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3).
-Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$, $m=1,...,M$.
-By Condition (3) of Lemma 2, the sets $\mathcal{T}_{m}$ and $\tilde{\mathcal{T}}_{m}$ are nonempty for all $m=1,...,M$.
-That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
-At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$.
-Consequently, Equation (B.4) can be written as
-$\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
-where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$ at the time $t$.
-Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$.
-Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true.
-Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t}}=\phi_{m_{t}}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (as a countable union of measure zero sets has a measure zero).
-It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$.
-Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for a given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$.
-For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches.
-Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$.
-Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-This concludes that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
+That is, $d_{ii}^{2}=1$ for all i=1,,di=1,...,d, so $d_{ii}=\pm 1$.
+Since diagonal elements of $D$ are, hence, either $1$ or 1-1, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
 Logistic weights.
-Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$.
-Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$.
-The proof proceeds in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then (3) we use these results to show that the intercepts and autoregression matrices are uniquely identified.
-By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$ (irrespective of the weight function dynamics).
-By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
-$\tilde{\phi}_{1}-\phi_{1}+\alpha_{2,t}(\phi_{1}-\phi_{2})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1}-\tilde{\phi}_{2})=0$ (B.7)
-and
-$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-A_{2,i})-\tilde{\alpha}_{2,t}(\tilde{A}_{1,i}-\tilde{A}_{2,i})=0,\ \ i=1,...,p,$ (B.8)
-respectively.
-Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in two variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$ (interpreting the AR parameters as coefficients here).
-Suppose first that all of them are identically true, i.e., that all the coefficients are zero.
-In (B.7), this requires that $\tilde{\phi}_{1}=\phi_{1}$, $\phi_{1}=\phi_{2}$, and $\tilde{\phi}_{1}=\tilde{\phi}_{2}$.
-Similarly in (B.8), zero coefficients require that $\tilde{A}_{1,i}=A_{1,i}$, $A_{1,i}=A_{2,i}$ and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$.
-But as $\phi_{1}=\phi_{2}$ and $A_{1,i}=A_{2,i}$ for all $i=1,...,p$ violates Condition (2) of Lemma 2, it must be that not all scalar equations in (B.7) and (B.8) are identically true.
-For each scalar equation that is not identically true, the coefficients are not all zero, and each of these scalar equations can be expressed in the form $a_{2,j}\alpha_{2,t}=a_{0,j}+a_{3,j}\tilde{\alpha}_{2,t}$ for some $a_{0,j},a_{2,j},a_{3,j}\in\mathbb{R}$ for the $j$th not identically true equation.
-If $a_{2,j}=0$, then $\tilde{\alpha}_{2,t}$ is the constant $-a_{0,j}/a_{3,j}$ for all $t$, which violates Condition (3) of Lemma 2.
-Hence, $a_{2,j}\neq 0$ (in the not identically true equations), implying that $\alpha_{2,t}=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}$, where $a_{1,j}\equiv a_{3,j}a_{2,j}^{-1}$.
-Thus, the solution to Equations (B.7) and (B.8) must satisfy this relationship (for all $j$).
-Making the dependence of the transition weights of $z_{t}$ explicit in the notation, suppose $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds in a set of positive Lebesque measure in $z_{t}\in\mathbb{R}$.
-Since $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are logistic, $\alpha_{2,t}(z_{t})$ and $a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ are real analytic functions.
-Thus, by the Identity theorem, we have that $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}$.
-Since this must hold for all $z_{t}$, it follows from $\underset{z_{t}\rightarrow-\infty}{\lim}\alpha_{2,t}(z_{t})=0$ and $\underset{z_{t}\rightarrow-\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}$ that $a_{0,j}=0$.
-Similarly, assuming $a_{1,j}\neq 0$, it follows from $a_{0,j}=0$, $\underset{z_{t}\rightarrow+\infty}{\lim}\alpha_{2,t}(z_{t})=1$, and $\underset{z_{t}\rightarrow+\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}+a_{1,j}$ that $a_{1,j}=1$.
-If $a_{1,j}=0$, then clearly $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds only in the limit $z_{t}\rightarrow-\infty$, and thereby not in a set of positive Lebesgue measure.
-Thus, it must be that $a_{0,j}=0$ and $a_{1,j}=1$ in sets of positive Lebesgue measure, concluding that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ almost everywhere in $z_{t}\in\mathbb{R}$.
-Next, we make use of the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ to show that $\tilde{c}=c$ and $\tilde{\gamma}=\gamma$.
-By substituting the logistic weight function and simplifying, the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ can be expressed in the form
-$(\tilde{\gamma}-\gamma)z_{t}=\tilde{\gamma}\tilde{c}-\gamma c.$ (B.9)
-If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$.
-If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$.
-Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
-We have shown that for each $t$, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
-If the $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$ (rather than exogenous), we have that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ (as a countable union of measure zero sets has a measure zero).
-Since this holds for each $t$, combining it for $t=1,....,T$ yields that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
-If $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$, this result holds almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-Before showing that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are identified, we present the following lemma that is used in the proof:
-Lemma B.2.
-Consider the $T$ linear equations in $M$ variables
-$\alpha_{1,t}x_{1}+...+\alpha_{M,t}x_{M}=c_{t},\quad t=1,...,T,$ (B.10)
-where $x_{1},...,x_{M}$ are scalar variables, and $\alpha_{1,t},...,\alpha_{M,t}$ and $c_{t}$ are scalar constants.
-Suppose there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
-Then, there exists at most one solution to the system of linear equations (B.10).
-Proof.
-By assumption, there exists $M$ indices $t_{1},...,t_{M}$ such that the coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
-Now, the rows of the $(M\times M)$ matrix
-$A^{\alpha}\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\\ \end{bmatrix}$ (B.11)
-are linearly independent, and thus $A^{\alpha}$ is invertible.
-The $M$ linear equations corresponding to these indices form the system $A^{\alpha}x=c$, where $x=(x_{1},...,x_{M})$ and $c=(c_{t_{1}},...,c_{t_{M}})$.
-Because $A^{\alpha}$ is invertible, this system has exactly one solution.
-Now, any solution $x$ of the full system (B.10) for $t=1,....,T$ must satisfy the above subsystem, which has exactly one solution. Therefore, the full system (B.10) has at most one solution. ∎
-We have established that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ (and $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ for all $t$) almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
-Using these identities, Equations (B.7) and (B.8) simplify to
-$(1-\alpha_{2,t})\tilde{\phi}_{1}+\alpha_{2,t}\tilde{\phi}_{2}=(1-\alpha_{2,t})\phi_{1}+\alpha_{2,t}\phi_{2}$ (B.12)
-and
-$(1-\alpha_{2,t})\tilde{A}_{1,i}+\alpha_{2,t}\tilde{A}_{2,i}=(1-\alpha_{2,t})A_{1,i}+\alpha_{2,t}A_{2,i},\ \ i=1,...,p,$ (B.13)
-respectively.
-Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as ($\mathcal{F}_{t-1}$-measurable or nonrandom) transition weights $\alpha_{2,t}$ as coefficients here, at each $t$, each scalar equation in (B.12)-(B.13) is a linear equation in two variables given by the corresponding entries of $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$.
-Each scalar equation in (B.12)-(B.13) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and two variables.
-By Condition (3) of Lemma 2 (with $M=2$), there exists two indices $t_{1},t_{2}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},\alpha_{2,t_{i}})$, or equally $(1-\alpha_{2,t_{i}},\alpha_{2,t_{i}})$, $i=1,2$, are linearly independent.
-Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td(1+pd)$ linear equations (B.12)-(B.13), $t=1,....,T$, in $2d(1+pd)$ variables.
-Since one solution is $\tilde{\phi}_{m}=\phi_{m},\tilde{A}_{m,i}=A_{m,i}$, $i=1,...,p$, $m=1,2$, this is the unique solution.
-This concludes the proof that the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
+Next, we show the ”moreover” part for the logistic transition weights defined in (2.3) with $M=2$ and endogenous $z_{t}$, in particular, a lagged endogenous variable obtained as $z_{t}=a^{\prime}\boldsymbol{y}_{t-1}$ with a fixed $a\neq 0$ (where $a$ has exactly one element equal to one and others equal to zero).
+Consider Equation (C.5) with two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, m=1,,Mm=1,...,M, $\tilde{c},\tilde{\gamma}$.
+The proof proceeds in the following steps: (1) we show that the AR parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M are identified (without relying on possible identification of the transition weights), and then, (2) given identification of the AR parameters, the identification of the transition weights and the parameters $c,\gamma$ is established.
+Part 1.
+Denote $\boldsymbol{A}_{m}\equiv[A_{m,1}:...:A_{m,p}]$ $(d\times dp)$ and $\tilde{\boldsymbol{A}}_{m}\equiv[\tilde{A}_{m,1}:...:\tilde{A}_{m,p}]$ $(d\times dp)$, $m=1,2$.
+By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, write (C.5) as
+$\begin{aligned} \alpha_{1,t}\big{[}(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}_{t-1}+(\phi_{1}-\phi_{2})\big{]}&-\tilde{\alpha}_{1,t}\big{[}(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}_{t-1}+(\tilde{\phi}_{1}-\tilde{\phi}_{2})\big{]}\\ &+\big{[}(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}_{t-1}+(\phi_{2}-\tilde{\phi}_{2})\big{]}=0.\end{aligned}$ (C.7)
+Since by assumption the transition weights depend only on the scalar switching variable $z_{t}=a^{\prime}\boldsymbol{y}_{t-1}$ (a single component of $\boldsymbol{y}_{t-1}$), we can analyze (C.7) on level sets of $z_{t}$ where the weights become constant.
+On each such set, Equation (C.7) reduces to an affine relation in $\boldsymbol{y}_{t-1}$ with fixed coefficients, which allows us to exploit linear-algebraic arguments to identify the AR matrices and intercepts.
+Because Equation (C.7) holds with probability one for $\boldsymbol{y}_{t-1}$, and Assumption 1 implies that the distribution of $\boldsymbol{y}_{t-1}$ has a density strictly positive almost everywhere in $\mathbb{R}^{dp}$, it follows that Equation (C.7) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$.
+Below, we drop the time index from $\boldsymbol{y}_{t-1}$ and write simply $\boldsymbol{y}\in\mathbb{R}^{dp}$ for the generic lag vector when considering Equation (C.7) as a functional equation holding for almost every $\boldsymbol{y}\in\mathbb{R}^{dp}$.
+For $v\in\mathbb{R}$, let $H_{v}\equiv\{\boldsymbol{y}\in\mathbb{R}^{dp}:a^{\prime}\boldsymbol{y}=v\}$ be the level set of $z_{t}$.
+Since (C.7) holds almost everywhere in $\boldsymbol{y}\in\mathbb{R}^{dp}$, it follows that for almost every value of $v=a^{\prime}\boldsymbol{y}$, (C.7) holds for almost every $\boldsymbol{y}\in H_{v}$.
+On $H_{v}$, the weights $\alpha_{1,t}=\alpha_{1}(v)$ and $\tilde{\alpha}_{1,t}=\tilde{\alpha}_{1}(v)$ are constants, so (C.7) becomes
+$\begin{aligned} \alpha_{1}(v)\big{[}(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}+(\phi_{1}-\phi_{2})\big{]}&-\tilde{\alpha}_{1}(v)\big{[}(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}+(\tilde{\phi}_{1}-\tilde{\phi}_{2})\big{]}\\ &+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}+(\phi_{2}-\tilde{\phi}_{2})=0.\end{aligned}$ (C.8)
+Since the left side of (C.8) is affine in $\boldsymbol{y}$, if it vanishes almost everywhere in $H_{v}$, it must vanish for all $\boldsymbol{y}\in H_{v}$.
+Thus, for almost every $v\in\mathbb{R}$, Equation (C.8) holds for all $\boldsymbol{y}\in H_{v}$, i.e., for all values of the dp1dp-1 free entries orthogonal to $a$.
+To make the dependence on $v$ explicit, define
+$L(v)$ $\equiv\alpha_{1}(v)(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})-\tilde{\alpha}_{1}(v)(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2}),$ (C.9)
+$b(v)$ $\equiv\alpha_{1}(v)(\phi_{1}-\phi_{2})-\tilde{\alpha}_{1}(v)(\tilde{\phi}_{1}-\tilde{\phi}_{2})+(\phi_{2}-\tilde{\phi}_{2}),$ (C.10)
+so that (C.8) reads $L(v)\boldsymbol{y}+b(v)=0$ for all $\boldsymbol{y}\in H_{v}$. Every $\boldsymbol{y}\in H_{v}$ can be written as
+$\boldsymbol{y}=\boldsymbol{y}_{v}(v)+U\boldsymbol{y}_{\smallsetminus v},$ (C.11)
+where $\boldsymbol{y}_{\smallsetminus v}\in\mathbb{R}^{dp-1}$, $U\in\mathbb{R}^{dp\times(dp-1)}$ has columns spanning $\mathcal{S}\equiv\{u:a^{\prime}u=0\}$ (so $a^{\prime}U=0$), and $\boldsymbol{y}_{v}(v)$ is any particular vector satisfying $a^{\prime}\boldsymbol{y}_{v}(v)=v$, e.g., $\boldsymbol{y}_{v}(v)\equiv\frac{v}{\|a\|^{2}}a$.
+Substituting gives, for almost every $v$,
+$L(v)U\boldsymbol{y}_{\smallsetminus v}+(L(v)\boldsymbol{y}_{v}(v)+b(v))=0$ (C.12)
+for all $\boldsymbol{y}_{\smallsetminus v}$.
+Hence, for almost every $v$, $L(v)U=0$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)=0$.
+Each of entry of $L(v)U$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)$ is a real-analytic function of $v$ (since $\alpha_{1}(v),\tilde{\alpha}_{1}(v)$ are logistic and hence real-analytic, and $\boldsymbol{y}_{v}(v)$ is affine in $v$).
+It follows that by the Identity Theorem, both $L(v)U=0$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)=0$ hold for all $v\in\mathbb{R}$.
+Consequently, (C.8) holds for all $v$ and $\boldsymbol{y}\in H_{v}$.
+Subtracting (C.8) for two (distinct) points $\boldsymbol{y}^{(1)},\boldsymbol{y}^{(2)}\in H_{v}$ eliminates the intercepts, giving
+$L(v)u=0\quad\text{for all }u\equiv\boldsymbol{y}^{(1)}-\boldsymbol{y}^{(2)}\in\mathcal{S},$ (C.13)
+where $\mathcal{S}=\{u:a^{\prime}u=0\}$ is the (dp1)(dp-1)-dimensional subspace of vectors orthogonal to $a$, i.e., the set of displacements in $\mathbb{R}^{dp}$ that keep $a^{\prime}\boldsymbol{y}$ (and hence $z_{t}$) fixed.
+Note that the difference $u$ does not depend on the level $v$ of switching variable $z_{t}$, since the related component of $u$ is always zero by construction.
+Now, observe that if $M\in\mathbb{R}^{d\times dp}$ satisfies $Mu=0$ for all $u\in\mathcal{S}$, then there exists a unique $\beta\in\mathbb{R}^{d}$ such that [^11] To see this, pick any $w$ with $a^{\prime}w=1$ and decompose $\boldsymbol{y}=u+(a^{\prime}\boldsymbol{y})w$ with $u\in\mathcal{S}$.
+Then $M\boldsymbol{y}=(a^{\prime}\boldsymbol{y})Mw$, and setting $\beta\equiv Mw$ gives $M\boldsymbol{y}=\beta a^{\prime}\boldsymbol{y}$ for all $\boldsymbol{y}$, i.e., $M=\beta a^{\prime}$.
+To show uniqueness, suppose $M=\tilde{\beta}a^{\prime}$ holds also for some $\tilde{\beta}\in\mathbb{R}^{d}$.
+Then, $(\beta-\tilde{\beta})a^{\prime}=0$, and multiplying this by $w$ from the right gives $\beta-\tilde{\beta}=Mw-\tilde{\beta}=0$, showing $\tilde{\beta}=Mw$ and hence $\tilde{\beta}=\beta$.
+$M\boldsymbol{y}=\beta a^{\prime}\boldsymbol{y}\ \ \text{for all}\ \ \boldsymbol{y}.$ (C.14)
+For any $u\in\mathcal{S}$, Equation (C.13) can be expressed as
+$\alpha_{1}(v)(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u-\tilde{\alpha}_{1}(v)(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u=0.$ (C.15)
+For each component k=1,,dk=1,...,d, define
+$u^{(1)}_{k}\equiv[(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u]_{k},\quad u^{(2)}_{k}\equiv[(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u]_{k},\quad u^{(0)}_{k}\equiv[(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u]_{k}.$ (C.16)
+Then, the $k$th component of (C.15) is
+$\alpha_{1}(v)u^{(1)}_{k}-\tilde{\alpha}_{1}(v)u^{(2)}_{k}+u^{(0)}_{k}=0\quad\forall v\in\mathbb{R}.$ (C.17)
+Let $r(v)\equiv(\alpha_{1}(v),-\tilde{\alpha}_{1}(v),1)\in\mathbb{R}^{3}$ and $w_{k}\equiv(u^{(1)}_{k},u^{(2)}_{k},u^{(0)}_{k})\in\mathbb{R}^{3}$, so that (C.17) can be expressed as $r(v)^{\prime}w_{k}=0$, i.e., $w_{k}$ is orthogonal to the set $\{r(v):v\in\mathbb{R}\}$.
+If $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ for all $v$, then $r(v)=(\alpha_{1}(v),-\alpha_{1}(v),1)$, so the span of $r(v)$ is only $2$-dimensional.
+Then, evaluating (C.17) at two distinct values of $v$ yields linear conditions that force $w_{k}=0$.
+If $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ does not hold identically, the set $\{r(v):v\in\mathbb{R}\}$ spans the whole $\mathbb{R}^{3}$.
+[^12] If not, then there exists a vector $(c_{1},c_{2},c_{3})\neq 0$ with $c_{1}\alpha_{1}(v)-c_{2}\tilde{\alpha}_{1}(v)+c_{3}=0$ for all $v$.
+Taking the limit $v\rightarrow\infty$ yields $c_{3}=0$, as $\underset{v\rightarrow\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow\infty}{\lim}\tilde{\alpha}_{1}(v)=0$.
+Then $c_{1}\alpha_{1}(v)=c_{2}\tilde{\alpha}_{1}(v)$ for all $v$, and taking the limit $v\rightarrow$-$\infty$ yields $c_{1}=c_{2}$, as $\underset{v\rightarrow-\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow-\infty}{\lim}\tilde{\alpha}_{1}(v)=1$.
+This forces $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$, which is a contradiction.
+Therefore the only vector $w_{k}$ orthogonal to all $r(v)$ is the zero vector, so $w_{k}=0$.
+Since $w_{k}=0$ for every k=1,,dk=1,...,d, we have
+$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u=0,\ \ (\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u=0,\ \ \text{and}\ \ (\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u=0$ (C.18)
+for all $u\in\mathcal{S}$. Applying the result given in (C.14) then shows
+$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=\beta_{1}a^{\prime}\boldsymbol{y},\ \ (\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=\beta_{2}a^{\prime}\boldsymbol{y},\ \ \text{and}\ \ (\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=\beta_{0}a^{\prime}\boldsymbol{y}$ (C.19)
+for some $\beta_{0},\beta_{1},\beta_{2}\in\mathbb{R}^{d}$ and all $\boldsymbol{y}$.
+Substituting these identities into (C.8) and taking $\boldsymbol{y}\in H_{v}$ (so $a^{\prime}\boldsymbol{y}=v$) gives
+$\alpha_{1}(v)(\beta_{1}v+\phi_{1}-\phi_{2})-\tilde{\alpha}_{1}(v)(\beta_{2}v+\tilde{\phi}_{1}-\tilde{\phi}_{2})+(\beta_{0}v+\phi_{2}-\tilde{\phi}_{2})=0$ (C.20)
+for all $v$.
+For $v\neq 0$, dividing (C.20) by $v$ and taking the limit $v\rightarrow\infty$ yields $\beta_{0}=0$, as $\underset{v\rightarrow\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow\infty}{\lim}\tilde{\alpha}_{1}(v)=0$.
+Plugging this into the last equation in (C.19) gives $(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=0$ for all $\boldsymbol{y}$, and thus, $\boldsymbol{A}_{2}=\tilde{\boldsymbol{A}}_{2}$.
+With $\beta_{0}=0$, dividing (C.20) by $v$ and taking the limit $v\rightarrow$-$\infty$ yields $\beta_{1}=\beta_{2}$, as $\underset{v\rightarrow-\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow-\infty}{\lim}\tilde{\alpha}_{1}(v)=1$.
+Plugging this and $\boldsymbol{A}_{2}=\tilde{\boldsymbol{A}}_{2}$ in the first two equations in (C.19) yields $(\boldsymbol{A}_{1}-\tilde{\boldsymbol{A}}_{1})\boldsymbol{y}=0$ for all $\boldsymbol{y}$, and hence, $\boldsymbol{A}_{1}=\tilde{\boldsymbol{A}}_{1}$.
+Also, $\alpha_{1}(v)v\rightarrow 0$ as $v\rightarrow\infty$, since for $v>0$ we have $0\leq v\alpha_{1}(v)\leq ve^{-\gamma(v-c)}=\frac{v}{e^{\gamma v}}e^{\gamma c}\rightarrow 0$ as $v\rightarrow\infty$ by L’Hôpital’s rule.
+Consequently, with $\beta_{0}=0$, taking the limit $v\rightarrow\infty$ in (C.20) gives $\phi_{2}=\tilde{\phi}_{2}$.
+Since the denominator of $v(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))=\frac{v\exp\{\tilde{\gamma}(v-\tilde{c})\}-v\exp\{\gamma(v-c)\}}{(1+\exp\{\gamma(v-c)\})(1+\exp\{\tilde{\gamma}(v-\tilde{c})\})}$ tends to one as $v\rightarrow$-$\infty$, it follows that $\underset{v\rightarrow-\infty}{\lim}v(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))=e^{-\tilde{\gamma}\tilde{c}}\underset{v\rightarrow-\infty}{\lim}ve^{\tilde{\gamma}v}-e^{-\gamma c}\underset{v\rightarrow-\infty}{\lim}ve^{\gamma v}=0$ by applying L’Hôpital’s rule.
+Therefore, with $\beta_{0}=0$, $\beta_{1}=\beta_{2}$, and $\phi_{2}=\tilde{\phi}_{2}$, taking the limit $v\rightarrow$-$\infty$ in (C.20) gives $\phi_{1}=\tilde{\phi}_{1}$.
+Thus, the AR parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are uniquely identified.
+Part 2.
+Given identification of $\phi_{m}$ and $A_{m,i}$, we can substitute $\phi_{m}=\tilde{\phi}_{m}$ and $A_{m,i}=\tilde{A}_{m,i}$, $m=1,2$, i=1,,pi=1,...,p, to Equation (C.8) to obtain that, for every $v\in\mathbb{R}$,
+$(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))\big{[}\mu_{1}(\boldsymbol{y})-\mu_{2}(\boldsymbol{y})\big{]}=0\ \ \text{for all}\ \ \boldsymbol{y}\in H_{v}=\{\boldsymbol{y}\in\mathbb{R}^{dp}:a^{\prime}\boldsymbol{y}=v\},$ (C.21)
+where $\mu_{m}(\boldsymbol{y})\equiv\phi_{m}+\boldsymbol{A}_{m}\boldsymbol{y}$, $m=1,2$.
+Hence, for each fixed $v$, either $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ or $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ for all $\boldsymbol{y}\in H_{v}$.
+Considering the latter case, the equality $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ can be written as
+$(\phi_{1}-\phi_{2})+(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=0.$ (C.22)
+Subtracting for two (distinct) points $\boldsymbol{y}^{(1)},\boldsymbol{y}^{(2)}\in H_{v}$ eliminates the intercepts and gives
+$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u=0\ \ \text{for all}\ \ \boldsymbol{u}=\boldsymbol{y}^{(1)}-\boldsymbol{y}^{(2)}\in\mathcal{S},$ (C.23)
+where $\mathcal{S}=\{u:a^{\prime}u=0\}$ as before. Then, applying (C.14) with $M=\boldsymbol{A}_{1}-\boldsymbol{A}_{2}$ yields
+$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=\beta_{5}a^{\prime}\boldsymbol{y}$ (C.24)
+for some $\beta_{5}\in\mathbb{R}^{d}$ and all $\boldsymbol{y}$.
+Substituting this to (C.22) and taking $\boldsymbol{y}\in H_{v}$ so that $a^{\prime}\boldsymbol{y}=v$ implies $(\phi_{1}-\phi_{2})+\beta_{5}v=0.$
+Thus, if $\beta_{5}\neq 0$, the equality $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ can hold only at a single level $v_{0}$ that satisfies $(\phi_{1}-\phi_{2})+\beta_{5}v_{0}=0$.
+If $\beta_{5}=0$, it follows from (C.24) that $\boldsymbol{A}_{1}=\boldsymbol{A}_{2}$ and thus from (C.22) that also $\phi_{1}=\phi_{2}$, contradicting Condition (1) of Lemma 2.
+Therefore, it must be that $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ for all $v\neq v_{0}$.
+However, since $\alpha_{1}(v),\tilde{\alpha}_{1}(v)$ are real-analytic in $v$ and coincide on $\mathbb{R}\setminus\{v_{0}\}$ that has a positive Lebesgue measure, they must agree everywhere by the Identity Theorem, concluding that $\alpha_{1,t}=\tilde{\alpha}_{1,t}$ identically.
+To show identification of the parameters $\gamma$ and $c$, observe that the identity $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ can be expressed as $(\tilde{\gamma}-\gamma)v=\tilde{\gamma}\tilde{c}-\gamma c$.
+Since this holds for all $v$, it holds for two distinct values of $v$, implying that $\gamma=\tilde{\gamma}$ and $c=\tilde{c}$, concluding the identification proof logistic weights.
 Exogenous nonrandom weights.
-Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom).
-If the transition weights are binary, i.e., $\alpha_{m,t}\in\{0,1\}$ for all $m=1,...,M$ and $t=1,...,T$, then arguments similar to those used for threshold weights can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
-Next, suppose the transition weights are not binary.
-By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$.
-Substituting the exogenous weights ($\alpha_{m,t}=\tilde{\alpha}_{m,t}$ is given for all $m=1,...,M$ and $t=1,...,T$) to these equations yields
-$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})$ $=0,\quad\text{and}$ (B.14)
-$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{A}_{m,i}-A_{m,i})$ $=0,\quad i=1,...,p,$ (B.15)
-respectively.
-Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as transition weights $\alpha_{1,t},...,\alpha_{M,t}$ as constants, each scalar equations in (B.14)-(B.15) defined together for all $t=1,...,T$ defines a system of $T$ linear equations in $M$ variables.
-Therefore, arguments similar to those used for logistic weights but applied for any finite number of regimes regimes, making use of Lemma B.2, can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.∎
-B.3 Proof of Proposition 1
-We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that models before and after reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ are not observationally equivalent.
-Specifically, we say that two models are observational equivalent if and only if their joint densities are equal.
-After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
-First, we establish the following lemma and its corollary, the latter of which is used in the proof:
-Lemma B.3.
-Suppose $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ (Condition (b) of Proposition 1).
-Then, $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$, where the density function $st(\cdot;\nu_{i},\lambda_{i})$ is given in Equation (4.1).
-Moreover, if $\lambda_{i}\neq 0$, then the equation $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$, and thus, $st(x;\nu_{i},\lambda_{i})\neq st(-x;\nu_{i},\lambda_{i})$ almost everywhere in $x\in\mathbb{R}$.
-Proof.
-Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$, which is not identically zero by $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$.
-It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable.
-It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is a one-dimensional manifold locally in a neighbourhood of $x^{0}$.
-Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds.
-Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero.
-On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$.
-Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$.
-That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
-To prove the moreover part, suppose $\lambda_{i}\neq 0$ and $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$, which simplifies to
-$\left(\frac{b_{i}x+a_{i}}{1-\mathbbm{1}\{x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}=\left(\frac{-b_{i}x+a_{i}}{1-\mathbbm{1}\{-x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{-x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2},$ (B.16)
-where $a_{i}\neq 0$ and $b_{i}\neq 1$.
-If $x<-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$, a direct calculation shows that the only solution to Equation (B.16) is $x=0$.
-If $x<-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$, Equation (B.16) is a second degree polynomial equation in $x$ and has thereby at most two real solutions.
-Combining all the four possible cases, in can be concluded that $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$.
-∎
-Corollary B.1.
-Suppose $D$ is a diagonal matrix with ±$\pm 1$ diagonal entries and $P$ is a permutation matrix, and denote by $I_{d,i}$ the $i$th column of the $d$-dimensional identity matrix.
-Moreover, suppose the entries $x$ are distinct as is and after changing any of their signs.
-If $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j=1,...,d$ and $\lambda_{i}\neq 0$ for all $i=1,...,d$, and in addition, $D\neq I_{d}$ or $P\neq I_{d}$, then
-$\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})\neq\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$ (B.17)
-almost everywhere in $x\in\mathbb{R}^{d}$.
-Proof.
-Suppose $\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})=\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$. Rearranging and taking logarithm yields
-$\sum_{i=1}^{d}(\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i}))=0.$ (B.18)
-Since $DPx$ reorders or swaps the signs of some of the elements of $x$ (but does not leave all the elements unchanged), by Lemma B.3 (and strict monotonicity of logarithm),
-$f_{i}(x)\equiv\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})\neq 0$ (B.19)
-for at least one $i\in\{1,...,d\}$ over sets of positive Lebesgue measure. Therefore, Equation (B.18) can hold almost everywhere only if the nonzero terms in the sum cancel each other out.
-Denote $\mathcal{I}(x)\equiv\{i\in\{1,...,d\}|f_{i}(x)\neq 0\}$ and $g(x^{*})\equiv\sum_{i\in\mathcal{I}}f_{i}(x)$, where $x^{*}$ is the $(d^{*}\times 1)$-dimensional ($d^{*}\leq d$) subvector of $x$ containing the components of $x$ for which $I_{d,i}^{\prime}x\neq\pm I_{d,i}^{\prime}DPx$, $i=1,...,d$ (i.e., $x^{*}$ captures all variables contributing to any nonzero $f_{i}(x)$).
-It is straightforward to show (by induction) that the function $\log st(x_{i};\nu_{i},\lambda_{i})$, $x_{i}\in\mathbb{R}$, is infinitely continuously differentiable.
-It follows that also $g(x^{*})$ is infinitely continuously differentiable, and thus $d^{*}$ times continuously differentiable.
-Then, arguments based on Implicit function theorem and Sard’s theorem (Sard, 1994) can be used similarly to the proof of Lemma B.3 to conclude that the set of solutions to $g(x^{*})=0$ has a Lebesgue measure zero in $x^{*}\in\mathbb{R}^{d^{*}}$.
-Hence, as sets of zero Lebesgue measure in $\mathbb{R}^{d^{*}}\subset\mathbb{R}^{d}$ have zero Lebesgue measure in $\mathbb{R}^{d}$, Equation (B.18) can hold only in sets of zero Lebesgue measure in $x\in\mathbb{R}^{d}$, concluding the proof.
-∎
-Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model.
-Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{1},...,y_{T})$.
-Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
-$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (B.20)
-where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $st(\cdot;\nu_{i},\lambda_{i})$ is the density function the $i$th shock given in Equation (4.1).
-At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $r_{m-1}<z_{t}\leq r_{m}$, where the threshold parameters, as well as $B_{y,t}$ up to ordering and signs of its columns, are identified by Lemma 2 regardless of whether $B_{1},...,B_{M}$ are identified.
-Consequently the density (B.20) can be expressed as
-$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}),\ \ \text{and}$ (B.21)
-$f_{y}(y_{t}-\mu_{y,t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i}),$ (B.22)
-where $\mathcal{T}_{m}$, $m=1,...,M$, is the set of time indices (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and nonempty by Condition (3) of Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
-Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$.
-Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns.
-Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well as the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
-By the invertibility of $B_{m}$, almost everywhere in $y_{t}\in\mathbb{R}^{d}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{y,t})$ are equal to each other as is nor after changing the signs of some of the entries.
-Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ are in a different order.
-Likewise, changing the signs of the columns of $B_{m}$ changes the signs of the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ have different signs (in addition to possibly being in a different order).
-By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1.
-Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$.
-Therefore, as $\mathcal{T}_{1},...,\mathcal{T}_{M}$ and $\mu_{y,t}$, $t=1,...,T$ are identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ by Lemma 2 and the shocks and thereby $y_{t}$ follow continuous distributions, we have
-$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.23)
-and thus also
-$f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\boldsymbol{\tau}_{2}),...,B_{M}(\boldsymbol{\tau}_{M}))\neq f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\tilde{\boldsymbol{\tau}}_{2})...,B_{M}(\tilde{\boldsymbol{\tau}}_{M}))$ (B.24)
-almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$.
-Since this result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.
-By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$.
-Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns.
-Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed.
-Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different joint densities of the model almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, i.e., to models that are not observationally equivalent.
-Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.∎
-B.4 Proof of Proposition 2
-The proof consists of two parts.
-First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix.
-In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$.
-Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified.
-For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
-For models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
-By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},$ $y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
-With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given.
-In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
+Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom), i.e., they are given.
+Consider Equation (C.5) with the two sets of parameters $\phi_{m},A_{m,1},$ $A_{m,p}$, m=1,,Mm=1,...,M, and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, m=1,,Mm=1,...,M (and $\alpha_{m,t}=\tilde{\alpha}_{m,t}$ for all m=1,,Mm=1,...,M and $t$).
+Because Equation (C.5) holds with probability one for $\boldsymbol{y}_{t-1}$, and Assumption 1 implies that the distribution of $\boldsymbol{y}_{t-1}$ has a density strictly positive almost everywhere in $\mathbb{R}^{dp}$, it follows that Equation (C.5) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$.
+Therefore,
+$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})+\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\boldsymbol{A}}_{m}-\boldsymbol{A}_{m})\boldsymbol{y}_{t-1}=0$ (C.25)
+almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where $\boldsymbol{A}_{m}=[A_{m,1}:...:A_{m,p}]$ $(d\times dp)$ and $\tilde{\boldsymbol{A}}_{m}=[\tilde{A}_{m,1}:...:\tilde{A}_{m,p}]$ $(d\times dp)$, m=1,,Mm=1,...,M as before.
+Because the left side of (C.25) is affine in $\boldsymbol{y}_{t-1}$, if it vanishes almost everywhere, it must vanish for all $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$. Thus, for each $t$,
+$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})=0\ \ \text{and}\ \ \sum_{m=1}^{M}\alpha_{m,t}(\tilde{\boldsymbol{A}}_{m}-\boldsymbol{A}_{m})=0.$ (C.26)
+By Condition (2) of Lemma 2, there exists $M$ indices $t_{1},...,t_{M}$ for which the $(M\times M)$ matrix of weights
+$W\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\end{bmatrix}$ (C.27)
+is invertible.
+Stacking the equations in (C.26) for $t\in\{t_{1},...,t_{M}\}$ yields two full-rank systems of linear equations $(W\otimes I_{d^{2}p})(\text{vec}(\tilde{\boldsymbol{A}}_{1}-\boldsymbol{A}_{1}),...,\text{vec}(\tilde{\boldsymbol{A}}_{M}-\boldsymbol{A}_{M}))=0$ and $(W\otimes I_{d})(\tilde{\phi}_{1}-\phi_{1},...,\tilde{\phi}_{M}-\phi_{M})=0$.
+Hence, $\tilde{\boldsymbol{A}}_{m}=\boldsymbol{A}_{m}$ and $\tilde{\phi}_{m}=\phi_{m}$ for all m=1,,Mm=1,...,M, showing that the AR parameters are uniquely identified and concluding the proof.∎
+C.3 Proof of Proposition 1
+By Lemma 2, the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are uniquely identified regardless of whether $B_{1},...,B_{M}$ are identified.
+Hence, the transition weights can be treated as uniquely identified in this proof, which consists of two parts.
+Part 1.
+We show that if the ordering and signs of the columns of $B_{1}$ are fixed, then changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact matrix $B_{y,t}$ that, at some $t$, cannot be obtained by reordering or changing the signs of the columns of the impact matrix associated with the original (i.e., unmodified) matrices $B_{1},...,B_{M}$.
+Part 2.
+We combine this result with Lemma 2 to show that the matrices $B_{1},...,B_{M}$ are uniquely identified.
+Part 1.
 Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively.
 Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the invertible impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns.
-Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
-We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
-Suppose $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
-At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
-$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.25)
-for some $(d\times d)$ diagonal matrix $D_{t}$ with ±$\pm 1$ diagonal elements and some $(d\times d)$ permutation matrix $P_{t}$ (that may vary in time). Rearranging the above equation gives
-$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})D_{t}P_{t})=-\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t}).$ (B.26)
-Suppose first $D_{t}P_{t}=I_{d}$. Equation (B.27) then simplifies to
-$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))=0.$ (B.27)
-By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
-At each such time period $t$, Equation (B.27) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other.
-Hence, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) forms a subspace in $\mathbb{R}^{d\times d(M-1)}$ with the dimension of this subspace being less than $(M-1)d^{2}$, as it must satisfy a nontrivial linear equation.
-Consequently, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) has a Lebesgue measure zero in $[B_{2}:\cdots:B_{M}]\in\mathbb{R}^{d\times d(M-1)}$, implying that the set of matrices $B_{1},...,B_{M}$ satisfying this equation have a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
-Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.26) as $C_{t}\in\mathbb{R}^{d\times d}$.
-By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
-At each such time period $t$, the right side of Equation (B.26) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros.
-Equation (B.26) is thereby nontrivial and can be expressed as
-$B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t})=\alpha_{1,t}^{-1}C_{t}$ (B.28)
-The set of matrices $B_{1}$ that satisfy Equation (B.28) forms a subspace in $\mathbb{R}^{d\times d}$ with the dimension of this subspace being less than $d^{2}$, as it must satisfy a nontrivial linear equation.
-It follows that the set of matrices $B_{1}$ that satisfy Equation (B.28) has Lebesgue measure zero on $\mathbb{R}^{d\times d}$, and therefore, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.28) has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$.
-By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.26) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
-Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.27) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
-This concludes that, almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, the matrix $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$
-Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
+Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well as the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
+We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $B_{1},...,B_{M}]\in\mathbb{R}^{d\times dM}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any m=2,,Mm=2,...,M.
+[…]
+$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (C.28)
+for some $(d\times d)$ diagonal matrix $D_{t}$ with $\pm 1$ diagonal elements and some $(d\times d)$ permutation matrix $P_{t}$ (that may depend on $t$). Rearranging the above equation gives
+$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})D_{t}P_{t})=-\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t}).$ (C.29)
+In the logistic case with $\gamma\in(0,\infty)$, $\alpha_{2,t}=[1+\exp\{-\gamma(z_{t}-c)\}]^{-1}\in(0,1)$ for all $t$, and hence also $\alpha_{1,t}=1-\alpha_{2,t}\in(0,1)$ for all $t$.
+From this, and in the case of exogenous weights from Condition (b) of Proposition 1, it follows that the weight vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$, say, $t^{*}$, such that none of its entries is zero.
+If $D_{t^{*}}P_{t^{*}}=I_{d}$, Equation (C.29) implies that
+$\sum_{m=2}^{M}\alpha_{m,t^{*}}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))=0.$ (C.30)
+Because $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})\neq B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ for some m=2,,Mm=2,...,M, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (C.30) is the solution space of a nontrivial homogeneous linear system in the entries of $B_{2},...,B_{M}$.
+This solution space is a proper linear subspace of $\mathbb{R}^{d\times d(M-1)}$ and therefore has Lebesgue measure zero.
+Consequently, the set of $B_{1}:...:B_{M}]$ satisfying (C.30) has Lebesgue measure zero in $\mathbb{R}^{d\times dM}$.
+Suppose then $D_{t^{*}}P_{t^{*}}\neq I_{d}$ and denote the left side of Equation (C.29) as $C_{t}\in\mathbb{R}^{d\times d}$.
+At time period $t^{*}$, the right side of Equation (C.29) is nonzero (as $B_{1}$ is invertible), thus, implying that also $C_{t^{*}}$ is not a zero matrix.
+Equation (C.29) is thereby nontrivial and can be expressed as
+$B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t^{*}}P_{t^{*}})=\alpha_{1,t^{*}}^{-1}C_{t^{*}}$ (C.31)
+Since $\text{rank}(I_{d}-D_{t^{*}}P_{t^{*}})\geq 1$, the set of matrices $B_{1}$ that satisfy Equation (C.31) defines either an empty set or an affine subspace in $\mathbb{R}^{d\times d}$ with the dimension of this subspace being less than $d^{2}$.
+It follows that the set of matrices $B_{1}$ that satisfy Equation (C.31) has Lebesgue measure zero in $\mathbb{R}^{d\times d}$, and therefore, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (C.31) has Lebesgue measure zero in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+[…]
+Part 2.
+Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
 Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$.
 By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$).
-Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A) of Proposition 2), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$.
-Hence, models before and after reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ are not be observationally equivalent.
-The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A) of Proposition 2, implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$).
-Thus, $B_{y,t}$ is uniquely identified at all $t$.
-Since $B_{y,t}$ is uniquely identified and two models are observationally equivalent if and only if they are observationally equivalent at all $t$, a unique identification of the matrices $B_{1},...,B_{M}$ is achieved if an impact matrix that is identical to $B_{y,t}$ for all $t$ cannot be obtained with any other invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$.
-Suppose an impact matrix identical to $B_{y,t}$ for all $t$ is obtained with some invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$.
+Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (a) of Proposition 1), we then obtain from the above results that almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}$.
+Hence, models before and after reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ are not observationally equivalent.
+The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under the conditions of Proposition 1, implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$.
+Thus, $B_{y,t}$ is uniquely identified at all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
+Consider the impact matrix $B_{y,t}$ based on invertible $(d\times d)$ matrices $B_{1},...,B_{M}$, and suppose an impact matrix identical to $B_{y,t}$ for all $t$ is obtained with some other invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$.
 Then,
-$\sum_{m=1}^{M}\alpha_{m,t}(B_{m}-\tilde{B}_{m})=0,$ (B.29)
-for all $t$.
-Denote the $ij$th element of $\tilde{B}_{m}$ as $\tilde{b}_{ij}^{(m)}$.
-At each $t$, Equation (B.29) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$.
-The weights $\alpha_{m,t}$ are treated as constant coefficients here (some of which may be zero), as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here.
-Each scalar equation in (B.29) is a linear equation in $M$ variables.
-Thus, each scalar equation in (B.29) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and $M$ variables.
-By Condition (3) of Lemma 2 (assumed in Proposition 2), there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
-Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td^{2}$ linear equations (B.29), $t=1,....,T$, in $Md^{2}$ variables.
-Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified.
-For exogenous models, the identification of the transition weights is given.
-For logistic models, by Lemma 2, the transition weights weights are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$, concluding the proof.∎
-B.5 Proof of Theorem A.1
-Saikkonen (2008) establishes stationarity, ergodicity, and mixing properties of a STVEC model by making use of the theory of Markov chains.
-Kheifets and Saikkonen (2020), in turn, show that the results of Saikkonen (2008) are applicable to their STVAR model, as it can be obtained as a special case of the STVEC model of Saikkonen (2008).
-However, the results of Saikkonen (2008) and Kheifets and Saikkonen (2020) are not directly applicable to our structural STVAR model, because our model defines the conditional covariance matrix differently to the STVEC model of Saikkonen (2008) and STVAR model of Kheifets and Saikkonen (2020).
-Hence, certain steps need to be taken show the applicability of the results of Saikkonen (2008) to our structural STVAR model.
-Note that by Assumption A.1(II), Assumption 1 in Saikkonen (2008) is satisfied.
+$\sum_{m=1}^{M}\alpha_{m,t}\Delta B_{m}=0,$ (C.32)
+for all $t$, where $\Delta B_{m}\equiv B_{m}-\tilde{B}_{m}$.
+In the case of logistic weights, $M=2$ so $\alpha_{1,t}=1-\alpha_{2,t}$ and Equation (C.32) reads
+$(1-\alpha_{2,t})\Delta B_{1}+\alpha_{2,t}\Delta B_{2}=0.$ (C.33)
+Because $0<\gamma<\infty$ and (endogenous) $z_{t}$ is real-valued and non-constant, the logistic map $\alpha_{2,t}$ attains at least two distinct values in $(0,1)$.
+Evaluating (C.33) at two such values, $\alpha^{a}\neq\alpha^{b}$, and subtracting gives $(\alpha^{a}-\alpha^{b})(\Delta B_{2}-\Delta B_{1})=0$, implying that $\Delta B_{2}=\Delta B_{1}$.
+Substituting back to (C.33) then yields $\Delta B_{1}=\Delta B_{2}=0$, i.e., $\tilde{B}_{1}=B_{1}$ and $\tilde{B}_{2}=B_{2}$.
+In the exogenous case (with any $M\geq 2$), by Condition (2) of Lemma 2 there exist $M$ indices $t_{1},...,t_{M}$ such that the $(M\times M)$ weight matrix $W$
+$W\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\end{bmatrix}$ (C.34)
+is invertible. Writing Equation (C.32) at these $M$ indices gives the linear system
+$\sum_{m=1}^{M}\alpha_{m,t_{i}}\Delta B_{m}=0,\quad i=1,...,M,$ (C.35)
+which can be expressed as $(W\otimes I_{d^{2}})(\text{vec}(\Delta B_{1}),...,\text{vec}(\Delta B_{M}))=0$.
+Since $W$ is invertible, $W\otimes I_{d^{2}}$ has full rank, so $\Delta B_{m}=0$, i.e., $\tilde{B}_{m}=B_{m}$, for all m=1,,Mm=1,...,M.
+This concludes that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.∎
+C.4 Proof of Proposition A.1
+We start by showing that under the conditions of Proposition A.1, the AR and threshold parameters are identified.
+Then, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns.
+In the case of exogenous binary weights, unique identification of $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, is given by Lemma 2 (whose conditions for exogenous weights are satisfied under conditions of Proposition A.1).
+Consider the threshold transition weights defined in (A.1) with the switching variable $z_{t}=a^{\prime}\boldsymbol{y}_{t-1}$, where $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ and $a$ is a $(dp\times 1)$ vector with exactly one of its elements equal to one and others equal to zero.
+At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the active regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$.
+Consider Equation (C.5) with two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, m=1,,Mm=1,...,M, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
+Note that $r_{0}=\tilde{r}_{0}=-\infty,r_{M}=\tilde{r}_{M}=\infty$ and $\infty<r_{1}<...<r_{M-1}<\infty$ by assumption.
+Denote $J_{m}\equiv\{z_{t}\in\mathbb{R}:r_{m-1}<z_{t}\leq r_{m}\},\tilde{J}_{n}\equiv\{z_{t}\in\mathbb{R}:\tilde{r}_{n-1}<z_{t}\leq\tilde{r}_{n}\}$, and $\dot{J}_{m,n}\equiv J_{m}\cap\tilde{J}_{n}$ for m,n=1,,Mm,n=1,...,M.
+Because $z_{t}$ has strictly positive density almost everywhere (by Assumption 1 and invertibility of the impact matrix in the active regime), each $J_{m}$ has positive Lebesgue measure.
+Consequently, since $\{\tilde{J}_{n}\}_{n=1}^{M}$ partition $\mathbb{R}$, for every m=1,,Mm=1,...,M, there exists at least one $n$ where $\dot{J}_{m,n}$ has positive Lebesgue measure.
+In this overlap event $\alpha_{m,t}=\tilde{\alpha}_{n,t}=1$, so Equation (C.5) gives
+[…]
+where $\boldsymbol{A}_{m}=[A_{m,1}:...:A_{m,p}]$ $(d\times dp)$ and $\tilde{\boldsymbol{A}}_{n}=[\tilde{A}_{n,1}:...:\tilde{A}_{n,p}]$ $(d\times dp)$.
+By Assumption 1 and invertibility of the impact matrix in the active regime, the lag vector $\boldsymbol{y}_{t-1}$ has strictly positive density almost everywhere in $\mathbb{R}^{dp}$, implying that it has strictly positive density also a.e.
+in $\{\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}:z_{t}\in\dot{J}_{m,n}\}$.
+Since an affine function that equals zero a.e.
+on a set with positive Lebesgue measure must have all coefficients zero, we have $\tilde{\phi}_{n}=\phi_{m}$ and $\tilde{\boldsymbol{A}}_{n}=\boldsymbol{A}_{m}$ for $\dot{J}_{m,n}$ with positive measure.
+Now, if there are two positive measure sets $\dot{J}_{m,n_{1}}$ and $\dot{J}_{m,n_{2}}$ with $n_{1}\neq n_{2}$, it follows that $(\tilde{\phi}_{n_{1}},\tilde{\boldsymbol{A}}_{n_{1}})=(\tilde{\phi}_{n_{2}},\tilde{\boldsymbol{A}}_{n_{2}})$ violating Condition (A) of Proposition A.1.
+Hence, for each $m$ there exists a unique $n=n(m)$ with $\dot{J}_{m,n(m)}$ that has positive measure.
+Since $\bigcup_{n=1}^{M}\dot{J}_{m,n}=J_{m}$ and only single $\dot{J}_{m,n(m)}$ can have positive Lebesgue measure, it follows that $J_{m}\subseteq\tilde{J}_{n(m)}$ almost everywhere.
+By symmetric arguments, we get that also $\tilde{J}_{n(m)}\subseteq J_{m}$ almost everywhere, so $J_{m}=\tilde{J}_{n(m)}$ almost everywhere.
+Because the symmetric difference between the half-open intervals $J_{m}$ and $\tilde{J}_{n(m)}$ has thereby Lebesgue measure zero, the end points must coincide, as otherwise the symmetric difference would contain a non-degenerate interval.
+Thus, $r_{m-1}=\tilde{r}_{n(m)-1}$ and $r_{m}=\tilde{r}_{n(m)}$.
+Since the thresholds are ordered by assumption so that regimes cannot be ’relabelled’, it follows that $n(m)=m$.
+Hence, $(\tilde{\phi}_{m},\tilde{A}_{1,m},...,\tilde{A}_{p,m})=(\phi_{m},A_{1,m},...,A_{p,m})$ for all m=1,,Mm=1,...,M and $\tilde{r}_{m}=r_{m}$ for all m=1,,M1m=1,...,M-1, concluding that the AR and threshold parameters are uniquely identified.
+By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$, given that it is invertible.
+Since the transition weights are binary, $B_{y,t}=B_{m}$ for some $m$ for all $t$.
+Since each $B_{m}$ is assumed invertible, it follows that $B_{y,t}$ is invertible for all $t$.
+Hence, for each regime $m$, the associated impact matrix $B_{m}$ is identified up to ordering and signs of its columns, concluding the proof.∎
+C.5 Proof of Theorem B.1
+As explained in Appendix B, our STVAR model is not nested to the STVEC model of Saikkonen (2008), and hence, additional steps are needed to show the applicability of the results of Saikkonen (2008) to our structural STVAR model.
+Note that by Assumption B.1(II), Assumption 1 in Saikkonen (2008) is satisfied.
 Moreover, by the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$, Assumption 2(a) in Saikkonen (2008) satisfied, and as we assume that the set $\text{l}_{2}$ in Saikkonen (2008) is an empty set, also Assumption 2(b) is satisfied.
-Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption A.1(I) also Condition (19).
+Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption B.1(I) also Condition (19).
 The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
@@
 Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ (with the notation of this paper).
-That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), reduce to the first term on the right hand side of Equation (3.2) defining our structural STVAR model.
-The third term on the right side of Equation (17) in Saikkonen (2008), however, reduces to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes.
+That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), simplify to the first term on the right hand side of Equation (3.2) defining our structural STVAR model.
+The third term on the right side of Equation (17) in Saikkonen (2008), however, simplifies to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes.
 This is different to the second term of Equation (3.2), which takes the form $\sum_{m=1}^{M}\alpha_{m,t}B_{m}e_{t}$ when the impact matrix (3.3) is substituted to it.
@@
 Saikkonen (2008) argues that thereby by Theorem 2.2(ii) of Cline and Phu (1998), the Markov chain $Z_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, which is then used to conclude further results.
-Concerning our model, note that by Lemma 1, $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, implying that the conditional covariance matrix $B_{y,t}(\boldsymbol{y}_{t-1})B_{y,t}(\boldsymbol{y}_{t-1})^{\prime}$ is positive definite, where the dependence of the conditional covariance matrix on the $(dp\times 1)$ random vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is made explicit in the notation.
-Hence, the smallest eigenvalue of $B_{y,t}(x)B_{y,t}(x)^{\prime}$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$.
+Concerning our model, note that by Assumption B.1(III), the regime weights are either of the threshold or logistic form.
+In the threshold case, denoting a generic lag vector as $\boldsymbol{y}\in\mathbb{R}^{dp}$, the impact matrix $B_{y,t}(\boldsymbol{y})=\sum_{m=1}^{M}\alpha_{m,t}(\boldsymbol{y})B_{m}$ is exactly $B_{m}$ within each regime, and since each $B_{m}$ is assumed invertible, the conditional covariance matrix $B_{y,t}(\boldsymbol{y})B_{y,t}(\boldsymbol{y})$’ is positive definite.
+In the logistic two-regime case, Condition (IV) rules out singular convex combinations of $B_{1}$ and $B_{2}$, which likewise guarantees that $B_{y,t}(\boldsymbol{y})$ is invertible for all $\boldsymbol{y}$.
+[^13] Indeed, denoting $C\equiv B_{1}^{-1}B_{2}$ and $B(\alpha)=(1-\alpha)B_{1}+\alpha B_{2}=B_{1}((1-\alpha)I+\alpha C)$, the following conditions are equivalent: (1) $B(\alpha)$ is singular for some $\alpha^{*}\in(0,1)$ and (2) $C$ has a negative real eigenvalue.
+To prove this, observe first that $\det B(\alpha)=\det(B_{1})\det((1-\alpha)I+\alpha C)$.
+As $B_{1}$ is invertible, it follows that $B(\alpha)$ is singular iff $(1-\alpha)I+\alpha C$ is singular, i.e., iff there exists $v\neq 0$ with $((1-\alpha)I+\alpha C)v=0$, which is equivalent to $Cv=-\frac{1-\alpha}{\alpha}v$.
+For $\alpha\in(0,1)$, the scalar $\frac{1-\alpha}{\alpha}$ is real and negative.
+Hence, if $B(\alpha^{*})$ is singular for some $\alpha^{*}\in(0,1)$, then $C$ has a negative real eigenvalue $\tfrac{1-\alpha^{*}}{\alpha^{*}}$.
+Conversely, if $Cv=\lambda v$ with $\lambda<0$, choose $\alpha^{*}=\frac{1}{1-\lambda}\in(0,1)$.
+Then, $(1-\alpha^{*})+\alpha^{*}\lambda=0$, so $(1-\alpha^{*})I+\alpha^{*}C$ (and hence $B(\alpha^{*}))$ is singular, establishing the equivalence.
+Therefore, and because in the logistic case $B_{y,t}(\boldsymbol{y})$ is continuous in $\boldsymbol{y}$, the smallest eigenvalue of $B_{y,t}(\boldsymbol{y})B_{y,t}(\boldsymbol{y})$’ attains a strictly positive minimum on every compact subset of $\mathbb{R}^{dp}$.
+Thus, the conditional covariance matrix is uniformly bounded away from singularity on compact subsets as required.
 It follows from this, Equation (3.2), and the invertibility of $B_{y,t}$ that we can apply Theorem 2.2(ii) of Cline and Phu (1998) to conclude that the Markov chain $\boldsymbol{y}_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, similarly to the process $Z_{t-1}$ in Saikkonen (2008).
-Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p.
-358) is satisfied.
-To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$ is bounded from above by a finite real number, where $j$ is a positive integer, $\mathcal{A}=\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\}$, and $\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}$ as well as the set $\mathcal{A}^{j}$ are defined in Appendix A.
+[…]
 Similarly to Saikkonen (2008, p. 313), clearly,
-$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (B.30)
-where |||||$\cdot$||| | ⋅ | | is Eucledian norm, $N$ is a positive integer, and $c$ a finite real number.
-Hence, because $B_{y,t}(x)$ is a constant here, by Jensen’s inequality and submultiplicativity of norms, we have
-$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E[||e_{t}||]^{2},$ (B.31)
-where the right side is bounded from above by some finite constant (by (B.30) and the finite second moment of $e_{t}$ given by Assumption 1).
+$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (C.37)
+where $||\cdot||$ is Euclidean norm, $N$ is a positive integer, and $c$ a finite real number. By Jensen’s inequality and submultiplicativity of norms, we have
+$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(\boldsymbol{y})e_{t}||^{2}\middle|\boldsymbol{y}=x\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E||e_{t}||^{2},$ (C.38)
+where the right side is bounded from above by some finite constant (by (C.37) and the finite second moment of $e_{t}$ given by Assumption 1).
 Thus, an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model.
 Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
-Appendix C Penalized nonlinear least squares estimation of STVAR models
+Appendix D Penalized nonlinear least squares estimation of STVAR models
 This appendix describes how the parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of nonlinear least squares (NLS) as explained in Hubrich and Teräsvirta (2013, Section 4.2) for logistic STVAR models.
 After describing the NLS estimation, penalized nonlinear least squares (PNLS) estimation is considered by introducing a penalization term to the sum of squares of residuals function.
-C.1 Nonlinear least squares estimation
+D.1 Nonlinear least squares estimation
 Consider the following sum of squares of residuals function
-$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (C.1)
+$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (D.1)
 where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters.
-[^5] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$.
+[^14] With threshold weights (A.1), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.3) weights, $\alpha=(c,\gamma)$.
 If the transition weights are exogenous, we simply treat them as known constants.
-The NLS estimates of the parameters $\phi_{m},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
-If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ (they depend on) needs to be estimated.
+The NLS estimates of the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
+If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ they depend on needs to be estimated.
 Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only.
-However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are not sufficient.
-Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Appendix 4.3, the following procedure is adopted.
-First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model.
-As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model at all.
+However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are insufficient.
+Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Section 4.3, the following procedure is adopted.
+First, determine the values of the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, t=1,,Tt=1,...,T, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model.
+As an extreme example, transition weights with $\alpha_{1,t}=0$ for all t=1,,Tt=1,...,T are not reasonable, as Regime 1 would not be involved in the model at all.
 Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
-One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$.
-In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime.
-Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$.
-Denote the large enough sum as $\tilde{T}$.
-We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough.
-[^6] We acknowledge that our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
+One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, m=1,,Mm=1,...,M.
+In the special case of discrete transition weights such as the threshold weights (A.1), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime.
+It can then be determined whether each regime contributes enough to the model by checking, for all m=1,,Mm=1,...,M, whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime.
+Denoting the large enough sum as $\tilde{T}$, transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all m=1,,Mm=1,...,M are deemed reasonable enough.
+[^15] In our empirical application (Section 5), we assume that there should be at least three times as many observations from all the $d$ variables combined, implying that $\tilde{T}=3kd^{-1}$, where $k$ is the number of parameters in a regime.
+We acknowledge that this assumption is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
 Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range.
-For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches.
+For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number to some large positive number that implies almost discrete regime-switches.
 With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$.
@@
 Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$.
-Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$.
+Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all m=1,,Mm=1,...,M.
 Denote this subset as $\mathcal{W}$.
-Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1},....,\phi_{M}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$.
+Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, and $\alpha$.
 Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
-C.2 Penalized nonlinear least squares estimation
-Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (C.1), the penalized sum of squares function is minimized.
+D.2 Penalized nonlinear least squares estimation
+Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (D.1), the penalized sum of squares function is minimized.
 We define the penalized sum of squares function as
-$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (C.2)
+$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (D.2)
 where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space.
 To focus on avoiding estimates that do not satisfy the usual stability condition for all regimes, we define the penalty term as
-$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (C.3)
+$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (D.3)
 where
-$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (C.4)
+$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (D.4)
 is the minimized residual sum of squares, $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization.
-Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues.
-The penalty term (C.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
-The penalized NLS estimates of $\phi_{m},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2).
+Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (D.3) is greater than zero, and it increases in the modulus of these eigenvalues.
+The penalty term (D.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
+The penalized NLS estimates of $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$ (where $\mathcal{W}$ is the set of grid point defined in Section D.1), then calculating the penalty term $P(\alpha)$ for all $\alpha\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (D.2).
 In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
-Appendix D A Monte Carlo study
+Appendix E A Monte Carlo study
 This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4.
-We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable.
-To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$.
-For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
-For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1.
-In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1).
-In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region.
-To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus).
-The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
+We consider two specifications of a simple bivariate (structural) LSTVAR model with $p=1$ and $M=2$, with the first lag of the first variable as the switching variable.
+To maintain some constancy, we use the same impact matrices across both specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, the same location and scale parameter values $c=0.8$ and $\gamma=5.0$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$.
+We consider two specifications of the AR parameters, presented in Table E.1.
+In the first specification (LSTVAR 1), the AR matrices are inside the stability region (see Section 4.1).
+In the second specification (LSTVAR 2), the AR matrices are in the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region.
+To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table E.1.
+As in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one of the roots of the companion form AR matrix is larger than $0.95$ in modulus.
+The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table E.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
 Regime 1 Regime 2
-$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
-[…]
-[…]
-[…]
-[…]
-Table D.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study.
-Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
-For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$.
-Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4.
-After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values.
+$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $\boldsymbol{A}_{2}||$
+LSTVAR 1 $0.30$ $0.70$ 0.30-0.30 $0.00$ $0.58$ $\phantom{-}1.20$ $0.50$ $0.20$ $\phantom{-}2.00$ $0.74$
+$0.60$ $0.20$ $\phantom{-}0.40$ $1.00$ $0.58$ 1.10-1.10 $0.30$ $0.50$ 1.00-1.00 $0.26$
+LSTVAR 2 $0.30$ $1.10$ 0.30-0.30 $0.00$ $0.97$ $\phantom{-}0.72$ $0.74$ $0.20$ $\phantom{-}2.00$ $0.98$
+$0.20$ $0.20$ $\phantom{-}0.80$ $1.00$ $0.97$ 0.87-0.87 $0.30$ $0.73$ 1.00-1.00 $0.49$
+Table E.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the LSTVAR models LSTVAR 1 and LSTVAR 2 in the Monte Carlo study.
+Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $\boldsymbol{A}_{m}||$, $m=1,2$.
+For each of the two model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$.
+Then, we fit the first-order two-regime LSTVAR model to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4.
+After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is negative and the second one positive, in line with the true parameter values.
 To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
-$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
-TVAR 1 $\phi_{1}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$0.00$ $(0.07)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$\phi_{2}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
-$0.01$ $(0.28)$ $0.00$ $(0.19)$ $0.00$ $(0.13)$ $0.00$ $(0.09)$ $0.00$ $(0.04)$
-$A_{1,1}$ $-0.01$ $(0.08)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$A_{2,1}$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$
-$0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$B_{1}$ $0.20$ $(0.52)$ $0.09$ $(0.33)$ $0.05$ $(0.21)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
-$-0.08$ $(0.34)$ $-0.04$ $(0.19)$ $-0.02$ $(0.11)$ $-0.01$ $(0.06)$ $0.00$ $(0.02)$
-$-0.02$ $(0.09)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.03$ $(0.10)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$B_{2}$ $0.30$ $(0.61)$ $0.14$ $(0.39)$ $0.07$ $(0.24)$ $0.03$ $(0.13)$ $0.00$ $(0.04)$
-$0.04$ $(0.15)$ $0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
-$-0.03$ $(0.18)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$r_{1}$ $0.00$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$ $0.00$ $(0.00)$
-$\nu_{1}$ $0.05$ $(0.53)$ $0.02$ $(0.34)$ $0.01$ $(0.23)$ $0.00$ $(0.16)$ $0.00$ $(0.07)$
-$\nu_{2}$ $165.03$ $(700.13)$ $105.25$ $(757.09)$ $13.98$ $(174.19)$ $1.48$ $(6.32)$ $0.16$ $(1.40)$
-$\lambda_{1}$ $-0.03$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$\lambda_{2}$ $0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-TVAR 2 $\phi_{1}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
-$-0.05$ $(0.14)$ $-0.04$ $(0.11)$ $-0.02$ $(0.07)$ $-0.02$ $(0.04)$ $-0.01$ $(0.02)$
-$\phi_{2}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
-$-0.01$ $(0.30)$ $0.00$ $(0.18)$ $0.01$ $(0.12)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$
-$A_{1,1}$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.00)$
-$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.03$ $(0.06)$ $-0.02$ $(0.04)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
-$A_{2,1}$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$0.00$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.02$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$B_{1}$ $0.38$ $(0.64)$ $0.16$ $(0.40)$ $0.08$ $(0.24)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
-$-0.15$ $(0.43)$ $-0.07$ $(0.25)$ $-0.04$ $(0.15)$ $-0.02$ $(0.08)$ $0.00$ $(0.03)$
-$-0.04$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.05$ $(0.18)$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
-$B_{2}$ $0.55$ $(0.73)$ $0.24$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
-$0.06$ $(0.17)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.13)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
-$-0.05$ $(0.25)$ $-0.01$ $(0.08)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$r_{1}$ $0.08$ $(0.45)$ $0.02$ $(0.24)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
-$\nu_{1}$ $0.05$ $(2.77)$ $-0.07$ $(0.29)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
-$\nu_{2}$ $107.17$ $(542.02)$ $93.32$ $(1273.27)$ $10.09$ $(81.02)$ $1.29$ $(5.10)$ $0.09$ $(1.47)$
-$\lambda_{1}$ $-0.04$ $(0.11)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
-$\lambda_{2}$ $-0.01$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4.
-The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
-The results are based on $500$ Monte Carlo repetitions.
 $T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
@@
 $0.00$ $(0.25)$ $0.00$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
-$\phi_{2}$ $-0.01$ $(0.17)$ $-0.01$ $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
-$-0.02$ $(0.39)$ $-0.02$ $(0.24)$ $-0.01$ $(0.15)$ $0.00$ $(0.11)$ $0.00$ $(0.05)$
+$\phi_{2}$ 0.01-0.01 $(0.17)$ 0.01-0.01 $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
+0.02-0.02 $(0.39)$ 0.02-0.02 $(0.24)$ 0.01-0.01 $(0.15)$ $0.00$ $(0.11)$ $0.00$ $(0.05)$
 $A_{1,1}$ $0.00$ $(0.15)$ $0.01$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.02$ $(0.26)$ $-0.01$ $(0.13)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
+0.02-0.02 $(0.26)$ 0.01-0.01 $(0.13)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
 $0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.01$ $(0.10)$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.10)$ 0.01-0.01 $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
 $A_{2,1}$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $0.00$ $(0.14)$ $0.00$ $(0.09)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
-$-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.08)$ 0.01-0.01 $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $B_{1}$ $0.24$ $(0.60)$ $0.10$ $(0.37)$ $0.04$ $(0.20)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
-$-0.09$ $(0.46)$ $-0.05$ $(0.24)$ $-0.02$ $(0.13)$ $-0.01$ $(0.08)$ $0.00$ $(0.03)$
-$-0.03$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.05$ $(0.14)$ $-0.03$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+0.09-0.09 $(0.46)$ 0.05-0.05 $(0.24)$ 0.02-0.02 $(0.13)$ 0.01-0.01 $(0.08)$ $0.00$ $(0.03)$
+0.03-0.03 $(0.12)$ 0.02-0.02 $(0.08)$ 0.01-0.01 $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+0.05-0.05 $(0.14)$ 0.03-0.03 $(0.08)$ 0.01-0.01 $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
 $B_{2}$ $0.44$ $(0.72)$ $0.19$ $(0.42)$ $0.08$ $(0.22)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
 $0.05$ $(0.17)$ $0.02$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
-$-0.03$ $(0.19)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+0.03-0.03 $(0.19)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $c$ $0.03$ $(0.19)$ $0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $\gamma$ $28.86$ $(121.66)$ $2.90$ $(25.52)$ $0.09$ $(0.75)$ $0.05$ $(0.47)$ $0.01$ $(0.19)$
-$\nu_{1}$ $-0.02$ $(0.49)$ $-0.02$ $(0.32)$ $-0.02$ $(0.21)$ $-0.01$ $(0.15)$ $0.00$ $(0.07)$
+$\nu_{1}$ 0.02-0.02 $(0.49)$ 0.02-0.02 $(0.32)$ 0.02-0.02 $(0.21)$ 0.01-0.01 $(0.15)$ $0.00$ $(0.07)$
 $\nu_{2}$ $152.80$ $(597.75)$ $43.27$ $(266.76)$ $8.00$ $(52.02)$ $1.46$ $(5.55)$ $0.11$ $(1.37)$
-$\lambda_{1}$ $-0.04$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$\lambda_{1}$ 0.04-0.04 $(0.09)$ 0.02-0.02 $(0.05)$ 0.01-0.01 $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $\lambda_{2}$ $0.02$ $(0.10)$ $0.01$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-LSTVAR 2 $\phi_{1}$ $-0.01$ $(0.18)$ $-0.01$ $(0.11)$ $-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
-$-0.10$ $(0.19)$ $-0.06$ $(0.13)$ $-0.03$ $(0.08)$ $-0.02$ $(0.06)$ $-0.02$ $(0.02)$
+LSTVAR 2 $\phi_{1}$ 0.01-0.01 $(0.18)$ 0.01-0.01 $(0.11)$ 0.01-0.01 $(0.08)$ 0.01-0.01 $(0.05)$ $0.00$ $(0.02)$
+0.10-0.10 $(0.19)$ 0.06-0.06 $(0.13)$ 0.03-0.03 $(0.08)$ 0.02-0.02 $(0.06)$ 0.02-0.02 $(0.02)$
 $\phi_{2}$ $0.07$ $(0.67)$ $0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
 $0.12$ $(1.18)$ $0.02$ $(0.19)$ $0.01$ $(0.13)$ $0.01$ $(0.08)$ $0.01$ $(0.04)$
-$A_{1,1}$ $-0.05$ $(0.13)$ $-0.02$ $(0.07)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $0.00$ $(0.01)$
-$-0.03$ $(0.11)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+$A_{1,1}$ 0.05-0.05 $(0.13)$ 0.02-0.02 $(0.07)$ 0.01-0.01 $(0.03)$ 0.01-0.01 $(0.02)$ $0.00$ $(0.01)$
+0.03-0.03 $(0.11)$ 0.02-0.02 $(0.08)$ 0.01-0.01 $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $0.03$ $(0.13)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
-$-0.03$ $(0.08)$ $-0.03$ $(0.06)$ $-0.02$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
-$A_{2,1}$ $-0.02$ $(0.11)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.02$ $(0.21)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.06)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
-$-0.03$ $(0.13)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+0.03-0.03 $(0.08)$ 0.03-0.03 $(0.06)$ 0.02-0.02 $(0.03)$ 0.01-0.01 $(0.02)$ 0.01-0.01 $(0.01)$
+$A_{2,1}$ 0.02-0.02 $(0.11)$ 0.01-0.01 $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+0.02-0.02 $(0.21)$ 0.01-0.01 $(0.06)$ 0.01-0.01 $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.06)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
+0.03-0.03 $(0.13)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $B_{1}$ $0.44$ $(0.74)$ $0.16$ $(0.41)$ $0.08$ $(0.24)$ $0.03$ $(0.12)$ $0.01$ $(0.05)$
-$-0.14$ $(0.55)$ $-0.06$ $(0.27)$ $-0.04$ $(0.18)$ $-0.02$ $(0.09)$ $0.00$ $(0.03)$
-$-0.04$ $(0.13)$ $-0.02$ $(0.08)$ $-0.01$ $(0.06)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
-$-0.03$ $(0.19)$ $-0.02$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$
+0.14-0.14 $(0.55)$ 0.06-0.06 $(0.27)$ 0.04-0.04 $(0.18)$ 0.02-0.02 $(0.09)$ $0.00$ $(0.03)$
+0.04-0.04 $(0.13)$ 0.02-0.02 $(0.08)$ 0.01-0.01 $(0.06)$ 0.01-0.01 $(0.03)$ $0.00$ $(0.01)$
+0.03-0.03 $(0.19)$ 0.02-0.02 $(0.09)$ 0.02-0.02 $(0.05)$ 0.01-0.01 $(0.04)$ $0.00$ $(0.02)$
 $B_{2}$ $0.64$ $(0.84)$ $0.25$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
 $0.08$ $(0.19)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
-$-0.01$ $(0.09)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
-$-0.04$ $(0.23)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
+0.01-0.01 $(0.09)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
+0.04-0.04 $(0.23)$ 0.01-0.01 $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
 $c$ $0.23$ $(0.80)$ $0.05$ $(0.31)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
 $\gamma$ $79.82$ $(349.54)$ $22.77$ $(90.88)$ $5.71$ $(44.05)$ $0.39$ $(4.28)$ $0.10$ $(0.30)$
-$\nu_{1}$ $-0.11$ $(0.57)$ $-0.07$ $(0.30)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
+$\nu_{1}$ 0.11-0.11 $(0.57)$ 0.07-0.07 $(0.30)$ 0.05-0.05 $(0.21)$ 0.03-0.03 $(0.15)$ $0.00$ $(0.07)$
 $\nu_{2}$ $192.21$ $(889.55)$ $38.30$ $(233.50)$ $9.58$ $(79.39)$ $1.37$ $(4.71)$ $0.15$ $(1.42)$
-$\lambda_{1}$ $-0.05$ $(0.09)$ $-0.02$ $(0.06)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
+$\lambda_{1}$ 0.05-0.05 $(0.09)$ 0.02-0.02 $(0.06)$ 0.01-0.01 $(0.04)$ 0.01-0.01 $(0.03)$ $0.00$ $(0.01)$
 $\lambda_{2}$ $0.00$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
-Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4.
+Table E.2: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4.
 The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
 The results are based on $500$ Monte Carlo repetitions.
-The results of the Monte Carlo study are presented in Table D.2 for the TVAR models and in Table D.3 for the LSTVAR models.
-The results show that there is some but mostly very small bias, which diminishes when the sample size increases.
-Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with consistency of the PML estimator.
+The results of the Monte Carlo study are presented in Table E.2.
+They show that there is some but mostly small bias, which diminishes when the sample size increases.
+Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with (possible) consistency of the PML estimator.
 Nonetheless, for certain parameters, there appears to be a notable bias in small samples.
@@
 This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution.
-Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values.
+Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of appropriately discriminating between large and very large parameter values.
 Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation.
-Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the large upward bias should not be a concern.
-Second, in the LSTVAR models, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small.
-A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are smooth in the true data generating process ($\gamma=5$).
+However, the large upward bias does not seem concerning, as there is virtually no difference in the economic interpretations of large and very large degrees-of-freedom parameter values (with the latter amplifying the bias).
+Second, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small.
+A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are relatively smooth in the true data generating process ($\gamma=5$).
 The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
-Third, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small.
-A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters.
+Third, the estimates of the first element of the first row of $B_{1}$ and $B_{2}$ exhibit a notable upward bias when the sample is small.
+A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$ for the related shock).
 In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.
-In general, the small-sample bias of the PML estimator seems to be slightly larger the specifications where the true AR matrices are in the penalization region (TVAR 2 and LSTVAR 2) than the specifications where the true AR matrices lie well inside the stability region (TVAR 1 and LSTVAR 1).
-The differences are, however, only small, and even in small samples, there does not appear to be any particular issues with the estimation accuracy of the AR matrices lying inside the penalization region.
-Also, the TVAR models seem to be slightly more accurately estimated than the LSTVAR models, possibly because, unlike in TVAR models, in LSTVAR models also the smoothness of regime-switches needs to be estimated.
-Appendix E Details on the empirical application
-The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a) in which the introduced methods have been implemented.
+In general, the small-sample bias of the PML estimator seems to be slightly larger the specification where the true AR matrices are in the penalization region (LSTVAR 2) than the specification where the true AR matrices lie well inside the stability region (LSTVAR 1).
+However, the differences are quite small, and even in small samples, there does not appear to be notably issues with the estimation accuracy of the AR matrices lying inside the penalization region.
+Appendix F Details on the empirical application
+The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a), which implements the introduced methods.
 The R package sstvars also contains the dataset studied in the empirical application.
-In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because labelling them effectively requires labelling the corresponding structural shocks, and only the CPU shock has been labelled.
-To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure E.1.
-There is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the lag zero CCFs between the fourth and fifth residual series.
+The upper bound for the joint spectral radius (see Appendix B), in turn, is computed with the JSR toolbox (Jungers, 2023) in MATLAB.
+To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure F.1.
+There is not much autocorrelation left in the residuals, but there are large correlation coefficients (CC) in the lag zero CCFs between the residuals of CPUI and EPUI.
 In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small.
-Hence, with the exception of the single large CC in the lag zero CCFs of the fourth and fifth residual series, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
-To study our model’s adequacy to capture the conditional heteroscedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.2.
-There is a very large CC in the lag zero CCFs between the second and third residual series, which, however, disappears if the COVID-19 residual from March 2020 is excluded (not shown).
+Hence, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
+To study our model’s adequacy to capture the conditional heteroskedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure F.2.
+There are particularly large CC in the lag one CCFs between the IPI and EPUI, in the lag one CCF between the RATE and EPUI, and the lag zero CCF between RATE and EPUI.
 In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series.
-Thus, there are some inadequacies in capturing the conditional heteroscedasticity, but with the exception of the large CC in CCFs between the second and third residual series (due to the extreme COVID-19 volatility), the inadequacies do not seem particularly severe.
-Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates are presented in Figure E.3 for each residual series.
-They show that the marginal distributions of the series are reasonably well captured, although there is a single large outlier in the third residual series (which disappears if the March 2020 residual is removed; not shown).
+Thus, there are clearly some inadequacies in capturing the conditional heteroskedasticity, but the vast majority of the CCs are small.
+Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom and skewness parameter values given by their estimates are presented in Figure F.3 for each residual series.
+They show that the marginal distributions of the series are reasonably well captured, although there is a single very large outlier in the IPI’s residuals (from March 2020), which is related to the vast and fast drop in production caused by the COVID-19 lockdown.
 Therefore, the overall adequacy of the model seems, in our view, reasonable.
-Figure E.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$.
+Figure F.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime second-order LSTVAR model for the lags 0,1,,240,1,...,24.
 The lag zero autocorrelation coefficients are omitted, as they are one by convention.
-The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
-Figure E.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$.
+The dashed lines are the $95\%$ bounds $\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
+Figure F.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime second-order LSTVAR model for the lags 0,1,,240,1,...,24.
 The lag zero autocorrelation coefficients are omitted, as they are one by convention.
-The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
-Figure E.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime sixth-order LSTVAR model.
-The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates.
+The dashed lines are the $95\%$ bounds $\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
+Figure F.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime second-order LSTVAR model.
+The quantile-quantile plots are based on tt–distributions with zero mean, variance one, and degrees-of-freedom and skewness parameter values given by their estimates.